### v1

# On the Dirichlet problem for the degenerate $k$-Hessian equation

Abstract.
This paper investigates the existence of a global $C^{1,1}$ solution to the Dirichlet problem for the $k$-Hessian equation with a nonnegative right-hand side $f$, focusing on the required conditions for $f$. A well-known counterexample demonstrates the sharpness of the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in a domain $\Omega_{0}\Supset\Omega$. We establish the existence for all $2\leq k\leq n-1$ under the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$, and under an a priori assumption that $\Delta u\geq 1$. Surprisingly, higher smoothness of $f$ permits a better exponent to be achieved. The condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$ is also sharp, as demonstrated by the same counterexample. For the Monge-Ampère equation (the case $k=n$), we establish the existence under the optimal condition $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$. For any $5\leq k\leq n-1$, we prove the existence under the condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$, and under the a priori assumption $\Delta u\geq 1$. Moreover, we obtain the existence for all $2\leq k\leq n-1$ under the condition $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$.
Key words and phrases:
Monge-Ampère equation, $k$-Hessian equation, Dirichlet problem, Existence theory, Optimal degeneracy.
2020 Mathematics Subject Classification:
35J60, 35J70, 35J25, 35J96.
1. Introduction
In this paper, we consider the Dirichlet problem for the degenerate $k$-Hessian equation:
$\begin{cases}\sigma_{k}\big[D^{2}u\big]=f\geq 0\quad\text{in}\ \Omega,\\ u=\varphi\quad\text{on}\ \partial\Omega,\end{cases}$ (1.1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ ($n\geq 2$), $2\leq k\leq n$, and $D^{2}u$ denotes the Hessian matrix of the function $u$. For an $n\times n$ real symmetric matrix $A$, we denote by $\sigma_{k}[A]$ the $k$-th elementary symmetric polynomial of its eigenvalues $\lambda(A)=(\lambda_{1},\dots,\lambda_{n})$. That is,
$\sigma_{k}[A]=\sigma_{k}(\lambda(A)):=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}},\quad\text{for}\ k=1,2,\dots,n.$
Equivalently, $\sigma_{k}[A]$ is the sum of all the $k\times k$ principal minors of $A$. By convention, we set $\sigma_{0}(\lambda)\equiv 1$.
Following the work of Caffarelli-Nirenberg-Spruck [2], a function $u\in C^{2}(\Omega)\cap C^{0}(\overline{\Omega})$ is called $k$-admissible if its Hessian eigenvalues satisfy
$\lambda\big(D^{2}u(x)\big)\in\Gamma_{k}\quad\text{for all}\ x\in\Omega,$
where $\Gamma_{k}$ denotes the Gårding cone defined by
$\Gamma_{k}:=\left\{\lambda\in\mathbb{R}^{n}:\ \sigma_{j}(\lambda)>0\ \text{for}\ j=1,2,\dots,k\right\}.$
They established that for any $\lambda\in\Gamma_{k}$ and each $1\leq i\leq n$,
$\frac{\partial\sigma_{k}}{\partial\lambda_{i}}>0;$
that $\sigma_{k}^{1/k}(\lambda)$ is a concave function in $\Gamma_{k}$; and that
$\text{the operator}\ \,\sigma_{k}^{1/k}[A]\,\ \text{is elliptic and concave in}\ \ \left\{A\in\mathscr{S}^{n}:\ \lambda(A)\in\Gamma_{k}\right\},$
where $\mathscr{S}^{n}$ denotes the space of $n\times n$ real symmetric matrices. A hypersurface in $\mathbb{R}^{n}$ is called $k$-convex (for $k\in\{1,2,\dots,n-1\}$) if its principal curvatures $\kappa=(\kappa_{1},\kappa_{2},\dots,\kappa_{n-1})$ satisfy $\sigma_{j}(\kappa)\geq 0$ for all $j\leq k$ everywhere. It is strictly $k$-convex if $\sigma_{j}(\kappa)\geq\delta>0$ for some constant $\delta$ uniformly on the hypersurface for all $j\leq k$. As shown in [2], a necessary condition for the existence of a $C^{2}(\overline{\Omega})$ solution to (1.1) is that the boundary $\partial\Omega$ be strictly (k1)(k-1)-convex when viewed as a hypersurface in $\mathbb{R}^{n}$.
The $k$-Hessian equation is called degenerate if the nonnegative function $f$ is allowed to vanish at some points in $\overline{\Omega}$, and non-degenerate if $\inf_{\Omega}f>\delta$ for some positive constant $\delta$. This paper focuses on the degenerate case. The condition $f^{1/k}\in C^{1,1}(\overline{\Omega})$ is natural when exploiting the concavity of the operator. However, the condition $f^{1/k}\in C^{1,1}$ may not be optimal. Regarding the optimal condition, a counterexample for the Monge-Ampère equation was given by Wang [32]:
$\begin{cases}\det D^{2}u=\bar{f}(x):=\eta\left(\frac{x_{n}}{|x^{\prime}|^{\alpha}}\right)\left|x^{\prime}\right|^{\beta}\quad\text{in}\ B_{1},\\ u=0\quad\text{on}\ \partial B_{1},\end{cases}$ (1.2)
where $\alpha>1$, $\beta>0$, and
$\eta(t):=\begin{cases}e^{-1/(1-t^{2})},\quad|t|<1,\\ 0,\quad|t|\geq 1.\end{cases}$
Wang [32] proved that if $\beta<2(n-1)(\alpha-1)$, then (1.2) admits no $C^{1,1}$ solution. Taking $\beta=2(n-1)(\alpha-1)-1$, we have $\bar{f}\geq 0$ in $B_{2}$, and
$\bar{f}^{\frac{1}{n-1-\frac{2n-1}{2\alpha}}}\in C^{1,1}(\overline{B_{2}}),\quad\bar{f}^{\frac{1}{n-1}}\in C^{1,1-\frac{2n-1}{\alpha(n-1)}}(\overline{B_{2}}),\quad\forall\alpha\geq 4,$
and
$\bar{f}^{\frac{1}{\frac{2(n-1)}{3}-\frac{2n-1}{3\alpha}}}\in C^{2,1}(\overline{B_{2}}),\quad\bar{f}^{\frac{3}{2(n-1)}}\in C^{2,1-\frac{6n-3}{2\alpha(n-1)}}(\overline{B_{2}}),\quad\forall\alpha\geq 6.$
Ivochkina-Trudinger-Wang [16] pointed out that a modification of the counterexample (1.2) also applies to the $k$-Hessian equations for $2\leq k\leq n-1$, although no proof was provided. A detailed proof of this fact was given by Dinew-Pliś-Zhang [5]; namely, the problem
$\begin{cases}\sigma_{k}\big[D^{2}u\big]=\bar{f}\quad\text{in}\ B_{1},\\ u=0\quad\text{on}\ \partial B_{1},\end{cases}$
has no $C^{1,1}$ solution whenever $\beta<2(k-1)(\alpha-1)$, for all $2\leq k\leq n-1$. Consequently, for any $k\in\{2,3,\dots,n\}$,
$f^{\frac{1}{k-1}}\in C^{1,1}\quad\text{and}\quad f^{\frac{3}{2(k-1)}}\in C^{2,1}$
are two best possible conditions.
For the Monge-Ampère equation, Guan [9] reduced the estimate of second-order derivatives to the boundary under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, and obtained the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ under the homogeneous boundary condition ($\varphi=0$). Subsequently, by introducing a completely new method, Guan-Trudinger-Wang [12] established a global $C^{2}$ estimate for the Monge-Ampère equation under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, which led to the existence of a convex solution in $C^{1,1}(\overline{\Omega})$. For the case $2\leq k\leq n-1$ of equation (1.1), Ivochkina-Trudinger-Wang [16] proposed the following open problem:
[…]
Dong [6] solved the problem (1.3) under the homogeneous boundary condition. For general boundary conditions, Jiao-Wang [17] derived a global $C^{2}$ estimate for convex solutions to problem (1.3). However, since smooth approximations of convex viscosity solutions do not necessarily preserve convexity, they [17] were unable to prove the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for equation (1.1).
In this paper, we investigate for equation (1.1) both problem (1.3) and the following problem: for any $k\in\{2,3,\dots,n\}$,
[…]
where the exponent p=2(k1)/3p=2(k-1)/3 is best possible. We solve problem (1.4) for the Monge-Ampère equation under the optimal condition p=2(n1)/3p=2(n-1)/3, and for general $k$-Hessian equations ($2\leq k\leq n-1$) with $p=2k/3$. Furthermore, we address problem (1.3) and problem (1.4) ($5\leq k\leq n-1$) under their respective best possible conditions, and under the following additional a priori assumption: there exist small positive constants $\delta_{0}$ and $r_{0}$ such that
$\Delta u\geq\delta_{0}\quad\text{in}\ \Omega\cap B_{r_{0}}(y),$ (1.5)
for all $y\in\partial\Omega$ at which $\Delta u(y)\gg 1$.
In order to find a global $C^{1,1}$ $k$-admissible solution in degenerate problems, the usual strategy is to establish global $C^{2}$ a priori estimates for the non-degenerate case that are independent of the positive lower bound of $f$. Then, based on the existence theory of Caffarelli-Nirenberg-Spruck [2] for the non-degenerate case, the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for the degenerate case is obtained by an approximation argument. There is an extensive body of work on this problem. In addition to the references cited, we refer the reader to [3, 15, 18, 19, 28, 29, 30, 11, 8, 13, 31, 25, 4]. This is, however, a far from complete list of important contributions.
Below we state the main results. First, we establish global $C^{1}$ estimate and boundary estimates for the second-order derivatives in the tangential and tangential-normal directions.
Theorem 1.1.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{1}(\overline{\Omega})}+\sup_{\partial\Omega}\left(|u_{\nu\eta}|+|u_{\eta\eta}|\right)\leq C,$
where $\nu$ denotes the unit interior normal vector, $\eta$ is any unit tangential vector; and $C$ depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Second, we reduce the estimate of second-order derivatives to the boundary.
Theorem 1.2.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\Omega}\left|D^{2}u\right|\leq\sup_{\partial\Omega}\left|D^{2}u\right|+C,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Under the homogeneous boundary condition, Dong [6] established the required second-order normal derivative estimate on the boundary. Combining this result with Theorems 1.1 and 1.2 yields the global $C^{2}$ estimate.
Theorem 1.3.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi=0$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.4.
Let $\partial\Omega\in C^{3}$ be strictly (k1)(k-1)-convex, $\varphi=0$, $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Results corresponding to Theorems 1.1, 1.2, 1.3 and Corollary 1.4 were established by Jiao-Wang [17] under the weaker condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega})$ with $f\geq 0$ in $\Omega$. In this paper, we focus on problems (1.3) and (1.4) with general boundary conditions. The primary challenge, therefore, is to estimate the second-order derivatives in the normal direction on the boundary. For general $k$-Hessian equations, we employ the framework of Ivochkina-Trudinger-Wang [16] to establish the boundary second-order normal derivative estimates.
Ivochkina-Trudinger-Wang [16] provided a PDE-based proof of Krylov’s results [21, 22, 23, 24], and roughly speaking, their ideas are similar. These results and ideas were clarified very well in [16], and we quote them below.
A global upper bound, independent of the positive lower bound of $f$, for the second-order derivatives of admissible solutions to the Dirichlet problem of Hessian equations, and more general Bellman equations, was obtained by Krylov [20], by a probabilistic argument, and by an analytic proof in a series of papers Krylov [21, 22, 23, 24]. Roughly speaking, Krylov’s proof [21, 22, 23, 24] consists of two steps. One is the weakly interior estimate, that is for any positive constants $\varepsilon,\ \delta>0$, there exists a constant $C_{\varepsilon,\delta}$, depending on $n$, $\varphi,\ \partial\Omega,\ \|u\|_{C^{1}(\overline{\Omega})},\ \|f\|_{C^{1,1}(\overline{\Omega})}$, and in particular the upper bound of the boundary second-order tangential and tangential-normal derivatives of $u$, such that
$\sup_{\Omega_{\delta}}\left|D^{2}u\right|\leq\varepsilon\sup_{\partial\Omega}\left|D^{2}u\right|+C_{\varepsilon,\delta},$ (1.6)
where $\Omega_{\delta}:=\{x\in\Omega:\ \operatorname{dist}(x,\partial\Omega)>\delta\}$. The other one is the boundary estimate in terms of the interior one, that is for any $\delta>0$, there exists a constant $C_{\delta}$, depending in addition on $C_{\varepsilon_{0},\delta}$ in (1.6), with arbitrarily given $\varepsilon_{0}>0$, such that
$\sup_{\partial\Omega}\left|D^{2}u\right|\leq C_{\delta}\left(1+\sup_{\partial\Omega_{\delta}}\left|D^{2}u\right|\right).$ (1.7)
Third, we establish the weakly interior estimate for degenerate $k$-Hessian equations.
Theorem 1.5.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. For any constant $\varepsilon>0$, there exists a positive constant $C_{\varepsilon}$ such that
$\left|D^{2}u(x)\right|\leq\frac{1}{d(x)}\left(\varepsilon\sup_{\partial\Omega}\left|D^{2}u\right|+C_{\varepsilon}\right),\quad\forall x\in\Omega,$ (1.8)
where $d(x):=\operatorname{dist}(x,\partial\Omega)$, and the constant $C_{\varepsilon}$ depends on $\varepsilon$, $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Fourth, we obtain the boundary estimate in terms of the interior one for degenerate $k$-Hessian equations under the condition (1.5). Then the global $C^{2}$ estimate is established.
Theorem 1.6.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Let the condition (1.5) hold. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Remark 1.7.
The condition (1.5) is naturally satisfied if $f$ has a positive lower bound near $\partial\Omega$.
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.8.
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, and $f\geq 0$ in $\Omega_{0}$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Let the condition (1.5) hold. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Fifth, if $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$, then the condition (1.5) can be omitted. As pointed out by Guan-Trudinger-Wang [12], for applications to problems in differential geometry, it is desirable to impose no restriction on the nonnegative function $f$ apart from smoothness. Guan [9] established the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ for the Monge-Ampère equation with the homogeneous boundary condition when $n=3$ and $0\leq f\in C^{3,1}(\mathbb{R}^{3})$. Compared to the case when $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it is surprising that the exponent of $f$ can be improved provided its smoothness is higher, for $k\geq 4$.
Theorem 1.9.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ and $\inf_{\Omega_{0}}f>0$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.10.
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Remark 1.11.
There is no inclusion relationship between the conditions $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{1/(k-q)}\in C^{2,1}(\overline{\Omega_{0}})$ for any $0<q<k$. This is demonstrated by two counterexamples. One counterexample is $f^{1/(k-q)}:=|x|^{2}\in C^{\infty}(\overline{\Omega_{0}})$, where the origin $0\in\partial\Omega$. It is clear that $f^{1/k}=|x|^{2(k-q)/k}\notin C^{1,1}(\overline{\Omega_{0}})$. The other counterexample is
$f^{1/k}:=\left(1+g(x)\right)e^{-1/|x-x_{0}|}$
in $\overline{\Omega_{0}}$, where the origin 0 and $x_{0}$ are two distinct points on $\partial\Omega$, and
$g(x):=\begin{cases}|x|^{4}\left(1+\sin\frac{1}{|x|}\right),\quad x\neq 0,\\ 0,\quad x=0.\end{cases}$
One can verify that $g\in W^{2,\infty}$ which implies $g\in C^{1,1}(\overline{\Omega_{0}})$, and that $D^{2}g$ is not continuous at the origin. Consequently, $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$, but $D^{2}(f^{1/(k-q)})$ fails to be continuous at the origin.
For the Monge-Ampère equation, using the convexity of the admissible solutions and of the domain, as well as the affine invariance of the equation, Guan-Trudinger-Wang [12] introduced a completely new method to estimate the second-order normal derivatives on the boundary. Recently, Jiao-Wang [17] extended this method to convex solutions of the $k$-Hessian equation in uniformly convex domains. Sixth, applying the method of Guan-Trudinger-Wang [12], we solve problem (1.4) for the Monge-Ampère equation under the optimal condition.
Theorem 1.12.
Let $u\in C^{3,1}(\overline{\Omega})$ be a convex solution of equation (1.1) with $k=n$, $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $\inf_{\Omega_{0}}f>0$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Remark 1.13.
The condition in $\Omega_{0}$ of Theorem 1.12 (rather than in $\Omega$) is only used to reduce the estimate of second-order derivatives to the boundary, as established in Theorem 1.2.
The existence of a convex solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.14.
Let $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique convex solution in $C^{1,1}(\overline{\Omega})$ for (1.1) with $k=n$.
The rest of the paper is organized as follows. Section 2 presents preliminary results. Section 3 establishes the global $C^{1}$ estimate and the boundary second-order tangential and tangential-normal estimates (Theorem 1.1). Section 4 reduces the estimate of second-order derivatives to the boundary (Theorem 1.2). Section 5 is devoted to the weakly interior estimate, namely Theorem 1.5. Section 6 develops the boundary estimate in terms of the interior one to finish Theorems 1.6 and 1.9. Section 7 solves problem (1.4) for the Monge-Ampère equation under the optimal condition, namely Theorem 1.12. The Appendix (Section 8) contains auxiliary lemmas.
In this paper, the dependence of constants on dimension $n$ and the number $k$ in (1.1) is omitted, and a constant depending only on $n$ and $k$ is called universal. This section ends by listing some notations.
Notation.
1. $e_{i}=(0,\dots,1,\dots,0)$ denotes the $i$-th standard coordinate vector.
2. $x^{\prime}=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$.
3. $B_{r}(x_{0})=\{x\in\mathbb{R}^{n}:\ |x-x_{0}|\leq r\}$ and $x_{0}$ is omitted when $x_{0}=0$.
4. $\operatorname{dist}(E,F)=\text{distance from $E$ to $F$}$, $\forall E,\ F\subset\mathbb{R}^{n}$.
5. $\mathbb{S}^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
6. $\mathbf{I}_{n}$ denotes the $n\times n$ identity matrix.
7. $A^{\prime}=A^{T}$ denotes the transpose of the matrix $A$.
8. $A^{1/2}$ denotes the square root of positive semi-definite matrix $A$.
9. $\nabla_{x^{\prime}}=(\partial_{1},\partial_{2},\dots,\partial_{n-1})$.
10. $D^{2}_{x^{\prime}x^{\prime}}$ denotes the $(n-1)\times(n-1)$ matrix with elements $\partial_{ij}$ where $1\leq i,\ j\leq n-1$.
11. $\nu=\nu(x)$ denotes the unit interior normal vector of $\partial\Omega$ at $x\in\partial\Omega$.
2. Preliminaries
The following pointwise inequalities are fundamental in this paper.
Lemma 2.1.
Let $g$ be a function satisfying $g>0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. We have
$(i)$ If $g\in C^{1,1}(\overline{\Omega_{0}})$, then
$\frac{\left|\nabla g(x)\right|^{2}}{g(x)}\leq K,\quad\forall x\in\overline{\Omega},$ (2.1)
where $K$ depends on $\|g\|_{C^{1,1}(\overline{\Omega_{0}})}$ and $\operatorname{dist}(\Omega,\partial\Omega_{0})$, but is independent of $\inf_{\Omega_{0}}g$.
$(ii)$ If $g\in C^{2,1}(\overline{\Omega_{0}})$, then for any $\alpha<1/2$,
$\partial_{ee}g(x)-\alpha\frac{\left|\partial_{e}g(x)\right|^{2}}{g(x)}\geq-Kg^{\frac{1}{3}}(x),\quad\forall x\in\overline{\Omega},\ e\in\mathbb{S}^{n-1},$ (2.2)
where $K$ depends on $\alpha$, $\|g\|_{C^{2,1}(\overline{\Omega_{0}})}$, and $\operatorname{dist}(\Omega,\partial\Omega_{0})$, but is independent of $\inf_{\Omega_{0}}g$.
Proof.
In fact, we only need to consider the case of one dimension. Let $h(x)$ be a function in $C^{2,1}(\mathbb{R})$ and $h>0$ on $\mathbb{R}$. Denote $\beta:=h(0)>0$.
Step 1: we prove conclusion $(i)$.
It holds that
$\begin{cases}0\leq h\big(\sqrt{\beta}\big)=\beta+h^{\prime}(0)\sqrt{\beta}+O(\beta)\\ 0\leq h\big(-\sqrt{\beta}\big)=\beta-h^{\prime}(0)\sqrt{\beta}+O(\beta).\end{cases}$
Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq 1+K,$
where $K=\|h\|_{C^{1,1}(\mathbb{R})}$. By virtue of the above, (2.1) are valid at points where $(g(x))^{1/2}<\operatorname{dist}(\Omega,\partial\Omega_{0})$. And (2.1) hold clearly at points where $(g(x))^{1/2}\geq\operatorname{dist}(\Omega,\partial\Omega_{0})$.
Step 2: we prove conclusion $(ii)$.
It holds that
$0\leq h\big(\beta^{t_{1}}\big)+h\big(-\beta^{t_{1}}\big)=2\beta+h^{\prime\prime}(0)\beta^{2t_{1}}+O\big(\beta^{3t_{1}}\big),$
where $t_{1}>0$ is to be determined. Thus
$h^{\prime\prime}(0)\geq-2\beta^{1-2t_{1}}+O\big(\beta^{t_{1}}\big).$
Taking $t_{1}=1/3$ yields
$h^{\prime\prime}(0)\geq-C\beta^{\frac{1}{3}},$ (2.3)
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$.
It holds that
$\begin{cases}0\leq h\big(\beta^{t_{2}}\big)=\beta+h^{\prime}(0)\beta^{t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{2t_{2}}+O\big(\beta^{3t_{2}}\big)\\ 0\leq h\big(-\beta^{t_{2}}\big)=\beta-h^{\prime}(0)\beta^{t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{2t_{2}}+O\big(\beta^{3t_{2}}\big),\end{cases}$
where $t_{2}>0$ is to be determined. Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\beta^{\frac{1}{2}-t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{t_{2}-\frac{1}{2}}+O\big(\beta^{2t_{2}-\frac{1}{2}}\big).$
Taking $t_{2}=1/3$ yields
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\frac{1}{2}h^{\prime\prime}(0)\beta^{-\frac{1}{6}}+C\beta^{\frac{1}{6}},$ (2.4)
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$.
Case 1: $h^{\prime\prime}(0)\leq K_{1}\beta^{1/3}$, where $K_{1}>0$ is to be determined. It follows from (2.4) that
$\frac{\left|h^{\prime}(0)\right|^{2}}{\beta}\leq\left(K_{1}/2+C\right)^{2}\beta^{\frac{1}{3}}.$
Combining this inequality and (2.3), we get
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq-\left(C+\alpha(K_{1}/2+C)^{2}\right)h^{\frac{1}{3}}(0)\quad\text{in Case 1}.$
Case 2: $h^{\prime\prime}(0)>K_{1}\beta^{1/3}$. It holds that
$\begin{cases}0\leq h\big(t_{3}\sqrt{\beta}\big)=\beta+h^{\prime}(0)t_{3}\sqrt{\beta}+\frac{1}{2}h^{\prime\prime}(0)\left(t_{3}\sqrt{\beta}\right)^{2}+O\big(\big(t_{3}\sqrt{\beta}\big)^{3}\big)\\ 0\leq h\big(-t_{3}\sqrt{\beta}\big)=\beta-h^{\prime}(0)t_{3}\sqrt{\beta}+\frac{1}{2}h^{\prime\prime}(0)\left(t_{3}\sqrt{\beta}\right)^{2}+O\big(\big(t_{3}\sqrt{\beta}\big)^{3}\big),\end{cases}$
where $t_{3}>0$ is to be determined. Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\frac{1}{t_{3}}+\frac{1}{2}h^{\prime\prime}(0)t_{3}+Ct_{3}^{2}\beta^{\frac{1}{2}},$
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$. Taking $t_{3}=\sqrt{2}\left(h^{\prime\prime}(0)\right)^{-1/2}$, which implies
$t_{3}\sqrt{\beta}=\sqrt{2}\left(h^{\prime\prime}(0)\right)^{-\frac{1}{2}}\sqrt{\beta}<\sqrt{2}K_{1}^{-\frac{1}{2}}\beta^{\frac{1}{3}},$
we have
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\sqrt{2}\sqrt{h^{\prime\prime}(0)}+2C\left(h^{\prime\prime}(0)\right)^{-\frac{3}{2}}\beta^{\frac{1}{2}}\sqrt{h^{\prime\prime}(0)}<\left(\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\right)\sqrt{h^{\prime\prime}(0)}.$
Given that $\alpha<1/2$, choose $K_{1}$ large such that
$\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\leq\frac{1}{\sqrt{\alpha}}.$
It follows that
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq 0\quad\text{in Case 2}.$
Combining Case 1 and Case 2 gives
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq-Kh^{\frac{1}{3}}(0),$
where $K$ depends on $\|h\|_{C^{2,1}(\mathbb{R})}$ and $\alpha$. By virtue of the above, (2.2) are valid at points in $\overline{\Omega}$ satisfying
$\max\left\{1,\sqrt{2}K_{1}^{-1/2}\right\}g^{\frac{1}{3}}(x)<\operatorname{dist}(\Omega,\partial\Omega_{0}).$
And (2.2) hold clearly at other points in $\overline{\Omega}$. The proof of Lemma 2.1 is complete. ∎
Remark 2.2.
In general, it is impossible to weaken the condition $g>0$ in $\Omega_{0}$ to $g>0$ in $\Omega$. A counterexample was provided by Błocki [1]: the function $g(x):=x$ on $[0,1]$ is smooth but $\sqrt{g}\notin C^{0,1}([0,1])$. However, the condition $g>0$ in $\Omega$ suffices to estimate the derivatives of $g$ in directions tangential to $\partial\Omega$.
Some fundamental properties of $\sigma_{k}$ are listed below, and one can refer to [27, 14, 26, 33].
Proposition 2.3.
For any $k\in\{1,2,\dots,n\}$ and $A\in\mathscr{S}^{n}$ with $\lambda(A)\in\Gamma_{k}$, the following hold for all $2\leq k\leq n$:
$\sigma_{k}^{\frac{1}{k}}[A]\leq C_{1}\sigma_{k-1}^{\frac{1}{k-1}}[A];$ (2.5)
$\sigma_{k-1}[A]\geq C_{2}\sigma_{1}^{\frac{1}{k-1}}[A]\sigma_{k}^{\frac{k-2}{k-1}}[A];$ (2.6)
$\sum_{i=1}^{n}\sigma_{k}^{ii}[A]=(n-k+1)\sigma_{k-1}[A],$ (2.7)
where $C_{1}>0$ and $C_{2}>0$ are universal.
The equation (1.1) is equivalent to
$F\big[D^{2}u\big]=g^{\frac{p}{k}}\quad\text{in}\ \Omega,$ (2.8)
where
$F[A]:=\sigma_{k}^{\frac{1}{k}}[A]\quad\text{and}\quad g:=f^{\frac{1}{p}},$
with
$p:=\begin{cases}k-1,\quad\text{if}\ f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}}),\\ (2k-2)/3,\quad\text{if}\ f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}}).\end{cases}$ (2.9)
Define $F^{ij}[A]:=\partial F/\partial A_{ij}$ and $F^{ij,st}[A]:=\partial^{2}F/\partial A_{ij}\partial A_{st}$. Differentiating equation (2.8) in the direction $e\in\mathbb{S}^{n-1}$ gives
$F^{ij}\big[D^{2}u\big]u_{eij}=\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}},\quad\forall e\in\mathbb{S}^{n-1}.$ (2.10)
A second differentiation combined with the concavity of $F$ yields
$\partial_{ee}\left(g^{\frac{p}{k}}\right)=F^{ij}\big[D^{2}u\big]u_{eeij}+F^{ij,st}\big[D^{2}u\big]u_{eij}u_{est}\leq F^{ij}\big[D^{2}u\big]u_{eeij},$
which implies
$F^{ij}\big[D^{2}u\big]u_{eeij}\geq\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{ee}g-\frac{k-p}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right),\quad\forall e\in\mathbb{S}^{n-1}.$ (2.11)
Define a matrix-valued function $G$ with entries
$G^{ij}(x):=\frac{F^{ij}\big[D^{2}u(x)\big]}{\operatorname{tr}F^{ij}\big[D^{2}u(x)\big]},\quad\forall x\in\Omega.$ (2.12)
Lemma 2.4.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), and $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$ hold. Then
$\operatorname{tr}F^{ij}\big[D^{2}u\big]\geq\frac{1}{C}\max\left\{1,\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}\right\};$ (2.13)
$0\leq\operatorname{tr}\left(GD^{2}u\right)\leq K;$ (2.14)
$\left|\operatorname{tr}\left(GD^{2}u_{\xi}\right)\right|\leq K\left|\xi\right|\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall\xi\in\mathbb{R}^{n},$ (2.15)
in $\Omega$, where $C>0$ is universal, and $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Step 1. It follows from (2.7) that
$\operatorname{tr}F^{ij}\big[D^{2}u\big]=\frac{1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sum_{i=1}^{n}\sigma_{k}^{ii}\big[D^{2}u\big]=\frac{n-k+1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sigma_{k-1}\big[D^{2}u\big].$
By (2.5) and (2.6), we get (2.13).
Step 2. It follows from (2.7) and the Euler theorem for homogeneous functions that
$\operatorname{tr}\left(GD^{2}u\right)=\frac{1}{n-k+1}\frac{\sigma_{k}^{ij}\big[D^{2}u\big]u_{ij}}{\sigma_{k-1}\big[D^{2}u\big]}=\frac{k}{n-k+1}\frac{\sigma_{k}\big[D^{2}u\big]}{\sigma_{k-1}\big[D^{2}u\big]}\geq 0.$
By (2.5) and (2.8), one derives
$\operatorname{tr}\left(GD^{2}u\right)\leq C\sigma_{k}^{\frac{1}{k}}\big[D^{2}u\big]=Cf^{\frac{1}{k}}\leq K,$
where $C$ is universal, and $K$ depends on $\|f\|_{C^{0}(\overline{\Omega})}$.
Step 3. The combination of (2.13) and (2.5) yields that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\left|F^{ij}\big[D^{2}u\big]u_{eij}\right|,\quad\forall e\in\mathbb{S}^{n-1}.$
Then it follows from (2.10), (2.9), and (2.1) that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\left|\nabla g\right|}{\sqrt{g}}\leq K\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall e\in\mathbb{S}^{n-1},$
where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Thus (2.15) is valid. ∎
Lemma 2.5.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), and $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Then
$\operatorname{tr}\left(GD^{2}u_{\xi\xi}\right)\geq-K\left|\xi\right|^{2}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall\xi\in\mathbb{R}^{n},$ (2.16)
in $\Omega$, where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
When $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it follows from (2.11), (2.13), and (2.1) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\frac{k-1}{k}g^{-\frac{1}{k}}\left|\partial_{ee}g-\frac{1}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right|\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
for all $e\in\mathbb{S}^{n-1}$, where $C$ is universal, and $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. When $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, which implies $(k+2)/3k<1/2$, it follows from (2.11) and (2.2) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)$ $\geq\frac{1}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}\frac{p}{k}g^{-\frac{2}{3k}-\frac{1}{3}}\left(\partial_{ee}g-\frac{k+2}{3k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)$
$\geq-\frac{1}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}K_{2}g^{-\frac{2}{3k}}$
for all $e\in\mathbb{S}^{n-1}$, where $K_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Combining this and (2.13) yields
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}K_{2}g^{-\frac{2}{3k}}=-CK_{2}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],$
where $C$ is universal. Therefore (2.16) holds. ∎
Below, we construct an auxiliary function which depends only on $\Omega$.
Lemma 2.6.
Let $\Omega\subset\mathbb{R}^{n}$ be a domain of class $C^{3,1}$ with $\partial\Omega$ strictly (k1)(k-1)-convex. Then there exists a function $\psi\in C^{3,1}(\overline{\Omega})$ satisfying
$\begin{cases}\operatorname{tr}\left(GD^{2}\psi\right)\leq-1&\quad\text{in}\ \Omega,\\ \psi>0&\quad\text{in}\ \Omega,\\ \psi=0&\quad\text{on}\ \partial\Omega,\\ \frac{D\psi(x)}{|D\psi(x)|}=\nu(x)&\quad\text{on}\ \partial\Omega,\\ |D\psi|\geq 1&\quad\text{on}\ \partial\Omega.\end{cases}$ (2.17)
Proof.
Let $\psi^{*}$ be the function constructed in Section 2 of Caffarelli-Nirenberg-Spruck [2], which satisfies
$\begin{cases}\lambda\big(D^{2}\psi^{*}(x)\big)\in\Gamma_{k},&\quad\forall x\in\overline{\Omega},\\ \psi^{*}<0&\quad\text{in}\ \Omega,\\ \psi^{*}=0&\quad\text{on}\ \partial\Omega,\\ \frac{D\psi^{*}(x)}{|D\psi^{*}(x)|}=-\nu(x)&\quad\text{on}\ \partial\Omega,\\ \left|D\psi^{*}\right|\geq\varepsilon_{0}&\quad\text{on}\ \partial\Omega,\end{cases}$
for some constant $\varepsilon_{0}>0$ depending on $\Omega$. By the compactness of $\overline{\Omega}$, there exists a small constant $\varepsilon_{1}>0$ such that
$\left\{\lambda\big(D^{2}\psi^{*}(x)\big):\ x\in\overline{\Omega}\right\}-\varepsilon_{1}\lambda(\mathbf{I}_{n})\Subset\Gamma_{k}.$
Using the concavity of $F$, we derive
$F^{ij}\big[D^{2}u\big]\left(K\big(\psi^{*}_{ij}-\varepsilon_{1}\delta_{ij}\big)-u_{ij}\right)$ $\geq F\big[K\big(D^{2}\psi^{*}-\varepsilon_{1}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=K\sigma_{k}^{\frac{1}{k}}\big[D^{2}\psi^{*}-\varepsilon_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0,$
where $K$ is sufficiently large. Consequently, by (2.14), we have
$F^{ij}\big[D^{2}u\big]\psi^{*}_{ij}\geq\varepsilon_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]+\frac{1}{K}F^{ij}\big[D^{2}u\big]u_{ij}\geq\varepsilon_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big].$
Thus
$\operatorname{tr}\left(GD^{2}\psi^{*}\right)\geq\varepsilon_{1}.$
Choosing $\psi=-t\psi^{*}$ with $t>0$ sufficiently large, we obtain (2.17). ∎
3. $C^{1}$ estimate and the boundary tangential-normal estimate
Following Section 2 of Caffarelli-Nirenberg-Spruck [2], one can construct a subsolution $\underline{u}$ to equation (1.1) satisfying
$\begin{cases}\sigma_{k}\big(\lambda\big(D^{2}\underline{u}\big)\big)\geq f\quad\text{in}\ \Omega,\\ \underline{u}=\varphi\quad\text{on}\ \partial\Omega.\end{cases}$
Let $h$ be the harmonic function in $\Omega$ which equals $\varphi$ on $\partial\Omega$. Using the maximum principle, we find
$\underline{u}\leq u\leq h\quad\text{in}\ \Omega.$
Combining this with $u=\varphi$ on $\partial\Omega$, we obtain
$\sup_{\overline{\Omega}}|u|+\sup_{\partial\Omega}|Du|\leq C,$ (3.1)
where $C$ depends on $\Omega$, $\|f\|_{C^{0}(\overline{\Omega})}$, and $\|\varphi\|_{C^{3}(\partial\Omega)}$. Recall that $g=f^{3/(2k-2)}$ and p=(2k2)/3p=(2k-2)/3.
Lemma 3.1.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\Omega}|Du|\leq\sup_{\partial\Omega}|Du|+C,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Without loss of generality, let $\Omega\subset\mathbb{R}^{n}\setminus B_{1}$. Suppose
$\sup_{x\in\overline{\Omega},e\in\mathbb{S}^{n-1}}\left(u_{e}+K|x|^{2}\right)$
is attained at an interior point $x_{0}\in\Omega$ in the direction $e=e_{1}$, where $K>0$ is a constant to be determined. Then at the point $x_{0}$,
$u_{1j}+2Kx_{j}=0,\quad\forall 1\leq j\leq n,$
and
$F^{ij}\big[D^{2}u\big]\left(u_{1ij}+2K\delta_{ij}\right)\leq 0.$ (3.2)
Since $\lambda(D^{2}u)\in\Gamma_{2}$, we derive at the point $x_{0}$,
$0\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{i,j=1}^{n}u_{ij}^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{j=1}^{n}u_{1j}^{2}=\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}|x|^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}.$ (3.3)
By (2.10) and (2.1), we obtain
$F^{ij}\big[D^{2}u\big]u_{1ij}=\frac{p}{k}\frac{\partial_{1}g}{\sqrt{g}}f^{\frac{1}{4(k-1)}-\frac{1}{k(k-1)}}\geq-K_{1}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$ (3.4)
where $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
It follows from (3.2), (3.4), (2.13), and (3.3) that at the point $x_{0}$,
$0$ $\geq 2K\operatorname{tr}F^{ij}\big[D^{2}u\big]-K_{1}f^{\frac{-1}{k(k-1)}}$
$\geq\frac{2K}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}-K_{1}f^{\frac{-1}{k(k-1)}}$
$\geq\left(\frac{2K}{C}(2K)^{\frac{1}{k-1}}-K_{1}\right)f^{\frac{-1}{k(k-1)}},$
where $C>0$ is universal. Taking $K$ large enough, we get a contradiction. The proof of Lemma 3.1 is complete. ∎
Lemma 3.2.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\partial\Omega}\left(|u_{\nu\eta}|+|u_{\eta\eta}|\right)\leq C,$
where $\eta$ is any unit tangential vector; and $C$ depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Let the origin be a boundary point satisfying $e_{n}=\nu(0)$, and $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)$ (3.5)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. It is clear that $\nabla_{x^{\prime}}\rho(0)=0$. Let $\varphi(0)=0$ and $\nabla\varphi(0)=0$ by subtracting an affine function. Differentiating $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ on $\partial\Omega$ twice, one gets
$u_{ij}(0)=\varphi_{ij}(0)-u_{n}(0)\rho_{ij}(0),\quad\forall 1\leq i,\ j\leq n-1.$
Thus the second-order tangential estimate is obtained.
For any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, define an operator
$T_{\alpha}:=\partial_{\alpha}+\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)(x_{\beta}\partial_{n}-x_{n}\partial_{\beta}),\quad\forall 1\leq\alpha\leq n-1.$
Since $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, we derive
$0=(u-\varphi)_{\alpha}+(u-\varphi)_{n}\rho_{\alpha}=(u-\varphi)_{\alpha}+(u-\varphi)_{n}\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)x_{\beta}+O\Big(\big|x^{\prime}\big|^{2}\Big)$
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for all $1\leq\alpha\leq n-1$. Hence
$T_{\alpha}(u-\varphi)=-x_{n}\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)(u-\varphi)_{\beta}+O\Big(\big|x^{\prime}\big|^{2}\Big)=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (3.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for all $1\leq\alpha\leq n-1$. It follows from (2.10) and (2.1) that
$\left|F^{ij}[D^{2}u]\left(T_{\alpha}(u-\varphi)\right)_{ij}\right|\leq\left|T_{\alpha}(g^{p})\right|+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\leq C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$ (3.7)
in $\Omega$, where $C_{1}$ depends on $\|\rho\|_{C^{3,1}}$, $\|\varphi\|_{C^{3}(\partial\Omega)}$; $C_{2}$ depends on $\|\rho\|_{C^{3,1}}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Define
$\omega_{r}:=\left\{x\in\Omega:\ \rho\big(x^{\prime}\big)<x_{n}<\rho\big(x^{\prime}\big)+r^{4},\ \big|x^{\prime}\big|<r\right\},\quad\forall r\leq r_{0}.$
Define a function
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{0}},$
where positive constants $\beta_{1}$ and $\beta_{2}$ are to be determined. It holds that
$D^{2}v=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]-\beta_{2}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big].$
Since $\partial\Omega\in C^{3,1}$ is strictly (k1)(k-1)-convex, one can fix $\beta_{1}$ sufficiently small such that $\lambda(D^{2}[x_{n}^{2}+\beta_{1}\rho](0))$ belongs to a compact subset of $\Gamma_{k}$ independent of $0\in\partial\Omega$. By
$|x|\rightarrow 0,\ \rho\big(x^{\prime}\big)\rightarrow 0,\ \nabla_{x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow 0,\ D^{2}_{x^{\prime}x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow D^{2}_{x^{\prime}x^{\prime}}\rho(0)\quad\text{as}\ r\rightarrow 0,$
there exist small positive constants $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\beta_{2}$, and $\delta_{1}$ such that
$\left\{\lambda\big(D^{2}v-\delta_{1}\mathbf{I}_{n}\big):\ x\in\omega_{r_{1}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big),$
and is thus contained in a compact subset of $\Gamma_{k}$. We emphasize that $\beta_{1}$, $r_{1}$, $\beta_{2}$, $\delta_{1}$ all depend only on $\Omega$. Using the concavity of $F$,
$F^{ij}\big[D^{2}u\big]\left(C_{3}(v_{ij}-\delta_{1}\delta_{ij})-u_{ij}\right)$ $\geq F\big[C_{3}\big(D^{2}v-\delta_{1}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=C_{3}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0$
in $\omega_{r_{1}}$, where $C_{3}$ is sufficiently large. Then by virtue of (2.14) and (2.13), we derive
$F^{ij}\big[D^{2}u\big]v_{ij}\geq\delta_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\geq\frac{\delta_{1}}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \omega_{r_{1}},$ (3.8)
where $C>0$ is universal.
We claim that
$\pm T_{\alpha}(u-\varphi)\leq-Kv\quad\text{in}\ \overline{\omega_{r_{1}}},\quad\forall 1\leq\alpha\leq n-1,$ (3.9)
where the constant $K>0$ is to be determined.
Step 1: verification of (3.9) on $\partial\omega_{r_{1}}$. The boundary $\partial\omega_{r_{1}}$ consists of three parts: $\partial_{1}\omega_{r_{1}}\cup\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$, where $\partial_{1}\omega_{r_{1}}$ and $\partial_{2}\omega_{r_{1}}$ are respectively the graph parts of $\rho$ and $\rho+r_{1}^{4}$, and $\partial_{3}\omega_{r_{1}}$ is the boundary part on $\{|x^{\prime}|=r_{1}\}$. On $\partial_{1}\omega_{r_{1}}$, using (3.6) one gets
$Kv=K\beta_{2}\big|x^{\prime}\big|^{2}\geq\pm T_{\alpha}(u-\varphi)$
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$. It is clear that
$\pm T_{\alpha}(u-\varphi)\leq C_{4},$
where $C_{4}$ depends on $\|\rho\|_{C^{2}}$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$. It holds that
$Kv\geq K\left(\beta_{1}r_{1}^{4}-r_{1}^{8}\right)\geq\frac{\beta_{1}}{2}Kr_{1}^{4}\quad\text{on}\ \partial_{2}\omega_{r_{1}},$
and
$Kv\geq K\beta_{2}r_{1}^{4}\quad\text{on}\ \partial_{3}\omega_{r_{1}}.$
Then we obtain $Kv\geq\pm T_{\alpha}(u-\varphi)$ on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$ by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$.
Step 2. Suppose, for contradiction, that (3.9) does not hold. Then the maximum
$\max_{\overline{\omega_{r_{1}}}}\left(\pm T_{\alpha}(u-\varphi)+Kv\right)$
is attained at an interior point $x_{0}\in\omega_{r_{1}}$. At the point $x_{0}$, we have
$0=\left(\pm T_{\alpha}(u-\varphi)+Kv\right)_{n}=\left(\pm T_{\alpha}u\right)_{n}-\left[-Kv_{n}+\left(\pm T_{\alpha}\varphi\right)_{n}\right]$ (3.10)
and
$F^{ij}\big[D^{2}u\big]\left(\pm T_{\alpha}(u-\varphi)+Kv\right)_{ij}\leq 0.$ (3.11)
It follows from (3.10) that
$\left|\left(T_{\alpha}u\right)_{n}\right|\geq-Kv_{n}-\left|\left(T_{\alpha}\varphi\right)_{n}\right|=K\left(\beta_{1}-2x_{n}+2\rho\big(x^{\prime}\big)\right)-\left|\left(T_{\alpha}\varphi\right)_{n}\right|\geq\frac{\beta_{1}}{2}K-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
at the point $x_{0}$. By taking $K$ large, which depends on $\Omega$ and $\|\varphi\|_{C^{2}(\partial\Omega)}$, we get
$\big|D^{2}u(x_{0})\big|\geq 1,$
which together with $\lambda(D^{2}u)\in\Gamma_{2}$ yields
$\sigma_{1}\big[D^{2}u(x_{0})\big]\geq\big|D^{2}u(x_{0})\big|\geq 1.$ (3.12)
Using (3.11), (3.7), (3.8), and (3.12), we obtain at the point $x_{0}$,
$\left(C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\right)+K\left(\frac{\delta_{1}}{2}\operatorname{tr}F^{ij}\big[D^{2}u\big]+\frac{\delta_{1}}{2C}f^{\frac{-1}{k(k-1)}}\right)\leq 0,$
which leads to a contradiction by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Therefore, claim (3.9) holds with $K$ depending on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Since $\nabla_{x^{\prime}}u(0)=\nabla_{x^{\prime}}\varphi(0)=0$, we have
$T_{\alpha}(u-\varphi)(0)=0,\quad\forall 1\leq\alpha\leq n-1.$
From (3.9), it follows that
$\pm\left(T_{\alpha}(u-\varphi)\right)_{n}(0)\leq-Kv_{n}(0)=K\beta_{1}.$
The proof of Lemma 3.2 is complete. ∎
Combining (3.1), Lemmas 3.1 and 3.2, we establish Theorem 1.1.
4. Reduction of the second-order estimate to the boundary
In this section, inspired by Guan [9], Guan-Trudinger-Wang [12], and Jiao-Wang [17], we reduce the estimate of second-order derivatives to the boundary second-order normal estimate, that is Theorem 1.2. Recall that $g=f^{3/(2k-2)}$ and p=(2k2)/3p=(2k-2)/3.
Proof of Theorem 1.2.
Define a function
$v:=\Delta u+|x|^{2}.$
Suppose $v$ attains its maximum at an interior point $x_{0}\in\Omega$, for otherwise we are done. Then at the point $x_{0}$, we have
$\partial_{x_{i}}(\Delta u)+2x_{i}=0,\quad\forall 1\leq i\leq n,$ (4.1)
and
$0\geq\sigma_{k}^{ij}\big[D^{2}u\big]v_{ij},$
which together with (2.7) and (2.6) implies
$0\geq\sigma_{k}^{ij}\big[D^{2}u\big](\Delta u)_{ij}+C\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]\sigma_{k}^{\frac{k-2}{k-1}}\big[D^{2}u\big],$ (4.2)
where $C>0$ is universal. Differentiating equation (1.1), we derive
$\sigma_{k}^{ij}\big[D^{2}u\big](\Delta u)_{ij}+\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}=\Delta(g^{p}).$
Combining this equality with (4.2), we obtain at the point $x_{0}$,
$0\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[(p-1)\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}.$ (4.3)
It follows from Lemma 3.2 of Guan-Li-Li [10] that computing
$0\geq\left(\left(\frac{\sigma_{k}}{\sigma_{1}}\right)^{1/(k-1)}\right)^{ij,st}u_{ije}u_{ste}$
yields
$\frac{\sigma_{k}^{ij,st}u_{ije}u_{ste}}{\sigma_{k}}\leq\frac{k-2}{k-1}\left(\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\right)^{2}+\frac{2}{k-1}\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}-\frac{k}{k-1}\left(\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}\right)^{2}$
for all $e\in\mathbb{S}^{n-1}$, where operators take values at arbitrary matrix $A$ with $\lambda(A)\in\Gamma_{k}$. Thus by (4.1), we infer at the point $x_{0}$,
$-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}$ $\geq-\sum_{l=1}^{n}\frac{k-2}{k-1}\frac{p^{2}g^{p-1}|\partial_{x_{l}}g|^{2}}{g}$ (4.4)
$\hskip 12.80373pt-\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{pg^{p-1}\partial_{x_{l}}g2x_{l}}{\Delta u}-\frac{kf}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right].$
Using (4.3), (4.4), (2.2), and (2.1), we get at the point $x_{0}$,
$0$ $\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[-\frac{1}{3}\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]$
$\hskip 12.80373pt-\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{pg^{p-1}\partial_{x_{l}}g2x_{l}}{\Delta u}-\frac{kf}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right]$
$\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}-K_{1}f^{1-\frac{1}{k-1}}$
$\hskip 12.80373pt-f^{1-\frac{1}{k-1}}\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{K_{2}f^{\frac{1}{4(k-1)}}2|x_{l}|}{\Delta u}-\frac{kf^{\frac{1}{k-1}}}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right],$
where $K_{1}$ and $K_{2}$ depend on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Therefore,
$\Delta u(x_{0})\leq K,$
where $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. The proof of Theorem 1.2 is complete. ∎
5. The weakly interior estimate
We recall the following comparison principle established by Ivochkina-Trudinger-Wang [16], and provide a proof for completeness.
Lemma 5.1 (Lemma 3.1 of [16]).
Let $w(x,y)$, $v(x,y)$ be functions in $\overline{\Omega}\times\mathbb{R}^{N}$ that are both $p$-homogeneous in $y\in\mathbb{R}^{N}$ for some $p\in\mathbb{R}$. Let $w$, $v\in C^{2}(\Omega\times\mathbb{S}^{N-1})\cap C^{0}(\overline{\Omega}\times\mathbb{S}^{N-1})$ with $v>0$ in $\overline{\Omega}\times\mathbb{S}^{N-1}$. Define the linear operator $\mathscr{L}$ by
$\mathscr{L}(w):=\operatorname{tr}\big(\widetilde{G}D^{2}w\big)+\langle b,Dw\rangle,$
where $\widetilde{G}=\widetilde{G}(x,y)$ is an $(n+N)\times(n+N)$ positive semi-definite matrix and $b=b(x,y)$ is an $(n+N)$-dimensional vector. If there exist two positive constants $\mu_{1}$, $\mu_{2}$ such that
$\begin{cases}\mathscr{L}(w)\geq-\mu_{1}\quad\text{in}\ \Omega\times\mathbb{S}^{N-1},\\ \mathscr{L}(v)\leq-\mu_{2}\quad\text{in}\ \Omega\times\mathbb{S}^{N-1},\end{cases}$ (5.1)
then
$\sup_{\overline{\Omega}\times\mathbb{S}^{N-1}}\frac{w}{v}\leq\frac{\mu_{1}}{\mu_{2}}+\sup_{\partial\Omega\times\mathbb{S}^{N-1}}\frac{w}{v}.$
Proof.
Since $w$, $v$ are $p$-homogeneous in $y\in\mathbb{R}^{N}$, we have
$\frac{w}{v}(x,y)\in C^{2}(Q)\cap C^{0}(\overline{Q})\quad\text{and}\quad\frac{w}{v}(x_{0},y_{0})=\sup_{\overline{Q}}\frac{w}{v},$
where $Q:=\Omega\times(B_{2}\setminus B_{1/2})$ and $(x_{0},y_{0})\in\overline{\Omega}\times\mathbb{S}^{N-1}$. We only need to consider the case of $(x_{0},y_{0})\in\Omega\times\mathbb{S}^{N-1}$ with $w(x_{0},y_{0})>0$. At the maximum point $(x_{0},y_{0})$,
$\partial_{i}\Big(\frac{w}{v}\Big)=\frac{\partial_{i}w}{v}-\frac{w\partial_{i}v}{v^{2}}=0,\quad\forall 1\leq i\leq n+N,$ (5.2)
and the Hessian
$\partial_{ij}\Big(\frac{w}{v}\Big)=\frac{\partial_{ij}w}{v}-\frac{\partial_{i}w\partial_{j}v+\partial_{j}w\partial_{i}v}{v^{2}}+\frac{2w\partial_{i}v\partial_{j}v}{v^{3}}-\frac{w\partial_{ij}v}{v^{2}}$
is an $(n+N)\times(n+N)$ negative semi-definite matrix. Then at the point $(x_{0},y_{0})$,
$0\geq\widetilde{G}^{ij}\partial_{ij}\Big(\frac{w}{v}\Big)$ $=\widetilde{G}^{ij}\left(\frac{\partial_{ij}w}{v}-\frac{w\partial_{ij}v}{v^{2}}+\left(\frac{2w\partial_{i}v}{v^{3}}-\frac{2\partial_{i}w}{v^{2}}\right)\partial_{j}v\right)$
[…]
$=\frac{\mathscr{L}(w)-b^{i}\partial_{i}w}{v}-\frac{w}{v^{2}}\left(\mathscr{L}(v)-b^{i}\partial_{i}v\right)$
[…]
$\overset{w,v>0}{\geq}\frac{1}{v}\left(-\mu_{1}+\frac{w}{v}\mu_{2}\right).$
This yields
$\frac{w}{v}(x_{0},y_{0})\leq\frac{\mu_{1}}{\mu_{2}}.$
The proof of Lemma 5.1 is complete. ∎
We now proceed to prove the weakly interior estimate.
Proof of Theorem 1.5.
We will apply Lemma 5.1 with $N=n+1$ and coordinates $(x,\xi,\eta)\in\overline{\Omega}\times\mathbb{R}^{n}\times\mathbb{R}$. Define functions
$w:=u_{\xi\xi}+2\eta u_{\xi}+\eta^{2}u$
and
$v:=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\varrho(\psi+\beta)\eta^{2},$
where $0<\alpha<1$, $0<\varrho<1$, and $0<\beta\ll 1$ are parameters to be determined; and the function $\psi$ is given in Lemma 2.6. Define a linear operator
$\mathscr{L}(w):=\operatorname{tr}\big(\widetilde{G}D^{2}w\big)+\langle b,Dw\rangle,$
where
$\widetilde{G}:=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix},\quad b:=\begin{pmatrix}0_{n\times 1}\\ -2Gq\\ 0_{1\times 1}\end{pmatrix}$,
with $G$ being the $n\times n$ matrix-valued function defined in (2.12). The scalar function $r$ and the $n$-dimensional vector-valued function $q$ are defined by
$r:=\frac{\alpha\psi_{\xi}}{\psi+\beta},\quad q:=\frac{1}{\psi+\beta}\left(\frac{\xi}{4}+\frac{\alpha\psi_{\xi}}{\psi+\beta}\nabla\psi\right).$ (5.3)
It holds that the quadratic form
$\left(\xi_{1},\xi_{2},\eta\right)^{T}\widetilde{G}\left(\xi_{1},\xi_{2},\eta\right)=\left(\xi_{1}+r\xi_{2}+\eta q\right)^{T}G\left(\xi_{1}+r\xi_{2}+\eta q\right)\geq 0$
is valid for all $(\xi_{1},\xi_{2},\eta)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$, which implies that $\widetilde{G}$ is positive semi-definite. Due to the substantial computations involved, we divide the proof into five steps for clarity.
Step 1. For convenience, denote
$\hat{v}:=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}.$
We claim that
$\mathscr{L}(\hat{v})\leq-\frac{1}{C_{\psi}}\frac{1}{8\alpha}|\xi|^{2}$ (5.4)
in $\Omega\times\mathbb{R}^{n}\times\mathbb{R}$ for all $\alpha\in(0,\alpha_{0})$, where constants $C_{\psi}$ and $\alpha_{0}$ depend on $\|\psi\|_{C^{3}(\overline{\Omega})}$ to be determined. It holds that
$\widetilde{G}D^{2}\hat{v}=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix}\begin{pmatrix}\partial_{x_{i}x_{j}}&\partial_{x_{i}\xi_{j}}&0\\ \partial_{\xi_{i}x_{j}}&\partial_{\xi_{i}\xi_{j}}&0\\ 0&0&0\end{pmatrix}[\hat{v}$],
where $1\leq i,\ j\leq n$. Since $G$ is symmetric, we derive
$\mathscr{L}(\hat{v})=G^{ij}\partial_{x_{i}x_{j}}\hat{v}+2rG^{ij}\partial_{x_{i}\xi_{j}}\hat{v}+r^{2}G^{ij}\partial_{\xi_{i}\xi_{j}}\hat{v}-2G^{ij}q_{j}\partial_{\xi_{i}}\hat{v}$. (5.5)
Compute derivatives
$\partial_{x_{i}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\left(\frac{1-\alpha}{4\alpha}|\xi|^{2}\psi_{x_{i}}+2\psi_{\xi}\psi_{\xi x_{i}}-\alpha\frac{\psi_{\xi}^{2}\psi_{x_{i}}}{\psi+\beta}\right);$
$\partial_{x_{i}x_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\left(\frac{1-\alpha}{4\alpha}|\xi|^{2}\psi_{x_{i}x_{j}}+2\psi_{\xi x_{i}}\psi_{\xi x_{j}}+2\psi_{\xi}\psi_{\xi x_{i}x_{j}}\right)$ (5.6)
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{x_{i}}\psi_{\xi x_{j}}-\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$r\partial_{x_{i}\xi_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{1-\alpha}{2}\psi_{\xi}\psi_{x_{i}}\xi_{j}+2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}+2\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$ (5.7)
$\hskip 12.80373pt-\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$\partial_{\xi_{i}}\hat{v}$ $=\frac{(\psi+\beta)^{1-\alpha}}{2\alpha}\xi_{i}+(\psi+\beta)^{-\alpha}2\psi_{\xi}\psi_{x_{i}};$
$q_{j}\partial_{\xi_{i}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\frac{\xi_{i}\xi_{j}}{8\alpha}+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{1}{2}\psi_{\xi}\xi_{i}\psi_{x_{j}}+\frac{1}{2}\psi_{\xi}\psi_{x_{i}}\xi_{j}\right)$ (5.8)
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$r^{2}\partial_{\xi_{i}\xi_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha+1}}\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}+\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}.$ (5.9)
Combining (5.5)–(5.9) with the symmetry of $G$ yields that
[…]
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}\psi_{x_{i}}\xi_{j}+3\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}+\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}\right)$
$\hskip 12.80373pt-\frac{G^{ij}}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+3\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}.$
Since $\operatorname{tr}\left(GD^{2}\psi\right)\leq-1$ from (2.17), $\operatorname{tr}G=1$, and $G$ is positive definite, we get that
[…]
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}G^{ij}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}G^{ij}\psi_{x_{i}}\xi_{j}-\frac{5}{2}\alpha\psi_{\xi}^{2}\right)$
$\hskip 12.80373pt-\frac{G^{ij}}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+3\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}$
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{1}\|\psi\|_{C^{3}(\overline{\Omega})}^{2}\right)|\xi|^{2}$
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha}}\left[-\frac{1}{4\alpha}\xi_{i}\xi_{j}-\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}-\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right],$ (5.10)
where the constant $C_{1}>0$ is universal. It holds that
$0$ $\leq\sum_{s=1}^{n}\left[\sum_{i=1}^{n}\Big(G^{1/2}\Big)^{si}\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{i}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{i}\right)\right]^{2}$
$=\sum_{i,j=1}^{n}G^{ij}\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{i}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{i}\right)\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{j}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{j}\right)$
$=\sum_{i,j=1}^{n}G^{ij}\left[\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}+\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}+\frac{(1+\alpha)^{2}}{4\alpha(3+\alpha)}\xi_{i}\xi_{j}\right].$ (5.11)
By (5.10) and (5.11), we obtain that
[…]
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{\psi}\right)|\xi|^{2}$
$\leq-\frac{1}{\psi+1}\frac{1}{8\alpha}|\xi|^{2}$
for all $\alpha\in(0,\alpha_{0})$ with $\alpha_{0}>0$ chosen sufficiently small, where constants $C_{\psi}$ and $\alpha_{0}$ depend on $\|\psi\|_{C^{3}(\overline{\Omega})}$. This establishes claim (5.4).
Step 2. It holds that
$\widetilde{G}D^{2}\left[(\psi+\beta)\eta^{2}\right]=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix}\begin{pmatrix}\partial_{x_{i}x_{j}}&0&\partial_{x_{i}\eta}\\ 0&0&0\\ \partial_{\eta x_{j}}&0&\partial_{\eta\eta}\end{pmatrix}\left[(\psi+\beta)\eta^{2}\right],$
where $1\leq i,\ j\leq n$. By the symmetry of $G$, we compute
$\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$ $=G^{ij}\left(\partial_{x_{i}x_{j}}+2q_{j}\partial_{\eta x_{i}}+q_{i}q_{j}\partial_{\eta\eta}\right)\left[(\psi+\beta)\eta^{2}\right]$
$=\eta^{2}G^{ij}\psi_{x_{i}x_{j}}+G^{ij}\left(\frac{1}{\psi+\beta}\eta\psi_{x_{i}}\xi_{j}+\frac{1}{(\psi+\beta)^{2}}4\alpha\eta\psi_{\xi}\psi_{x_{i}}\psi_{x_{j}}\right)$
$\hskip 12.80373pt+G^{ij}\left(\frac{1}{\psi+\beta}\frac{1}{8}\xi_{i}\xi_{j}+\frac{1}{(\psi+\beta)^{2}}\alpha\psi_{\xi}\psi_{x_{i}}\xi_{j}+\frac{1}{(\psi+\beta)^{3}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right).$
Since $\operatorname{tr}\left(GD^{2}\psi\right)\leq-1$ from (2.17), $\operatorname{tr}G=1$, and $G$ is positive definite, we get that
$\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$ $\leq-\eta^{2}+\frac{1}{\beta^{2}}C_{\psi}|\eta||\xi|+\frac{1}{\beta^{3}}C_{\psi}|\xi|^{2}$
$\leq-\frac{1}{2}\eta^{2}+\frac{1}{2\beta^{4}}C_{\psi}^{2}|\xi|^{2}+\frac{1}{\beta^{3}}C_{\psi}|\xi|^{2}$
$\leq-\frac{1}{2}\eta^{2}+\frac{1}{\beta^{4}}C_{\psi}^{2}|\xi|^{2},$ (5.12)
where the constant $C_{\psi}>0$ depends on $\|\psi\|_{C^{3}(\overline{\Omega})}$. Then it follows from (5.4) and (5.12) that
$\mathscr{L}(v)$ $\leq-\frac{1}{C_{\psi}}\frac{1}{8\alpha}|\xi|^{2}+\varrho\left(-\frac{1}{2}\eta^{2}+\frac{1}{\beta^{4}}C_{\psi}^{2}|\xi|^{2}\right)$
$=\left(-\frac{1}{8C_{\psi}}\frac{1}{\alpha}+C_{\psi}^{2}\right)|\xi|^{2}-\frac{1}{2}\beta^{4}\eta^{2}$
$\leq-\frac{1}{2}\beta^{4}\left(|\xi|^{2}+\eta^{2}\right)$ (5.13)
in $\Omega\times\mathbb{R}^{n}\times\mathbb{R}$ for all $\alpha\in(0,\alpha_{0})$, where we set
$\varrho:=\beta^{4}$
and $\alpha_{0}>0$ chosen sufficiently small depending on $\|\psi\|_{C^{3}(\overline{\Omega})}$. Fix $\alpha\in(0,\alpha_{0})$ depending on $\|\psi\|_{C^{3}(\overline{\Omega})}$ in subsequent arguments.
Step 3. It holds that
[…]
where $1\leq i,\ j\leq n$. By the symmetry of $G$, we compute
$\mathscr{L}(w)$ $=G^{ij}\left(\partial_{x_{i}x_{j}}+2r\partial_{x_{i}\xi_{j}}+2q_{j}\partial_{x_{i}\eta}+r^{2}\partial_{\xi_{i}\xi_{j}}+2rq_{j}\partial_{\xi_{i}\eta}+q_{i}q_{j}\partial_{\eta\eta}-2q_{j}\partial_{\xi_{i}}\right)[w]$
$=G^{ij}2q_{j}(\partial_{x_{i}\eta}w-\partial_{\xi_{i}}w)+G^{ij}\big(u_{\xi\xi x_{i}x_{j}}+2\eta u_{\xi x_{i}x_{j}}+\eta^{2}u_{x_{i}x_{j}}\big)$
$\hskip 12.80373pt+G^{ij}2r\big(2u_{\xi x_{i}x_{j}}+2\eta u_{x_{i}x_{j}}\big)+G^{ij}r^{2}2u_{x_{i}x_{j}}+G^{ij}2rq_{j}2u_{x_{i}}+G^{ij}q_{i}q_{j}2u$
$=G^{ij}\big(u_{\xi\xi x_{i}x_{j}}+2\eta u_{\xi x_{i}x_{j}}+\eta^{2}u_{x_{i}x_{j}}\big)+4rG^{ij}\big(u_{\xi x_{i}x_{j}}+\eta u_{x_{i}x_{j}}\big)$
$\hskip 12.80373pt+2r^{2}G^{ij}u_{x_{i}x_{j}}+4rG^{ij}u_{x_{i}}q_{j}+2uG^{ij}q_{i}q_{j}.$
It follows from (2.13)–(2.16), (5.3), and $0\leq G\leq\mathbf{I}_{n}$ that
$\mathscr{L}(w)$ $\geq-K_{1}|\xi|^{2}\sigma_{1}^{\frac{-1}{k-1}}-2|\eta|K_{1}|\xi|\sigma_{1}^{\frac{-1}{k-1}}-4|r|K_{1}|\xi|\sigma_{1}^{\frac{-1}{k-1}}-4|r||\eta|K_{1}$
$\hskip 12.80373pt-4|r|\|u\|_{C^{1}(\overline{\Omega})}|q|-2\|u\|_{C^{0}(\overline{\Omega})}|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+r^{2}\right)-K_{1}\left(4\eta^{2}+r^{2}\right)-4r^{2}-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-K_{1}\left(4\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-4\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}$
$\hskip 12.80373pt-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)\frac{C_{\psi}}{\beta^{4}}|\xi|^{2}$
$\geq-K\left(1+\sigma_{1}^{\frac{-1}{k-1}}\right)\frac{1}{\beta^{4}}\left(|\xi|^{2}+\eta^{2}\right),$ (5.14)
where $K>0$ depends on $\|\psi\|_{C^{1}(\overline{\Omega})}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Step 4. It holds that
$w$ $=\left(\theta\nu+\tau\sqrt{|\xi|^{2}-\theta^{2}}\right)^{T}D^{2}u\left(\theta\nu+\tau\sqrt{|\xi|^{2}-\theta^{2}}\right)+2\eta u_{\xi}+\eta^{2}u$
$\leq\theta^{2}\sup_{\partial\Omega}u_{\nu\nu}+C$ (5.15)
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$, where $\tau$ is a unit tangential direction, $\theta:=\langle\xi,\nu\rangle$, and the constant $C>0$ depends on $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Moreover, by (2.17) we get that
$v$ $=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\beta^{4}(\psi+\beta)\eta^{2}$
$=\frac{\beta^{1-\alpha}}{4\alpha}|\xi|^{2}+\frac{1}{\beta^{\alpha}}(\left|\nabla\psi\right|\langle\nu,\xi\rangle)^{2}+\beta^{5}\eta^{2}$
$\geq\frac{1}{\beta^{\alpha}}\theta^{2}+\beta^{5}\big(|\xi|^{2}+\eta^{2}\big)$
$=\frac{1}{\beta^{\alpha}}\theta^{2}+\beta^{5}$ (5.16)
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$. Therefore, it follows from (5.15) and (5.16) that
$\frac{w}{v}\leq\frac{\theta^{2}\sup_{\partial\Omega}u_{\nu\nu}+C}{\theta^{2}\beta^{-\alpha}+\beta^{5}}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{C}{\beta^{5}}$
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$. Thus
$\sup_{\partial\Omega\times\mathbb{S}^{n}}\frac{w}{v}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{C}{\beta^{5}}.$ (5.17)
Step 5. Let
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})=\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v},$
where $(x_{*},\xi_{*},\eta_{*})\in\overline{\Omega}\times\mathbb{S}^{n}$. Applying Lemma 5.1 with (5.13), (5.14), and (5.17), we derive
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{2K+C}{\beta^{8}}\left(1+\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u(x_{*})\big]\right).$ (5.18)
Suppose
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})>\frac{1}{\beta^{8}},$
then we have at the point $(x_{*},\xi_{*},\eta_{*})$,
$u_{\xi\xi}+2\eta u_{\xi}+\eta^{2}u$ $\geq\frac{1}{\beta^{8}}\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\beta^{4}(\psi+\beta)\eta^{2}\right)$
$\geq\frac{1}{\beta^{8}}\left(\frac{1}{4}\beta|\xi|^{2}+\beta^{5}\eta^{2}\right)$
$\geq\frac{1}{\beta^{3}}\left(|\xi|^{2}+\eta^{2}\right)$
$=\frac{1}{\beta^{3}}.$
Since $\lambda(D^{2}u(x_{*}))\in\Gamma_{2}$, one gets at the point $x_{*}$,
$\sigma_{1}\big[\lambda\big(D^{2}u\big)\big]\geq\lambda_{\operatorname{max}}\big(D^{2}u\big)\geq u_{\xi_{*}\xi_{*}}\geq\frac{1}{\beta^{3}}-2\eta_{*}u_{\xi_{*}}-\eta_{*}^{2}u\geq\frac{1}{\beta^{3}}-3\|u\|_{C^{1}(\overline{\Omega})}>1,$ (5.19)
where $0<\beta\ll 1$. Combining (5.18) and (5.19) yields
$\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}.$
Therefore, for any $x\in\Omega$ and $\xi\in\mathbb{S}^{n-1}$,
$w(x,\xi,0)$ $\leq v(x,\xi,0)\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v}$
$=\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}\right)\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right)$ (5.20)
$\leq\frac{C_{\psi}}{\psi^{\alpha}}\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right),$ (5.21)
where the constant $C_{\psi}>0$ depends on $\|\psi\|_{C^{3}(\overline{\Omega})}$. From the construction of $\psi^{*}=-t^{-1}\psi$ in Lemma 2.6, there exist small positive constants $\delta$ and $\delta^{*}$ such that
$\psi^{*}\leq-\delta^{*}d(x)\quad\text{in}\ \Omega_{\delta},$ (5.22)
and since $\psi^{*}$ is subharmonic,
$\psi^{*}\leq-\delta^{*}\delta\quad\text{in}\ \Omega\setminus\Omega_{\delta},$ (5.23)
where $t$, $\delta$, $\delta^{*}$ depend on $\Omega$. It follows from (5.21)–(5.23) that
$u_{\xi\xi}(x)\leq C_{*}\frac{1}{d(x)}\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{1}{\beta^{8}}\right),\quad\forall 0<\beta\ll 1,$
for all $x\in\Omega$ and $\xi\in\mathbb{S}^{n-1}$, where the constant $C_{*}>0$ depends on $\Omega$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Finally, taking
$\varepsilon:=C_{*}\beta^{\alpha}\quad\text{and}\quad C_{\varepsilon}:=C_{*}\beta^{-8},$
using Theorem 1.1, we complete the proof of Theorem 1.5. ∎
6. The boundary estimate in terms of the interior one
Let $A=(A_{ij})$ be a skew-symmetric matrix, which implies that the matrix $\exp{A}$ is orthogonal. Define a vector field $\bar{\tau}:\ \overline{\Omega}\rightarrow\mathbb{R}^{n}$ by
$\bar{\tau}(x):=Ax.$
We adopt notations
$u_{\bar{\tau}\bar{\tau}}:=u_{ij}\bar{\tau}_{i}\bar{\tau}_{j}\quad\text{and}\quad u_{(\bar{\tau})(\bar{\tau})}:=\big(u_{\bar{\tau}}\big)_{\bar{\tau}}=u_{ij}\bar{\tau}_{i}\bar{\tau}_{j}+u_{i}\bar{\tau}_{j}\partial_{x_{j}}(\bar{\tau}_{i}).$
The following derivative exchange formula is due to Ivochkina-Trudinger-Wang [16].
Lemma 6.1 (Lemma 2.1 of [16]).
For any function $u\in C^{4}(\overline{\Omega})$, it holds that
$F^{ij}\big[D^{2}u\big]\left(u_{(\tau)(\tau)}\right)_{ij}=\left(F\big[D^{2}u\big]\right)_{(\tau)(\tau)}-F^{ij,st}\big[D^{2}u\big](u_{\tau})_{ij}(u_{\tau})_{st},$
where $\tau(x):=\vec{a}+\bar{\tau}(x)$ for any fixed vector $\vec{a}\in\mathbb{R}^{n}$.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution to the Dirichlet problem (1.1). Let the origin be a boundary point satisfying $\nu(0)=e_{n}$, and $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)$ (6.1)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. It is clear that $\nabla_{x^{\prime}}\rho(0)=0$ and $x_{n}=O(|x^{\prime}|^{2})$. Let $\varphi(0)=0$ and $\nabla\varphi(0)=0$ by subtracting an affine function. It follows from (6.1) that
$\nu(x)=\frac{\left(-\rho_{1}(x^{\prime}),\dots,-\rho_{n-1}(x^{\prime}),1\right)}{\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}}$ (6.2)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. We introduce a tangential vector field $\eta$ defined by
$\eta(x):=\left(\frac{x^{\prime}}{|x^{\prime}|},\sum_{i=1}^{n-1}\frac{x_{i}}{|x^{\prime}|}\rho_{i}\big(x^{\prime}\big)\right)\quad\text{in}\ B_{r_{0}}\setminus\{0\}.$ (6.3)
We construct a vector field
$\tau=\tau(x):=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}+\begin{pmatrix}~&~&~&-\rho_{11}(0)\\ ~&0_{(n-1)\times(n-1)}\hskip-28.45274pt&~&\vdots\\ ~&~&~&-\rho_{1,n-1}(0)\\ \rho_{11}(0)\hskip-19.91692pt&\cdots\hskip-17.07164pt&\rho_{1,n-1}(0)\hskip-8.53581pt&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\end{pmatrix}=\begin{pmatrix}1-\rho_{11}(0)x_{n}\\ -\rho_{12}(0)x_{n}\\ \vdots\\ -\rho_{1,n-1}(0)x_{n}\\ \sum_{i=1}^{n-1}\rho_{1i}(0)x_{i}\end{pmatrix}$ (6.4)
in $\overline{\Omega}$. Define
$w(x):=u_{(\tau)(\tau)}(x)-u_{(\tau)(\tau)}(0)\quad\text{in}\ \overline{\Omega},$
and
$M:=\sup_{x\in\partial\Omega}u_{\nu\nu}(x).$
Without loss of generality, we assume $M\gg 1$, and $C_{\varepsilon}>1$ for all $\varepsilon>0$ in (1.8).
Lemma 6.2.
There exist a positive constant $K_{0}$ and a linear function $h$ of $x^{\prime}$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)$ (6.5)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but are independent of $\inf_{\Omega_{0}}f$.
Proof.
For any point $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, let $\xi(x)$ be the projection of $\tau(x)$ onto the tangent plane of $\partial\Omega$ at $x$. By (6.2) and (6.4), one has
$\langle\tau,\nu\rangle=\left[\left(-\rho_{1}\big(x^{\prime}\big)+\sum_{i=1}^{n-1}\rho_{1i}(0)x_{i}\right)+x_{n}\sum_{i=1}^{n-1}\rho_{i}\big(x^{\prime}\big)\rho_{1i}(0)\right]\frac{1}{\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}}}=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. It is clear that for any $i\in\{1,\dots,n\}$,
$u_{i}(x)-u_{i}(0)$ $=u_{i}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)-u_{i}(0,\rho(0))$
$=\nabla_{x^{\prime}}\partial_{x_{i}}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\cdot x^{\prime}+\partial_{x_{n}x_{i}}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\left(\nabla_{x^{\prime}}\rho\big(tx^{\prime}\big)\cdot x^{\prime}\right)$
$=e_{i}^{T}D^{2}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\eta\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\times\left|x^{\prime}\right|$
$=O\big(\big|x^{\prime}\big|\big),$ (6.7)
for all $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $t\in(0,1)$ depends on $x$, and the tangential vector field $\eta$ is defined in (6.3). Since $\nabla_{x^{\prime}}u(0)=\nabla_{x^{\prime}}\varphi(0)=0$, using (6.7) one gets
$|u_{\nu}(x)-u_{n}(x)|$ $=|\nabla u(x)\cdot(\nu(x)-e_{n})|$
$\leq\sum_{i=1}^{n-1}\left|\left(u_{i}(x)-u_{i}(0)\right)\rho_{i}\big(x^{\prime}\big)\right|+|u_{n}(x)|\frac{\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}{1+\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}}$
$=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.8)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
Combining (6.6), (6.4), and (6.7), one derives
$u_{(\tau)(\tau)}(x)$ $=(\xi+\langle\tau,\nu\rangle\nu)^{T}D^{2}u\left(\xi+\langle\tau,\nu\rangle\nu\right)+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=u_{\xi\xi}+O\Big(M\big|x^{\prime}\big|^{4}\Big)+O\Big(\big|x^{\prime}\big|^{2}\Big)+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=u_{\xi\xi}+u_{i}\partial_{x_{1}}(\tau_{i})+u_{i}\tau_{n}\partial_{x_{n}}(\tau_{i})+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}-\sum_{i=1}^{n-1}\bigg(\rho_{1i}(0)\sum_{k=1}^{n-1}\rho_{1k}(0)u_{k}(x)\bigg)x_{i}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}(x)+h_{1}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$ (6.9)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where
$h_{1}\big(x^{\prime}\big):=-\sum_{i=1}^{n-1}\bigg(\rho_{1i}(0)\sum_{k=1}^{n-1}\rho_{1k}(0)u_{k}(0)\bigg)x_{i}.$
Using $\partial\Omega\in C^{3,1}$ and $u=\varphi$ on $\partial\Omega$, one gets
$u_{\xi\xi}(x)$ $=\varphi_{\xi\xi}(x)+\kappa(x,\xi(x))\left(\varphi_{\nu}(x)-u_{\nu}(x)\right)$
$=\left[\varphi_{11}(0)+\rho_{11}(0)\varphi_{n}(0)\right]+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$
$=\left[u_{11}(0)+\rho_{11}(0)u_{n}(0)\right]+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$ (6.10)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $\kappa(x,\xi)$ is the curvature of $\partial\Omega$ along the direction $\xi$ at $x\in\partial\Omega$, and $h_{2}(x^{\prime})$ is the linear part in the Taylor expansion of $\varphi_{\xi\xi}(x)+\kappa(x,\xi(x))\varphi_{\nu}(x)$ with respect to $x^{\prime}$. By (6.9), (6), (6.8), and (6.7), we have
$w(x)$ $=u_{(\tau)(\tau)}(x)-\left[u_{11}(0)+\rho_{11}(0)u_{n}(0)\right]$
$=\rho_{11}(0)u_{n}(x)-\kappa(x,\xi(x))u_{\nu}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(0)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big)$
$=h_{3}\big(x^{\prime}\big)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}\Big)+O\Big(M\big|x^{\prime}\big|^{4}\Big),$
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h_{3}(x^{\prime})$ is the linear part in the Taylor expansion of $\kappa(x,\xi(x))u_{n}(0)$ with respect to $x^{\prime}$.
Consequently, there exists a constant $K_{0}>0$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)$
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h(x^{\prime}):=h_{1}(x^{\prime})+h_{2}(x^{\prime})+h_{3}(x^{\prime})$; $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Using Theorem 1.1, we complete the proof of Lemma 6.2. ∎
Define the linear elliptic operator $\mathcal{L}$ by
$\mathcal{L}(w):=F^{ij}\big[D^{2}u\big]w_{ij}.$
Define
$\omega_{r}:=\left\{x\in\Omega:\ \rho\big(x^{\prime}\big)<x_{n}<\rho\big(x^{\prime}\big)+r^{4},\ \big|x^{\prime}\big|<r\right\},\quad\forall r\leq r_{0}.$
6.1. The best possible condition
In this subsection, we assume either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ holds or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Additionally, we assume that the condition (1.5) is satisfied. Recall that $g=f^{1/p}$, where $p$ is given by (2.9).
Lemma 6.3.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. We have two conclusions:
$(i)$ If $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, then
$\mathcal{L}(w)\geq-C_{1}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$
where $C_{1}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
$(ii)$ If $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, then
$\mathcal{L}(w)\geq-C_{2}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$
where $C_{2}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
From Lemma 6.1, we have
$\mathcal{L}(w)$ $=F^{ij}\big[D^{2}u\big]\left(u_{(\tau)(\tau)}\right)_{ij}$
$\geq\left(g^{\frac{p}{k}}\right)_{\tau\tau}+\left(g^{\frac{p}{k}}\right)_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{\tau\tau}g-\frac{k-p}{k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+\left(\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$ (6.11)
in $\Omega$. When $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it follows from (6.11) and (2.1) that
$\mathcal{L}(w)$ $\geq-\frac{p}{k}g^{-\frac{1}{k}}\left|\partial_{\tau\tau}g-\frac{k-p}{k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right|-g^{-\frac{1}{k}}\left|\frac{p}{k}\partial_{x_{i}}g\tau_{j}\partial_{x_{j}}(\tau_{i})\right|$
$\geq-C_{1}f^{\frac{-1}{k(k-1)}}$
in $\Omega$, where $C_{1}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
When $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, which implies $(k+2)/3k<1/2$, it follows from (6.11), (2.2), and (2.1) that
$\mathcal{L}(w)$ $\geq\frac{p}{k}g^{\frac{-p}{k(k-1)}-\frac{1}{3}}\left(\partial_{\tau\tau}g-\frac{k+2}{3k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+g^{\frac{-p}{k(k-1)}}\left(\frac{p}{k}g^{\frac{1}{6}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$
$\geq-C_{2}f^{\frac{-1}{k(k-1)}}$
in $\Omega$, where $C_{2}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. This completes the proof of Lemma 6.3. ∎
Lemma 6.4.
There exist positive constants $r_{2}$ and $K_{2}$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)+K_{2}M\left(x_{n}-\rho\big(x^{\prime}\big)\right)\quad\text{in}\ \omega_{r_{2}},$
where $r_{2}$ depends on $\Omega$; and the constant $K_{2}$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Define a function
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}-K_{0}\big|x^{\prime}\big|^{4}\quad\text{in}\ \omega_{r_{0}},$
where the constant $\beta_{1}>0$ is to be determined. It holds that
$D^{2}v=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]-\frac{K_{0}}{M}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big]-D^{2}\Big[K_{0}\big|x^{\prime}\big|^{4}\Big].$
Since $\partial\Omega\in C^{3,1}$ is strictly (k1)(k-1)-convex, one can fix $\beta_{1}$ sufficiently small such that $\lambda(D^{2}[x_{n}^{2}+\beta_{1}\rho](0))$ belongs to a compact subset of $\Gamma_{k}$ independent of $0\in\partial\Omega$. By
$|x|\rightarrow 0,\ \rho\big(x^{\prime}\big)\rightarrow 0,\ \nabla_{x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow 0,\ D^{2}_{x^{\prime}x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow D^{2}_{x^{\prime}x^{\prime}}\rho(0)\quad\text{as}\ r\rightarrow 0,$
there exist two small positive constants $r_{2}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{2}$, and a large constant $C_{1}>0$ such that
$\left\{\lambda\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big):\ x\in\omega_{r_{2}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big),\quad\forall M\geq C_{1},$
and is thus contained in a compact subset of $\Gamma_{k}$ for all $M\geq C_{1}$. We emphasize that $\beta_{1}$, $r_{2}$, $\delta_{2}$, $C_{1}$ all depend only on $\Omega$. Using the concavity of $F$, we have
$F^{ij}\big[D^{2}u\big]\left(C_{2}(v_{ij}-\delta_{2}\delta_{ij})-u_{ij}\right)$ $\geq F\big[C_{2}\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=C_{2}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{2}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0$
in $\omega_{r_{2}}$, where $C_{2}$ is sufficiently large. Then by virtue of (2.14) and (2.13), we derive
[…]
in $\omega_{r_{2}}$, where $C>0$ is universal. Thus using (6.12) and Lemma 6.3, there exists a large constant $K>0$ such that
$\mathcal{L}(-KMv)<\mathcal{L}(-Kv)<\mathcal{L}(w-h)\quad\text{in}\ \omega_{r_{1}},$ (6.13)
where $K$ depends on $\delta_{0}$, $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
The boundary $\partial\omega_{r_{2}}$ consists of three parts: $\partial_{1}\omega_{r_{2}}\cup\partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}}$, where $\partial_{1}\omega_{r_{2}}$ and $\partial_{2}\omega_{r_{2}}$ are respectively the graph parts of $\rho$ and $\rho+r_{2}^{4}$, and $\partial_{3}\omega_{r_{2}}$ is the boundary part on $\{|x^{\prime}|=r_{2}\}$. On $\partial_{1}\omega_{r_{2}}$, using (6.5) one gets
$KMv=KK_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h.$
It follows from Theorems 1.1 and 1.2 that
$w-h\leq C_{3}M+C_{4}\quad\text{on}\ \partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}},$
where $C_{3}$ depends on $\Omega$; $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. It holds that
$KMv\geq KM\left(\beta_{1}r_{2}^{4}-r_{2}^{8}\right)\geq\frac{\beta_{1}}{2}KMr_{2}^{4}\quad\text{on}\ \partial_{2}\omega_{r_{2}},$
and
$KMv\geq KK_{0}\left(r_{2}^{2}+Mr_{2}^{4}\right)\quad\text{on}\ \partial_{3}\omega_{r_{2}}.$
Then we obtain $KMv\geq w-h$ on $\partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}}$ by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
Therefore, it follows from the maximum principle that
$KMv\geq w-h\quad\text{in}\ \omega_{r_{2}},$
which implies
$w-h\leq KM\left(\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)+\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}+K_{0}\big|x^{\prime}\big|^{4}\right)$
in $\omega_{r_{2}}$. This completes the proof of Lemma 6.4. ∎
Lemma 6.5.
For any $\sigma>0$, there exists a positive constant $K_{\sigma}$ such that
$\left(u_{(\tau)(\tau)}\right)_{n}(0)\leq\sigma M+K_{\sigma},$
where $K_{\sigma}$ depends on $\sigma$, $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Let $v$ be the function defined in the proof of Lemma 3.2, which is
$v=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{1}}.$
Recalling (3.8), we derive
[…]
where $C>0$ is universal. We emphasize that $\beta_{1}$, $\beta_{2}$, $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{1}$ all depend only on $\Omega$. It follows from Lemma 6.3 and (6.14) that there exists a large constant $K>0$ such that
$\mathcal{L}(-Kv)<\mathcal{L}(w-h)\quad\text{in}\ \omega_{r_{1}},$ (6.15)
where $K$ depends on $\delta_{0}$, $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
For any $r\leq\min\{r_{1},r_{2}\}$, the boundary of $\omega_{r}$ consists of three parts: $\partial_{1}\omega_{r}\cup\partial_{2}\omega_{r}\cup\partial_{3}\omega_{r}$, where $\partial_{1}\omega_{r}$ and $\partial_{2}\omega_{r}$ are respectively the graph parts of $\rho$ and $\rho+r^{4}$, and $\partial_{3}\omega_{r}$ is the boundary part on $\{|x^{\prime}|=r\}$. Using (6.5), one gets
$K(rM+1)v=K\beta_{2}(rM+1)\big|x^{\prime}\big|^{2}\geq K\beta_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h\quad\text{on}\ \partial_{1}\omega_{r},$ (6.16)
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from Lemma 6.4 that
$w-h\leq 2K_{2}Mr^{4}+K_{2}r^{2}\quad\text{on}\ \partial_{3}\omega_{r}.$
We have
$K(rM+1)v\geq K(rM+1)\beta_{2}r^{2}\geq K\beta_{2}Mr^{4}+K\beta_{2}r^{2}\geq w-h$ (6.17)
on $\partial_{3}\omega_{r}$, by taking $K$ large, which depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
Without loss of generality, we assume $|\nabla_{x^{\prime}}\rho|\leq 1$ on $\{|x^{\prime}|\leq r_{0}\}$. The distance function satisfies
$\operatorname{dist}(x,\partial\Omega)\geq\inf_{|y^{\prime}|\leq r_{0}}\left(\big|x^{\prime}-y^{\prime}\big|+r^{4}-\big|\nabla_{x^{\prime}}\rho\cdot\big(x^{\prime}-y^{\prime}\big)\big|\right)\geq r^{4},\quad\forall x\in\partial_{2}\omega_{r}.$
By virtue of the weakly interior estimate, namely Theorem 1.5, we infer
$w-h\leq\left|u_{ij}\tau_{i}\tau_{j}\right|+C_{1}\leq 2\sup_{\Omega_{r^{4}}}\left|D^{2}u\right|+C_{1}$ $\leq\frac{2}{r^{4}}\left(r^{9}\sup_{\partial\Omega}\left|D^{2}u\right|+C_{r^{9}}\right)+C_{1}$
$\leq 2Mr^{5}+\frac{2C_{r^{9}}}{r^{4}}+C_{1}$
on $\partial_{2}\omega_{r}$, where $C_{r^{9}}$ is given in (1.8); and $C_{1}>0$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Then we obtain
$K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)\frac{\beta_{1}}{2}r^{4}=\frac{K\beta_{1}}{2}Mr^{5}+\frac{K\beta_{1}}{2}\frac{C_{r^{9}}}{r^{4}}\geq w-h$ (6.18)
on $\partial_{2}\omega_{r}$, by taking the constant $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{1/(k-k/3-2/3)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Note that $\mathcal{L}(v)>0$ and $v<0$ in $\omega_{r}$. Applying the maximum principle with (6.15)–(6.18) yields
$K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq w-h\quad\text{in}\ \omega_{r},\quad\forall r\leq\min\{r_{1},r_{2}\},$
where the constant $K$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Therefore,
$\left(u_{(\tau)(\tau)}\right)_{n}(0)$ $=\lim_{x_{n}\rightarrow 0^{+}}\frac{u_{(\tau)(\tau)}(0,x_{n})-u_{(\tau)(\tau)}(0,0)}{x_{n}}$
$=\lim_{x_{n}\rightarrow 0^{+}}\frac{w(0,x_{n})-h(0)}{x_{n}}$
$\leq K\beta_{1}\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)$
for all $r\leq\min\{r_{1},r_{2}\}$. Let
$\sigma:=K\beta_{1}r\quad\text{and}\quad K_{\sigma}:=K\beta_{1}\frac{C_{r^{9}}}{r^{8}}.$
This completes the proof of Lemma 6.5. ∎
Lemma 6.6.
It holds that
$\left(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right)(x)\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$
for all points $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $K_{\sigma}$ is given in Lemma 6.5.
The proof of Lemma 6.6 is deferred to Appendix. We now proceed to prove Theorem 1.6.
Proof of Theorem 1.6.
Without loss of generality, let the maximum $M$ be attained at the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, (6.1) hold, $\varphi(0)=0$, and $\nabla\varphi(0)=0$.
Applying Taylor expansion to $h(t):=u_{n}(tx^{\prime},\rho(tx^{\prime}))$, we obtain that for any $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$,
$u_{n}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}=\frac{1}{2}u_{n(\eta)(\eta)}\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\left|x^{\prime}\right|^{2},$ (6.19)
where $t\in(0,1)$ depends on $x$, and the tangential vector field $\eta$ is defined in (6.3). Since $u=\varphi$ on $\partial\Omega$, the third-order tangential derivatives of $u$ are bounded (see Lemma 8.1 in the Appendix). Using Lemma 6.6, we derive
$u_{n(\eta)(\eta)}(y)$ $=\left[u_{n\eta\eta}+u_{ni}\eta_{j}\partial_{x_{j}}(\eta_{i})\right](y)$
$=\langle e_{n},\nu\rangle\left[u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right](y)+C_{1}$
$\leq\sigma M+K_{\sigma}+C_{1}$ (6.20)
for all $y\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ and for all $\sigma>0$, where by Theorem 1.1 the constant $C_{1}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from (6.19) and (6.1) that
$u_{n}(x)-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\leq(\sigma M+K_{\sigma})\big|x^{\prime}\big|^{2}$ (6.21)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, by increasing $K_{\sigma}$.
Let $v$ be the function defined in the proof of Lemma 3.2, which is
$v=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{1}}.$
Recalling (3.8), we derive
[…]
where $C>0$ is universal. We emphasize that $\beta_{1}$, $\beta_{2}$, $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{1}$ all depend only on $\Omega$. By (2.10) and (2.1), we infer
$\mathcal{L}(u_{n})=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{n}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}}\geq-C_{2}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$ (6.23)
where $C_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Define
$\tilde{v}:=-K(\sigma M+K_{\sigma})v\quad\text{in}\ \overline{\omega_{r_{1}}},$
where the constant $K>0$ is to be determined. By (6.22) and (6.23), we obtain
$\mathcal{L}(\tilde{v})\leq\mathcal{L}\left(u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\right)\quad\text{in}\ \omega_{r_{1}},$
by taking the constant $K$ large, which depends on $\delta_{0}$, $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Using (6.21), one has
$\tilde{v}=K\beta_{2}(\sigma M+K_{\sigma})\big|x^{\prime}\big|^{2}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{on}\ \partial_{1}\omega_{r_{1}},$
by taking $K$ large, which depends on $\Omega$. It follows that
$\tilde{v}\geq K(\sigma M+K_{\sigma})\min\left\{\frac{\beta_{1}}{2}r_{1}^{4},\beta_{2}r_{1}^{2}\right\}$
on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$. According to Theorem 1.1, we get
$\tilde{v}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{on}\ \partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}},$
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Hence applying the comparison principle yields
$\tilde{v}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{in}\ \omega_{r_{1}},$
where the constant $K$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Therefore,
$M=u_{nn}(0)=\lim_{x_{n}\rightarrow 0^{+}}\frac{u_{n}(0,x_{n})-u_{n}(0)}{x_{n}}\leq\partial_{x_{n}}\tilde{v}(0)=K\beta_{1}(\sigma M+K_{\sigma}),$
where $K_{\sigma}$ is given in Lemma 6.5. Taking $\sigma=1/(2K\beta_{1})$, we obtain the upper bound of $M$. This completes the proof of Theorem 1.6. ∎
6.2. The case of exponent $\mathbf{3/(2k)}$
In this subsection, we assume $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. In fact, Lemmas 2.5 and 6.3 can be improved.
Lemma 6.7.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$, and $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. Then
$\operatorname{tr}\left(GD^{2}u_{\xi\xi}\right)\geq-K|\xi|^{2},\quad\forall\xi\in\mathbb{R}^{n},$
in $\Omega$, where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
It follows from (2.11), (2.2), and (2.13) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq\frac{1}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}\frac{2}{3}g^{-\frac{1}{3}}\left(\partial_{ee}g-\frac{1}{3}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)\geq-\frac{K}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}\geq-CK$
for all $e\in\mathbb{S}^{n-1}$, where $g=f^{3/(2k)}$, $C>0$ is universal, and $K>0$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. ∎
Lemma 6.8.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$, and $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. Then
$\mathcal{L}(w)\geq-C\quad\text{in}\ \Omega,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
It follows from (6.11) that
$\mathcal{L}(w)\geq\frac{2}{3}g^{-\frac{1}{3}}\left(\partial_{\tau\tau}g-\frac{1}{3}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+\left(\frac{2}{3}g^{\frac{1}{6}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$
in $\Omega$, where $g=f^{3/(2k)}$. Using Lemma 2.1, we obtain
$\mathcal{L}(w)\geq-C,$
where $C$ depends on $\|\rho\|_{C^{3}}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. ∎
Using Lemmas 6.7 and 6.8, with the same proof as in Theorem 1.6, we prove Theorem 1.9.
7. The optimal condition for the Monge-Ampère equation
In this section, we assume
$f^{\frac{3}{2n-2}}\in C^{2,1}(\overline{\Omega}),$
and apply the method of Guan-Trudinger-Wang [12] to establish the boundary second-order normal derivative estimate.
Let the origin be a boundary point satisfying $\nu(0)=e_{n}$. By an affine transformation of $x^{\prime}$, let $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)=\frac{1}{2}\big|x^{\prime}\big|^{2}+\text{cubic of $x^{\prime}$}+O\Big(\big|x^{\prime}\big|^{4}\Big)$ (7.1)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. Then rotate $x^{\prime}$ such that $D^{2}_{x^{\prime}x^{\prime}}u(0)$ is a diagonal matrix with
$0<u_{11}(0)\leq u_{22}(0)\leq\dots\leq u_{n-1,n-1}(0).$
By subtracting an affine function, let
$u(0)=0\quad\text{and}\quad\nabla u(0)=0,$
which yields by the convexity of $u$,
$u(x)\geq 0,\quad\forall x\in\overline{\Omega}.$
Differentiating the boundary condition $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ yields
$D^{2}_{x^{\prime}x^{\prime}}u(0)=D^{2}_{x^{\prime}x^{\prime}}\varphi(0)+D^{2}_{x^{\prime}x^{\prime}}\rho(0)\partial_{n}\varphi(0).$
For convenience, denote
$b_{i}:=u_{ii}(0),\quad\forall 1\leq i\leq n-1.$
By Taylor expansion with respect to $x^{\prime}$ at the origin, we have
$\varphi\big(x^{\prime},\rho\big(x^{\prime}\big)\big)=\frac{1}{2}\sum_{i=1}^{n-1}b_{i}x_{i}^{2}+\text{cubic of $x^{\prime}$}+O\Big(\big|x^{\prime}\big|^{4}\Big)$ (7.2)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
The following lemma is the inequality (3.5) in Guan-Trudinger-Wang [12].
Lemma 7.1 ([12]).
Let $h\in C^{3,1}(\mathbb{R})$ be a nonnegative function on $\mathbb{R}$ satisfying
$h(t)=\alpha t^{2}+\beta t^{3}+R(t),\quad\forall t\in\mathbb{R},$
where $\alpha\geq 0$, $\beta\in\mathbb{R}$, and the remainder term satisfies $|R(t)|\leq\gamma|t|^{4}$ for some $\gamma>0$. Then
$|\beta|\leq(1+\gamma)\sqrt{\alpha}.$
Proof.
Lemma 7.1 is obtained by combining two facts $h(\sqrt{\alpha})\geq 0$ and $h(-\sqrt{\alpha})\geq 0$. ∎
7.1. Lower bounds for boundary second-order tangential derivatives
Lemma 7.2.
There exists a constant $\gamma_{0}>0$ such that
$\prod_{i=1}^{n-1}b_{i}\geq\gamma_{0}f(0),$
where $\gamma_{0}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
Proof.
Step 1. Define
$M:=\frac{1}{b_{1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (7.3)
where
$M_{i}:=\sqrt{b_{i}}M\ \big(\geq\sqrt{M}\big)\ \ \text{for}\ i=1,2,\dots,n-1,\quad\text{and}\quad M_{n}:=M.$
Without of loss generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=\frac{f(T^{-1}y)}{\prod_{i=1}^{n-1}b_{i}}=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
It follows from (7.1) and (7.3) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}y_{1}^{2}+\frac{1}{2}\sum_{i=2}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (7.4)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$d_{i}\leq 1\ \ \text{for}\ i=2,\dots,n-1,\quad\big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\quad\big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (7.2) and (7.3) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}\big|y^{\prime}\big|^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (7.5)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 7.1, we conclude that the coefficients of the third-order terms in (7.5) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Step 2. Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1,\ |y_{i}|<1\ \text{for}\ i=2,\dots,n-1\right\}.$
We analyze the geometry of $\omega$. Clearly, the set $\omega$ is convex. Looking at $\partial\omega\cap\{y_{n}=1\}$, by (7.4) we have
$1\geq\frac{1}{2}y_{1}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|y_{1}|=2$ must be outside of $\omega$. The connectedness of $\omega$ yields $|y_{1}|\leq 2$, so the set $\omega$ is contained in the ball $B_{\sqrt{n+3}}$. It is trivial that
$y_{\star}:=\left(0,\frac{1}{4n},\dots,\frac{1}{4n},\frac{1}{2}\right)\in\omega,$ (7.6)
and
[…]
Then, since $\tilde{\rho}(y^{\prime})\leq|y^{\prime}|^{2}$, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where
$r_{\star}:=\min\left\{\frac{1}{2},1-\frac{1}{4n},\min\left\{\frac{1}{4n},\frac{1}{2}-\frac{1}{4n}\right\}\right\}.$ (7.7)
Using the convexity of $v$, we derive that the maximum of $v$ in $\overline{\omega}$ is attained on $\partial\omega\cap\partial\widetilde{\Omega}$. Thus, by (7.5) we obtain
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$ (7.8)
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Step 3. We claim that there exists a constant $K>0$, independent of $M$, such that for any ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$ with arbitrary center $y_{*}$,
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\tilde{f}<K.$ (7.9)
Suppose, for contradiction, that (7.9) does not hold. Then
$\det D^{2}v\geq K\quad\text{in}\ B_{\frac{r_{\star}}{4n}}(y_{*}).$
Since (7.8) the function
$\frac{K^{1/n}}{2}\left((y-y_{*})^{2}-\frac{r_{\star}}{4n}\right)+C_{2}$
is a upper barrier of $v$ in $B_{\frac{r_{\star}}{4n}}(y_{*})$. Thus
$v(y_{*})\leq-K^{\frac{1}{n}}\frac{r_{\star}}{8n}+C_{2}<0$
by taking $K$ large, which depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$. But this contradicts with $v\geq 0$. Therefore, we obtain claim (7.9).
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
We have
$\left\|D_{y}^{3}\tilde{g}\right\|_{L^{\infty}(\widetilde{\Omega})}\leq\left(\prod_{i=1}^{n-1}b_{i}\right)^{\frac{-3}{2n-2}}b_{1}^{\frac{3}{2}}\left\|D_{x}^{3}\left(f^{\frac{3}{2n-2}}\right)\right\|_{L^{\infty}(\Omega)}\leq\left\|f^{\frac{3}{2n-2}}\right\|_{C^{2,1}(\overline{\Omega})}\leq C.$ (7.10)
From (7.9), there is a point $\tilde{y}\in B_{\frac{r_{\star}}{4n}}(y_{\star})$ satisfying $\tilde{f}(\tilde{y})\leq K$. Similar to (2.3), the fact
$0\leq\tilde{g}\Big(\tilde{y}+\frac{r_{\star}}{2}e\Big)+\tilde{g}\Big(\tilde{y}-\frac{r_{\star}}{2}e\Big)\leq 2K^{\frac{3}{2(n-1)}}+\partial_{ee}\tilde{g}(\tilde{y})\left(\frac{r_{\star}}{2}\right)^{2}+O\big(r_{\star}^{3}\big)$
yields
$\partial_{ee}\tilde{g}(\tilde{y})\geq-\frac{C}{r_{\star}^{2}}\left(K^{\frac{3}{2(n-1)}}+1\right),\quad\forall e\in\mathbb{S}^{n-1}.$ (7.11)
Combining (7.10) and (7.11) yields that there exists a constant $K_{1}>0$ such that
$D^{2}\tilde{g}+K_{1}\mathbf{I}_{n}\quad\text{is positive definite in $\omega$},$
that is $\tilde{g}$ is semi-convex in $\omega$. Define
$\bar{g}:=\tilde{g}+\frac{K_{1}}{2}|y|^{2}\quad\text{in}\ \overline{\omega}.$
Then $\bar{g}$ is convex in $\omega$, and there exists a constant $\overline{K}>0$ such that
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\bar{g}<\overline{K}$ (7.12)
for every ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$ with arbitrary center $y_{*}$. By (7.12) and the convexity of $\bar{g}$, it follows (see Figure 1) that
$|D\bar{g}|\leq\frac{2\overline{K}}{r_{\star}/4}\quad\text{in}\ B_{\frac{r_{\star}}{4}}(y_{\star}),$
and then by the mean value theorem,
$\partial_{ee}\bar{g}(y_{e})\leq\frac{4\overline{K}}{r_{\star}/4}\frac{1}{r_{\star}/4},\quad\forall e\in\mathbb{S}^{n-1},$
for some point $y_{e}\in\omega$ depending on $e$. Combining this and (7.10), we obtain
$\|\tilde{g}\|_{C^{2,1}(\overline{\omega})}\leq K_{2},$
where $K_{2}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
Figure 1. Use the convexity of $\bar{g}$ to get a uniform bound on its first-order derivatives in $B_{\frac{r_{\star}}{4}}(y_{\star})$, and further a uniform bound on its second-order derivatives at some points is obtained by the mean value theorem.
Therefore,
$K_{2}^{\frac{2(n-1)}{3}}\geq\tilde{f}(0)=\frac{f(0)}{\prod_{i=1}^{n-1}b_{i}}.$
This completes the proof of Lemma 7.2. ∎
7.2. Lower bounds for boundary tangential-normal derivatives
Lemma 7.3.
For any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\frac{1}{2}\sqrt{b_{n-1}}$, we have
$\left|u_{\nu\eta}(x)\right|\leq C\sqrt{b_{n-1}},$
where $\eta$ is any unit tangential vector to $\partial\Omega$ at $x$, and $C>0$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
Proof.
Define
$M:=\frac{1}{b_{n-1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y^{\prime}=\sqrt{M}x^{\prime}\quad\text{and}\quad y_{n}=Mx_{n}.$ (7.13)
Without of loss generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=M^{n-1}f\big(T^{-1}y\big)=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
It follows from (7.1) and (7.13) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}\big|y^{\prime}\big|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (7.14)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\quad\big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (7.2) and (7.13) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}\sum_{i=1}^{n-1}\tilde{b}_{i}y_{i}^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (7.15)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\tilde{b}_{i}\leq 1\ \ \text{for}\ i=1,2,\dots,n-1,\quad\text{and}\quad\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 7.1, we conclude that the coefficients of the third-order terms in (7.15) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1\right\}.$
Looking at $\partial\omega\cap\{y_{n}=1\}$, by (7.14) we have
$1=\frac{1}{2}\big|y^{\prime}\big|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|y^{\prime}|=2$ must be outside of $\omega$. So the set $\omega$ is contained in the ball $B_{2}$ by the convexity of $\omega$. Similar to Step 2 in the proof of Lemma 7.2, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where $y_{\star}$ and $r_{\star}$ are given in (7.6) and (7.7), respectively. Furthermore, using the convexity of $v$ and (7.15), we derive
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
By the same argument as in Step 3 of the proof of Lemma 7.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq K,$
where $K$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Note that $\partial\widetilde{\Omega}$ is uniformly convex on $\partial\widetilde{\Omega}\cap\partial\omega$ by (7.14). The first-order normal estimate on the boundary follows from Theorem 17.21 of Gilbarg-Trudinger [7] that
$\left|v_{\tilde{\nu}}(y)\right|\leq C_{3},\quad\forall y\in\partial\widetilde{\Omega}\cap\left\{y_{n}\leq 7/8\right\},$ (7.16)
where $\tilde{\nu}=\tilde{\nu}(y)$ denotes the unit interior normal vector to $\partial\widetilde{\Omega}$ at $y$, and $C_{3}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Using the convexity of $v$ and the uniform convexity of $\partial\widetilde{\Omega}$, it follows from (7.16) that
$|Dv|\leq C_{4}\quad\text{in}\ \widetilde{\Omega}\cap\left\{y_{n}\leq 3/4\right\},$
where $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. In view of Remark 2.2, to weaken the condition to $\inf_{\Omega}\tilde{g}>0$ when applying Lemma 2.1, we adapt the proof of Lemma 3.2 by differentiating with respect to a tangential vector filed to establish the boundary tangential-normal estimate. The convexity of $v$ and the uniform convexity of $\partial\widetilde{\Omega}$ play key roles in this argument. Define tangential vector fields
$\tilde{\eta}^{l}(y):=e_{l}+\tilde{\rho}_{l}\big(y^{\prime}\big)e_{n}\quad\text{for}\ l=1,2,\dots,n-1,$
in $\overline{\omega}$. Consider functions
$\Psi^{l}(y):=\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A|y|^{2}-B_{A}y_{n}\quad\text{in}\ \overline{\omega},$
for all $l=1,2,\dots,n-1$, where positive constants $A$ and $B_{A}$ are to be determined. We can obtain
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
and similarly
$\left|v_{\tilde{\nu}\tilde{\eta}^{l}}(y)\right|\leq C,\quad\forall y\in\partial\widetilde{\Omega}\cap\left\{y_{n}\leq 1/2\right\},\quad\forall 1\leq l\leq n-1.$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Pulling back to the $x$-coordinates, we finish the proof of Lemma 7.3. Since a more complicated version of these arguments will be needed later, in the absence of uniform convexity of $\partial\widetilde{\Omega}$, we omit the details here. ∎
We proceed from Lemma 7.3 by induction to obtain a refinement. Define
$\eta^{l}=\eta^{l}(x):=e_{l}+\rho_{l}\big(x^{\prime}\big)e_{n},\quad\forall x\in\partial\Omega\cap\left\{\big|x^{\prime}\big|\leq r_{0}\right\},$ (7.17)
for all $l=1,2,\dots,n-1$. In fact, $\eta^{l}(x)$ is the tangential direction of $\partial\Omega$, which lies in the two-dimensional plane parallel to the $x_{l}$- and $x_{n}$-axes and passes through the point $x$. Our induction hypothesis is the following.
Assumption 7.4.
For some fixed $k\in\{1,2,\dots,n-2\}$ and each $i=k+1,\dots,n-1$, there exists a constant $\theta_{i}>0$ such that for any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\theta_{i}\sqrt{b_{i}}$, we have
$\left|u_{\nu\eta^{j}}(x)\right|\leq C\sqrt{b_{i}},\quad\forall 1\leq j\leq i,$
where constants $\theta_{i}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$.
Lemma 7.5.
Let Assumption 7.4 hold for some fixed $k\in\{1,2,\dots,n-2\}$. Then there exists a constant $\theta_{k}>0$ such that for any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\theta_{k}\sqrt{b_{k}}$, we have
$\left|u_{\nu\eta^{j}}(x)\right|\leq C\sqrt{b_{k}},\quad\forall 1\leq j\leq k,$ (7.18)
where constants $\theta_{k}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$.
Proof.
Define
$M:=\frac{1}{b_{k}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}:=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (7.19)
where
$M_{i}:=\begin{cases}\sqrt{M},\quad i=1,\dots,k,\\ \sqrt{b_{i}}M\ (\geq\sqrt{M}),\quad i=k+1,\dots,n-1,\\ M,\quad i=n.\end{cases}$
Without of loss generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=\frac{M^{k}f(T^{-1}y)}{\prod_{i=k+1}^{n-1}b_{i}}=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
Denote
$\hat{y}:=(y_{1},\dots,y_{k})\quad\text{and}\quad\tilde{y}:=(y_{k+1},\dots,y_{n-1}).$
It follows from (7.1) and (7.19) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}|\hat{y}|^{2}+\frac{1}{2}\sum_{i=k+1}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (7.20)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$d_{i}=\frac{b_{k}}{b_{i}}\ (\leq 1)\ \ \text{for}\ i=k+1,\dots,n-1,\quad\big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\quad\big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (7.2) and (7.19) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}|\tilde{y}|^{2}+\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}y_{i}^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (7.21)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\tilde{b}_{i}=\frac{b_{i}}{b_{k}}\ (\leq 1)\ \ \text{for}\ i=1,2,\dots,k,\quad\text{and}\quad\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 7.1, we conclude that the coefficients of the third-order terms in (7.21) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<\beta^{2},\ |y_{i}|<\beta\ \text{for}\ i=k+1,\dots,n-1\right\},$
where $\beta$ is a small positive constant to be determined. We analyze the geometry of $\omega$. Clearly, the set $\omega$ is convex. Looking at $\partial\omega\cap\{y_{n}=\beta^{2}\}$, by (7.20) we have
$\beta^{2}\geq\frac{1}{2}|\hat{y}|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|\hat{y}|=2\beta$ must be outside of $\omega$. So the set $\omega$ is contained in the ball $B_{\sqrt{n+3}\beta}$ by the convexity of $\omega$. It is trivial that
$y_{\star}:=\left(0,\dots,0,\frac{\beta}{4n},\dots,\frac{\beta}{4n},\frac{\beta^{2}}{2}\right)\in\omega,$
and
[…]
Then, since $\tilde{\rho}(y^{\prime})\leq|y^{\prime}|^{2}$, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where
$r_{\star}:=\min\left\{\frac{\beta^{2}}{2},\beta-\frac{\beta}{4n},\min\left\{\frac{\beta}{4n},\frac{\beta^{2}}{2}-\frac{\beta^{2}}{4n}\right\}\right\}.$
Using the convexity of $v$ and (7.21), we derive
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$ (7.22)
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$. Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega},\quad\text{and}\quad p=\frac{2n-2}{3}.$
By the same argument as in Step 3 of the proof of Lemma 7.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq C_{3},$
where $C_{3}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
The main difficulty is that we cannot control the convexity of $\partial\widetilde{\Omega}$ in (7.20). An observation is that the directions, in which the convexity is under control, with respect to $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ complement each other.
Step 1: we prove $|Dv|\leq C$ in $\omega$. Define
$\underline{v}(y):=\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}+\frac{1}{2}Ky_{n}^{2}-Ky_{n},$
where constants $\sigma>0$ and $K>0$ are to be determined. Let $K$ large depending on $\sigma$ such that
$\det D^{2}\underline{v}=\sigma^{n-1}K\geq 2^{n}\sup_{\omega}\tilde{f}$. (7.23)
We now verify that
$\underline{v}\leq v\quad\text{on}\ \partial\omega.$ (7.24)
The boundary $\partial\omega$ consists of three parts: $\partial_{1}\omega\cup\partial_{2}\omega\cup\partial_{3}\omega$, where $\partial_{1}\omega:=\partial\omega\cap\partial\widetilde{\Omega}$, $\partial_{2}\omega:=\partial\omega\cap\{y_{n}=\beta^{2}\}$, and $\partial_{3}\omega:=\partial\omega\cap\{|y_{i}|=\beta\ \text{for some}\ i=k+1,\dots,n-1\}$. It follows from (7.21) that on $\partial_{1}\omega$,
$v=\frac{1}{2}|\tilde{y}|^{2}+\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}y_{i}^{2}+O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}|\tilde{y}|^{2}-|\hat{y}|^{2}$
by taking $\beta$ small. By (7.20), we have
$y_{n}=\frac{1}{2}|\hat{y}|^{2}+\frac{1}{2}\sum_{i=k+1}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}|\hat{y}|^{2}-\beta|\tilde{y}|^{2}$
on $\partial_{1}\omega$, since $M\gg 1$. Hence
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}-\frac{1}{2}Ky_{n}\leq\left(\frac{1}{2}\sigma+\frac{1}{2}K\beta\right)\left|\tilde{y}\right|^{2}+\left(\frac{1}{2}\sigma-\frac{1}{8}K\right)\left|\hat{y}\right|^{2}\leq v\quad\text{on}\ \partial_{1}\omega,$
by taking $\sigma<1/4$ and $\beta$ small depending on $K$. It holds that
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}-\frac{1}{2}K\beta^{2}\leq\frac{1}{2}\beta^{2}((n+3)\sigma-K)\leq 0\leq v\quad\text{on}\ \partial_{2}\omega.$
We only consider the piece $\partial^{\prime}_{3}\omega:=\partial\omega\cap\{y_{n-1}=\beta\}$, since other pieces of $\partial_{3}\omega$ can be handled in the same way. Let $\varepsilon_{0}>0$ be a small parameter to be determined later. It holds that
$\underline{v}\leq\frac{1}{2}\beta^{2}((n+3)\sigma-K\varepsilon_{0})\leq 0\leq v\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}\geq\varepsilon_{0}\beta^{2}\right\}$
by taking $\sigma$ small depending on $\varepsilon_{0}$.
We claim that
$v\geq\frac{1}{4}\beta^{2}\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\}.$ (7.25)
It follows from (7.21) that
$v\geq\frac{1}{2}\beta^{2}+O\big(\beta^{3}\big)\geq\frac{1}{4}\beta^{2}\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\}\cap\partial\widetilde{\Omega}$,
by taking $\beta$ small. It remains to consider points $p=(\hat{p},\tilde{p},p_{n})$ on $\partial^{\prime}_{3}\omega\cap\{y_{n}<\varepsilon_{0}\beta^{2}\}\cap\widetilde{\Omega}$, where $\hat{p}:=(p_{1},\dots,p_{k})$ and $\tilde{p}:=(p_{k+1},\dots,p_{n-1})$. Define
[…]
and
$p^{*}:=(0,\tilde{p},p_{n}+\delta),\quad p^{0}:=(0,\tilde{p},\tilde{\rho}(0,\tilde{p}))\in\partial\widetilde{\Omega}$.
From (7.20), we deduce
$\left|\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})\right|=O\big(\beta^{2}\big).$
Then $\delta=O(\beta^{3})$. Since $\widetilde{\Omega}$ is convex,
$p^{*}_{n}=p_{n}+\delta>p^{0}_{n}+\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})\cdot\hat{p}+\delta\geq p^{0}_{n},$
which implies $p^{*}\in\widetilde{\Omega}$. Extend the segment $\overline{pp^{*}}$ until it intersects $\partial\widetilde{\Omega}$ at a point denoted by $\bar{p}$ (see Figure 2).
Figure 2. The convexity of $\partial\widetilde{\Omega}$ in the direction $\tilde{y}$ is not under control. Cut $\widetilde{\Omega}$ by the hyperplane $\tilde{y}=\tilde{p}$ through the point $p$, and relate it to points on the boundary $\partial\widetilde{\Omega}$.
From (7.20), the graph of $\partial\widetilde{\Omega}$ near $p^{0}$ in the direction $\hat{y}$ is a perturbation of the paraboloid $|\hat{y}|^{2}/2$, by taking $\beta$ small. Hence we have
$\frac{1}{2}|p^{*}-p|\leq|\bar{p}-p^{*}|.$ (7.26)
Since the slope of $\overline{pp^{*}}$ is $|\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})|=O(\beta^{2})$, we solve
$\frac{1}{4}t^{2}\leq\left(\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)\right)+O\big(\beta^{2}\big)t$
to deduce $t=O(\beta)$. Thus
$\bar{p}_{n}\leq\left(\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)\right)+O\big(\beta^{3}\big)\leq 2\varepsilon_{0}\beta^{2}$ (7.27)
by taking $\beta$ small depending on $\varepsilon_{0}$. Using (7.20) and (7.27) we derive
$\frac{1}{2}\sum_{i=1}^{k}\left|\bar{p}_{i}\right|^{2}\leq\bar{p}_{n}+O\big(\beta^{3}\big)\leq 3\varepsilon_{0}\beta^{2}.$ (7.28)
For any $T^{-1}x=y=(0,\tilde{y},\tilde{\rho}(0,\tilde{y}))\in\partial\widetilde{\Omega}\cap B_{\theta_{k+1}}(0)$ with $\theta_{k+1}$ given in Assumption 7.4, we have
$\hat{x}=0\quad\text{and}\quad|x_{i}|=\left|\frac{y_{i}}{M_{i}}\right|\leq\theta_{k+1}\frac{b_{k}}{\sqrt{b_{i}}},\quad\forall k+1\leq i\leq n-1.$
Then by (6.8), (6.7), and Assumption 7.4, we infer
$\left|v_{y_{n}}(y)\right|=M\left|u_{x_{n}}(x)\right|$ $=\frac{1}{b_{k}}\left(\left|u_{\nu}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)\right|+O\Big(\big|x^{\prime}\big|^{2}\Big)\right)$
$=\frac{1}{b_{k}}\left|u_{n}(0)+\nabla_{x^{\prime}}(u_{\nu})\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\cdot x^{\prime}\right|+C\theta_{k+1}^{2}$
$=\frac{1}{b_{k}}\sum_{i=k+1}^{n-1}(u_{\nu\eta_{i}}+\partial_{x_{i}}\nu\cdot\nabla u)\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)x_{i}+C\theta_{k+1}^{2}$
$\leq\frac{1}{b_{k}}\sum_{i=k+1}^{n-1}\left(C\sqrt{b_{i}}x_{i}+C\big|x^{\prime}\big|^{2}\right)+C\theta_{k+1}^{2}$
$\leq C_{4}$
for some $t\in(0,1)$, where $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. By the convexity of $v$, (7.21), and (7.27), we get
$v(p^{*})\geq v\big(p^{0}\big)+\nabla v\big(p^{0}\big)\cdot\big(p^{*}-p^{0}\big)\geq\widetilde{\varphi}\big(p^{0}\big)-C_{4}\bar{p}_{n}\geq\frac{1}{2}\beta^{2}-2C_{4}\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)\geq\frac{1}{3}\beta^{2}$ (7.29)
by taking $\varepsilon_{0}$ small. Furthermore, we obtain from (7.21) and (7.27)–(7.29) that
$v(\bar{p})=\widetilde{\varphi}(\bar{p})=\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}\left|\bar{p}_{i}\right|^{2}+\widetilde{\varphi}\big(p^{0}\big)+O\big(\beta^{3}\big)$ $\leq v(p^{*})+C_{4}\bar{p}_{n}+4\varepsilon_{0}\beta^{2}$
$\leq v(p^{*})+(2C_{4}+4)\varepsilon_{0}\beta^{2}$
$\leq\frac{9}{8}v(p^{*})$ (7.30)
by taking $\varepsilon_{0}$ small. Using the convexity of $v$, (7.30), (7.26), and (7.29), we get
$v(p)\geq-\frac{|p^{*}-p|}{|\bar{p}-p^{*}|}v(\bar{p})+\frac{|\bar{p}-p|}{|\bar{p}-p^{*}|}v(p^{*})\geq\left(1-\frac{1}{8}\frac{|p^{*}-p|}{|\bar{p}-p^{*}|}\right)v(p^{*})\geq\frac{1}{4}\beta^{2}.$
In conclusion, claim (7.25) is valid.
We deduce from (7.25) that
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}\leq\frac{n+3}{2}\sigma\beta^{2}\leq v\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\},$
by taking $\sigma$ small. Thus (7.24) holds. It follows from the comparison principle that
$\underline{v}\leq v\quad\text{in}\ \overline{\omega}.$
Since $\underline{v}(0)=v(0)=0$,
$v_{n}(0)\geq\underline{v}_{n}(0).$
Using the convexity of $v$ and (7.22), we get
$v_{n}(0)\leq\frac{v((0,\dots,0,\beta^{2}))-v(0)}{\beta^{2}}\leq\frac{C_{2}}{\beta^{2}}.$
So $|v_{n}(0)|\leq C$. For any point $y_{0}\in\partial\omega\cap\partial\widetilde{\Omega}$, letting $y_{0}$ be the origin and $\tilde{\nu}(y_{0})=e_{n}$ by translation and rotation, letting $v(y_{0})=0$ and $\nabla v(y_{0})=0$ by subtracting an affine function, with similar arguments as above, we get
$|v_{\tilde{\nu}}(y_{0})|\leq C,\quad\forall y_{0}\in\partial\omega\cap\partial\widetilde{\Omega}$, (7.31)
where $\tilde{\nu}=\tilde{\nu}(y)$ denotes the unit interior normal vector to $\partial\widetilde{\Omega}$ at $y$. From (7.31) and the convexity of $v$, it follows that
$|Dv|\leq C\quad\text{in}\ \overline{\omega}$
for a smaller $\beta$, where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$.
Step 2. Define tangential vector fields
$\tilde{\eta}^{l}(y):=e_{l}+\tilde{\rho}_{l}\big(y^{\prime}\big)e_{n}\quad\text{for}\ l=1,2,\dots,n-1,$
in $\overline{\omega}$. Indeed, for any $l\in\{1,2,\dots,n-1\}$,
$T\eta^{l}=T(e_{l}+\partial_{x_{l}}\rho e_{n})=M_{l}e_{l}+\partial_{x_{l}}\rho M_{n}e_{n}=M_{l}\left(e_{l}+\partial_{y_{l}}\tilde{\rho}e_{n}\right)=M_{l}\tilde{\eta}^{l}.$ (7.32)
Construct functions
$\Psi^{l}(y):=\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A|y|^{2}-B_{A}\bigg(v-\frac{1}{2}\underline{v}+(C_{5}+1)y_{n}\bigg)\quad\text{in}\ \overline{\omega},$
for $l=1,2,\dots,n-1$, where positive constants $A$ and $B_{A}$ are to be determined, and $C_{5}:=\sup_{\omega}|D(v-\underline{v}/2)|$.
First, we choose $A$ large so that the functions $\Psi^{l}$ cannot attain its maximum in $\omega$ for $l=1,2,\dots,n-1$. Fix $l\in\{1,2,\dots,n-1\}$, and suppose, for contradiction, that $\Psi^{l}$ attains its maximum at an interior point $y_{*}\in\omega$. Then at the point $y_{*}$, we have
$\partial_{n}\Psi^{l}=0\quad\text{and}\quad F^{ij}\big[D^{2}v\big]\Psi^{l}_{ij}\leq 0.$ (7.33)
From (7.33), at the point $y_{*}$,
$\pm\partial_{n}\Big(\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})\Big)=B_{A}\bigg(\bigg(v-\frac{1}{2}\underline{v}\bigg)_{n}+C_{5}+1\bigg)-2Ay_{n}\geq B_{A}-2A,$
which yields $|D^{2}v(y_{*})|>1$ by taking $B_{A}$ large depending on $A$. Since $\lambda(D^{2}v(y_{*}))\in\Gamma_{2}$,
$\sigma_{1}\big[D^{2}v(y_{*})\big]\geq\left|D^{2}v(y_{*})\right|>1.$ (7.34)
For convenience, introduce a coordinate system $z$, obtained by an orthogonal transformation of the $y$-coordinates, such that at the point $y_{*}$ in the $z$-coordinates,
$D_{z}^{2}v\ \ \text{is diagonal, which implies}\ \ F^{ij}_{z}\big[D^{2}_{z}v\big]\ \ \text{is also diagonal}.$
Computing (7.33) at the point $y_{*}$ in the $z$-coordinates, by the concavity and homogeneity of $F$ and (7.23), we obtain
$0$ $\geq F^{ij}_{z}\big[D^{2}_{z}v\big]\Psi^{l}_{z_{i}z_{j}}$
$=F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\pm\tilde{\eta}^{l}_{z}\cdot\nabla_{z}(v-\widetilde{\varphi})\right)_{z_{i}z_{j}}+F^{ij}\big[D^{2}v\big]\left(A|y|^{2}+B_{A}\bigg(\frac{1}{2}\underline{v}-v\bigg)\right)_{ij}$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}(v-\widetilde{\varphi})\right)_{z_{i}z_{j}}+2A\operatorname{tr}F^{ij}\big[D^{2}v\big]+B_{A}\bigg(\frac{1}{2}F\big[D^{2}\underline{v}\big]-F\big[D^{2}v\big]\bigg)$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}z_{j}}+2\partial_{z_{j}}\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}}\right)+(2A-C)\operatorname{tr}F^{ij}\big[D^{2}v\big]$
$=\pm\tilde{\eta}^{l}\cdot\nabla\left(\tilde{g}^{\frac{p}{n}}\right)\pm 2\sum_{i=1}^{n}F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\times\partial_{z_{i}}\tilde{\eta}^{l,i}_{z}+(2A-C)\operatorname{tr}F^{ij}\big[D^{2}v\big]$
$\geq-\left|\frac{p}{n}\tilde{g}^{\frac{p}{n}-\frac{1}{2}}\frac{\tilde{\eta}^{l}\cdot\nabla\tilde{g}}{\sqrt{\tilde{g}}}\right|-C\sum_{i=1}^{n}\left|F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\right|+(2A-C)\operatorname{tr}F^{ij}\big[D^{2}v\big],$ (7.35)
[…]
$\tilde{f}^{\frac{1}{n}}=\sum_{j=1}^{n}F^{jj}_{z}\big[D^{2}_{z}v\big]v_{z_{j}z_{j}}\geq\left|F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\right|\quad\text{for}\ i=1,2,\dots,n,$ (7.36)
at the point $y_{*}$ in the $z$-coordinates. Combining (7.35), (7.36), Lemma 2.1, (2.13), and (7.34), we derive at the point $y_{*}$,
$0$ $\geq-C_{6}\tilde{f}^{\frac{1}{4(n-1)}-\frac{1}{n(n-1)}}-C\tilde{f}^{\frac{1}{n}}+\frac{2A-C}{C_{7}}\sigma_{1}^{\frac{1}{n-1}}\big[D^{2}v\big]\tilde{f}^{\frac{-1}{n(n-1)}}$
$\geq\frac{2A-C}{C_{7}}\tilde{f}^{\frac{-1}{n(n-1)}}-C_{6}\tilde{f}^{\frac{1}{4(n-1)}-\frac{1}{n(n-1)}}-C\tilde{f}^{\frac{1}{n}}$
$>0$
by taking $A$ large, which is a contradiction, where $C_{7}>0$ is universal, and $C_{6}$ depends on $\|\tilde{g}\|_{C^{1,1}(\overline{\omega})}$ . Thus $\Psi^{l}$ cannot attain its maximum in $\omega$ for $l=1,2,\dots,n-1$.
Second, we choose $B_{A}$ large, depending on $A$, so that
$\Psi^{l}\leq 0\quad\text{on}\ \partial\omega,\quad\forall 1\leq l\leq n-1.$
Fix $l\in\{1,2,\dots,n-1\}$, it follows from (7.24) that
$\Psi^{l}\leq\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A\big|y^{\prime}\big|^{2}-\frac{1}{2}B_{A}(v+y_{n}).$ (7.37)
Combining (7.20) and (7.21) yields
$v+y_{n}\geq\frac{1}{2}\big|y^{\prime}\big|^{2}+O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}\big|y^{\prime}\big|^{2}\quad\text{on}\ \partial_{1}\omega,$
which implies
$\Psi^{l}\leq A\big|y^{\prime}\big|^{2}-\frac{1}{8}B_{A}\big|y^{\prime}\big|^{2}\leq 0\quad\text{on}\ \partial_{1}\omega,$
by taking $B_{A}$ large. It follows from (7.25) that
$v+y_{n}\geq\min\left\{\frac{1}{4}\beta^{2},\varepsilon_{0}\beta^{2}\right\}=\varepsilon_{0}\beta^{2}\quad\text{on}\ \partial_{2}\omega\cup\partial_{3}\omega.$ (7.38)
Then using (7.37) and (7.38), we have
$\Psi^{l}\leq C+A-\frac{1}{2}B_{A}\varepsilon_{0}\beta^{2}\leq 0\quad\text{on}\ \partial_{2}\omega\cup\partial_{3}\omega,$
[…]
By the comparison principle,
$\Psi^{l}\leq 0\quad\text{in}\ \overline{\omega},\quad\forall 1\leq l\leq n-1.$
Since $\Psi^{l}(0)=0$, we get $\Psi^{l}_{n}(0)\leq 0$, which yields
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. For any point
$T^{-1}y_{0}=x_{0}\in\partial\Omega\cap\left\{\big|x^{\prime}\big|\leq r_{0}\right\}\cap\left\{|x|\leq\theta_{k}\sqrt{b_{k}}\right\},$
where $\theta_{k}>0$ is chosen small so that $|x_{0}|\leq\theta_{i}\sqrt{b_{i}}/2$ for all $k+1\leq i\leq n-1$, let $v(y_{0})=0$, $\nabla v(y_{0})=0$ by subtracting an affine function, and let $y_{0}$ be the origin, $\tilde{\nu}(y_{0})=e_{n}$, $\tilde{\eta}^{l}(y_{0})=e_{l}$ for all $1\leq l\leq n-1$ by translation and rotation. With similar arguments as above, we derive
$\left|v_{\tilde{\nu}\tilde{\eta}^{l}}(y_{0})\right|\leq C,\quad\forall 1\leq l\leq n-1,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. Pulling back to the $x$-coordinates, we obtain from (7.32) that
$C\geq M^{2}\left|\tilde{\nu}^{\prime}D^{2}_{y}u(y_{0})\tilde{\eta}^{l}\right|=M^{2}\left|\big(T^{-1}\tilde{\nu}\big)^{\prime}D^{2}_{x}u(x_{0})T^{-1}\tilde{\eta}^{l}\right|=M^{2}\left|u_{\nu\eta^{l}}(x_{0})\right|\frac{1}{|T\nu|}\frac{1}{M_{l}}$
for $l=1,2,\dots,n-1$, where $\tilde{\nu}=T\nu/|T\nu|$. Thus by (7.19), we have
$\left|u_{\nu\eta^{l}}(x_{0})\right|\leq C\frac{|T\nu|M_{l}}{M^{2}}\leq C\frac{1}{\sqrt{M}}=C\sqrt{b_{k}},\quad\forall 1\leq l\leq k,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. This completes the proof of Lemma 7.5. ∎
We emphasize here that for the Monge-Ampère equation, Theorem 1.1 also holds under the weaker conditions $\inf_{\Omega}f>0$ and $f^{3/(2n-2)}\in C^{1,1}(\overline{\Omega})$. Indeed, Lemma 3.1 follows directly from the convexity of the admissible solution, and Lemma 3.2 can be adapted by differentiating with respect to a tangential vector filed, with the same argument as in the proof of Lemma 7.3.
Theorem 7.6.
Let $u\in C^{3,1}(\overline{\Omega})$ be a convex solution of equation (1.1) with $k=n$, $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega}f>0$, and $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega})$. Then
$\sup_{\partial\Omega}|u_{\nu\nu}|\leq C,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
At the origin, we have
$u_{nn}(0)\prod_{i=1}^{n-1}b_{i}-\sum_{j=1}^{n-1}\bigg(\prod_{i\neq j}b_{i}\bigg)u_{nj}^{2}(0)=f(0).$
Lemma 7.3 ensures that Assumption 7.4 holds for k=n2k=n-2. Thus using Lemma 7.5, by induction we obtain
$u_{ni}(0)\leq C\sqrt{b_{i}}\quad\text{for}\ i=1,2,\dots,n-2.$ (7.39)
Then it follows from Lemma 7.2, Lemma 7.3, and (7.39) that
$0\leq u_{nn}(0)\leq\frac{1}{\prod_{i=1}^{n-1}b_{i}}f(0)+\sum_{j=1}^{n-1}\frac{1}{b_{j}}u_{nj}^{2}(0)\leq C,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. This completes the proof of Theorem 7.6. ∎
Combining Theorems 1.1, 1.2, and 7.6, we obtain Theorem 1.12.
8. Appendix
Proof of Lemma 6.6.
It is clear that the expression
$\left(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right)(x)$
is invariant under orthogonal transformations. The proof is divided into three steps.
Step 1. Let the origin $0\in\partial\Omega$, let $e_{n}=\nu(0)$, let $\partial\Omega$ be represented by $x_{n}=\tilde{\rho}(x^{\prime})$ near the origin, and let $\varphi(0)=0$ and $\nabla\varphi(0)=0$. Let $\tilde{\eta}(x)$ be an arbitrary vector field defined in $B_{r}$, where the radius $r>0$ can be arbitrarily small depending on the origin $0\in\partial\Omega$. Furthermore, let $\tilde{\eta}$ satisfy
$\tilde{\eta}(0)=e_{1}\quad\text{and}\quad\left|\partial_{x_{1}}(\tilde{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$
for some constant $C$.
Since $(-\tilde{\rho}_{1}(x^{\prime}),\dots,-\tilde{\rho}_{n-1}(x^{\prime}),1)$ is a normal vector at $x\in\partial\Omega$,
$\tilde{\eta}\big(x^{\prime},\tilde{\rho}\big(x^{\prime}\big)\big)\cdot\big(-\tilde{\rho}_{1}\big(x^{\prime}\big),\dots,-\tilde{\rho}_{n-1}\big(x^{\prime}\big),1\big)=0.$ (8.1)
Differentiating (8.1) with respect to $x_{1}$ at the origin yields
$\partial_{x_{1}}(\tilde{\eta}_{n})(0)-\tilde{\rho}_{11}(0)=0.$
Then we have
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)=u_{n11}(0)+u_{ni}(0)\partial_{x_{1}}(\tilde{\eta}_{i})(0)\leq u_{n11}(0)+u_{nn}(0)\tilde{\rho}_{11}(0)+K_{1},$
where $K_{1}$ depends on $C$ and second-order tangential-normal derivatives of $u$ on $\partial\Omega$. Define
$\tilde{\tau}=\tilde{\tau}(x):=\begin{pmatrix}1-\tilde{\rho}_{11}(0)x_{n}\\ -\tilde{\rho}_{12}(0)x_{n}\\ \vdots\\ -\tilde{\rho}_{1,n-1}(0)x_{n}\\ \sum_{i=1}^{n-1}\tilde{\rho}_{1i}(0)x_{i}\end{pmatrix}$.
It holds that
$\left(u_{(\tilde{\tau})(\tilde{\tau})}\right)_{n}(0)$ $=u_{n11}(0)+u_{ij}\partial_{x_{n}}\left[\tilde{\tau}_{i}\tilde{\tau}_{j}\right]+u_{in}(0)\tilde{\tau}_{j}\partial_{x_{j}}(\tilde{\tau}_{i})(0)+u_{i}\partial_{x_{n}}\left[\tilde{\tau}_{j}\partial_{x_{j}}(\tilde{\tau}_{i})\right]$
$\geq u_{n11}(0)+u_{nn}(0)\tilde{\rho}_{11}(0)-K_{2},$
where $K_{2}$ depends on $\Omega$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. As a conclusion,
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\left(u_{(\tilde{\tau})(\tilde{\tau})}\right)_{n}(0)+K_{1}+K_{2}\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$ (8.2)
by Lemma 6.5 and increasing $K_{\sigma}$.
Step 2. Return to the coordinate system (6.1). For any boundary point $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, we introduce a new coordinate system by $y=T(x-x_{0})$, where $T$ is an orthogonal matrix, such that
$\bar{\nu}(0)=e_{n}\quad\text{and}\quad\bar{\eta}(0)=\left|\eta(x_{0})\right|e_{1},$
where $\bar{\nu}$, $\bar{\eta}$ are defined by
$\bar{\nu}(y):=T\nu(x)=T\nu\big(T^{\prime}y+x_{0}\big),\quad\bar{\eta}(y):=T\eta(x)=T\eta\big(T^{\prime}y+x_{0}\big)$
with $\eta$ given in (6.3). It is clear that $\bar{\eta}$ is well-defined in $B_{|x_{0}|/2}$. Assume $\partial\Omega$ is represented by $y_{n}=\bar{\rho}(y^{\prime})$ near the origin in the new coordinates.
We claim that
$\left|\partial_{y_{1}}(\bar{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$ (8.3)
where $C$ depends only on $\Omega$ (the proof is deferred to Step 3). Define $\tilde{\eta}(y):=\bar{\eta}(y)/|\eta(x_{0})|$. Then
$\tilde{\eta}(0)=e_{1}\quad\text{and}\quad\left|\partial_{y_{1}}(\tilde{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$
because $1\leq|\eta(x_{0})|\leq 2$. By (8.2), we get
$\left(u_{nij}\bar{\eta}_{i}\bar{\eta}_{j}+u_{ni}\bar{\eta}_{j}\partial_{y_{j}}(\bar{\eta}_{i})\right)(0)=|\eta(x_{0})|^{2}u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\sigma M+4K_{\sigma/4},\quad\forall\sigma>0.$
Coming back to the original coordinate system (6.1), by increasing $K_{\sigma}$, we derive
$\left(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right)(x_{0})\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$
for all $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
Step 3: verification of (8.3). Compute
$\partial_{y_{1}}(\bar{\eta}_{i})(y)$ $=\sum_{j=1}^{n}\partial_{x_{j}}\left(\sum_{k=1}^{n}T_{ik}\eta_{k}(x)\right)\frac{\partial x_{j}}{\partial y_{1}}$
$=\sum_{j=1}^{n-1}\partial_{x_{j}}\left[T_{in}\sum_{k=1}^{n-1}\frac{x_{k}}{|x^{\prime}|}\rho_{k}\big(x^{\prime}\big)+\sum_{k=1}^{n-1}T_{ik}\frac{x_{k}}{|x^{\prime}|}\right]T^{\prime}_{j1}$
$=\sum_{j,k=1}^{n-1}T_{in}T^{\prime}_{j1}\left[\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+\rho_{k}\big(x^{\prime}\big)\left(\frac{\delta_{jk}}{|x^{\prime}|}-\frac{x_{k}x_{j}}{|x^{\prime}|^{3}}\right)\right]+T^{\prime}_{j1}T_{ik}\left(\frac{\delta_{jk}}{|x^{\prime}|}-\frac{x_{k}x_{j}}{|x^{\prime}|^{3}}\right)$
$=\sum_{j,k=1}^{n-1}T_{in}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+T_{in}\frac{1}{|x^{\prime}|}\left(\sum_{k=1}^{n-1}T_{1k}\rho_{k}\big(x^{\prime}\big)\right)$
$\quad\,+\sum_{j,k=1}^{n-1}T_{1j}T_{ik}\left(\frac{\delta_{jk}}{|x^{\prime}|}-\frac{x_{k}x_{j}}{|x^{\prime}|^{3}}\right)-\sum_{j=1}^{n-1}T_{1j}\frac{x_{j}}{|x^{\prime}|^{2}}\left(T_{in}\sum_{k=1}^{n-1}\rho_{k}\big(x^{\prime}\big)\frac{x_{k}}{|x^{\prime}|}\right)$ (8.4)
for all $1\leq i\leq n$. Since $T\nu(x_{0})=e_{n}$,
$T_{1n}+\sum_{k=1}^{n-1}T_{1k}\left(-\rho_{k}\big(x_{0}^{\prime}\big)\right)=0.$ (8.5)
It follows from (8) and (8.5) that
$\partial_{y_{1}}(\bar{\eta}_{i})(0)$ $=\sum_{j,k=1}^{n-1}T_{in}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+T_{in}T_{1n}\frac{1}{|x^{\prime}|}$
$\quad\,+\sum_{j,k=1}^{n-1}T_{1j}T_{ik}\left(\frac{\delta_{jk}}{|x^{\prime}|}-\frac{x_{k}x_{j}}{|x^{\prime}|^{3}}\right)-\sum_{j=1}^{n-1}T_{1j}\frac{x_{j}}{|x^{\prime}|^{2}}\left(T_{in}\sum_{k=1}^{n-1}\rho_{k}\big(x^{\prime}\big)\frac{x_{k}}{|x^{\prime}|}\right)$ (8.6)
for all $1\leq i\leq n$, where $x^{\prime}$ is evaluated at $x_{0}^{\prime}$. Since $T\eta(x_{0})=|\eta(x_{0})|e_{1}$, we have
$T_{in}\sum_{k=1}^{n-1}\rho_{k}\big(x^{\prime}\big)\frac{x_{k}}{|x^{\prime}|}+\sum_{k=1}^{n-1}T_{ik}\frac{x_{k}}{|x^{\prime}|}=0,\quad\forall 2\leq i\leq n,$ (8.7)
and
$T_{1n}\sum_{k=1}^{n-1}\rho_{k}\big(x^{\prime}\big)\frac{x_{k}}{|x^{\prime}|}+\sum_{k=1}^{n-1}T_{1k}\frac{x_{k}}{|x^{\prime}|}=|\eta(x_{0})|,$ (8.8)
where $x^{\prime}$ is evaluated at $x_{0}^{\prime}$.
Combining (8) and (8.7) gives that for any $2\leq i\leq n$,
$\partial_{y_{1}}(\bar{\eta}_{i})(0)$ $=\sum_{j,k=1}^{n-1}T_{in}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+T_{in}T_{1n}\frac{1}{|x^{\prime}|}+\sum_{j,k=1}^{n-1}T_{1j}T_{ik}\frac{\delta_{jk}}{|x^{\prime}|}$
$=\sum_{j,k=1}^{n-1}T_{in}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+\frac{1}{|x^{\prime}|}\sum_{k=1}^{n}T_{ik}T_{1k}$
$=\sum_{j,k=1}^{n-1}T_{in}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big),$ (8.9)
where $x^{\prime}$ is evaluated at $x_{0}^{\prime}$. The last equality in (8) is due to the fact that $T$ is an orthogonal matrix. By virtue of (8) and (8.8), one derives
$\partial_{y_{1}}(\bar{\eta}_{1})(0)$ $=\sum_{j,k=1}^{n-1}T_{1n}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+\frac{1}{|x^{\prime}|}-|\eta(x_{0})|\sum_{j=1}^{n-1}T_{1j}\frac{x_{j}}{|x^{\prime}|^{2}}$
$=\sum_{j,k=1}^{n-1}T_{1n}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+\frac{1}{|x^{\prime}|}+\frac{|\eta(x_{0})|}{|x^{\prime}|}\left(T_{1n}\sum_{k=1}^{n-1}\rho_{k}\big(x^{\prime}\big)\frac{x_{k}}{|x^{\prime}|}-|\eta(x_{0})|\right)$
$=\sum_{j,k=1}^{n-1}T_{1n}T_{1j}\frac{x_{k}}{|x^{\prime}|}\rho_{kj}\big(x^{\prime}\big)+\frac{1-|\eta(x_{0})|^{2}}{|x^{\prime}|}+|\eta(x_{0})|T_{1n}\sum_{k=1}^{n-1}\frac{x_{k}}{|x^{\prime}|}\frac{\rho_{k}\big(x^{\prime}\big)}{|x^{\prime}|},$ (8.10)
where $x^{\prime}$ is evaluated at $x_{0}^{\prime}$. Since $\nabla_{x^{\prime}}\rho(0)=0$, we obtain from (8) and (8) that
$\left|\partial_{y_{1}}(\bar{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n,$
where $C$ depends only on $\Omega$. The proof of Lemma 6.6 is complete. ∎
Lemma 8.1.
There exists a constant $C>0$ such that
$\sup_{x\in\partial\Omega}\left|u_{\eta_{1}\eta_{2}\eta_{3}}(x)\right|\leq C,$
where $\eta_{1}(x)$, $\eta_{2}(x)$, $\eta_{3}(x)$ are arbitrary unit tangential vectors of $\partial\Omega$ at $x$; and $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Let the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, and $\partial\Omega$ be represented by $x_{n}=\rho(x^{\prime})$ near the origin. Differentiating the boundary condition $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ with respect to $x_{i}$, $x_{j}$, and $x_{k}$, where $1\leq i,\ j,\ k\leq n-1$, we compute
$(u-\varphi)_{i}+\rho_{i}(u-\varphi)_{n}=0,$
$(\partial_{j}+\rho_{j}\partial_{n})(u-\varphi)_{i}+\rho_{ij}(u-\varphi)_{n}+\rho_{i}(\partial_{j}+\rho_{j}\partial_{n})(u-\varphi)_{n}=0,$
$\partial_{k}\Big((\partial_{j}+\rho_{j}\partial_{n})(u-\varphi)_{i}+\rho_{ij}(u-\varphi)_{n}+\rho_{i}(\partial_{j}+\rho_{j}\partial_{n})(u-\varphi)_{n}\Big)\Big|_{0}=0.$
Therefore, at the origin
$(u-\varphi)_{ijk}+\rho_{jk}(u-\varphi)_{in}+\rho_{ijk}(u-\varphi)_{n}+\rho_{ij}(u-\varphi)_{nk}+\rho_{ik}(u-\varphi)_{nj}=0.$
Combining this and Theorem 1.1 completes the proof of Lemma 8.1. ∎