### v2

# On the Dirichlet problem for the degenerate $k$-Hessian equation

Abstract.
This paper investigates the existence of a global $C^{1,1}$ solution to the Dirichlet problem for the $k$-Hessian equation with a nonnegative right-hand side $f$, focusing on the required conditions for $f$. The conditions $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, together with $f\geq 0$ in a domain $\Omega_{0}\Supset\Omega$, are optimal, as demonstrated by classical counterexamples. For the Monge-Ampère equation ($k=n$), we establish the existence under the optimal condition $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ together with $f\geq 0$ in $\Omega_{0}$. For the general $k$-Hessian equation, we establish the existence under the condition $f\geq 0$ in $\Omega_{0}$ together with one of the following three conditions:
$(i)\quad f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}}),\ \ \inf_{\Omega}\Delta u\geq 1,\ \ 2\leq k\leq n-1;$
$(ii)\quad f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ \inf_{\Omega}\Delta u\geq 1,\ \ 5\leq k\leq n-1;$
$(iii)\quad f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ 2\leq k\leq n-1.$
Key words and phrases:
Monge-Ampère equation, $k$-Hessian equation, Dirichlet problem, Existence theory, Optimal degeneracy.
2020 Mathematics Subject Classification:
35J60, 35J70, 35J25, 35J96.
1. Introduction
In this paper, we consider the Dirichlet problem for the degenerate $k$-Hessian equation
$\begin{cases}\sigma_{k}\big[D^{2}u\big]=f\geq 0\quad\text{in}\ \Omega,\\ u=\varphi\quad\text{on}\ \partial\Omega,\end{cases}$ (1.1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$, $2\leq k\leq n$, and $D^{2}u$ denotes the Hessian matrix of the function $u$. For an $n\times n$ real symmetric matrix $A$, we denote by $\sigma_{k}[A]$ the $k$-th elementary symmetric polynomial of the eigenvalues $\lambda(A)=(\lambda_{1},\dots,\lambda_{n})$; that is,
$\sigma_{k}[A]=\sigma_{k}(\lambda(A)):=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}$
for $k=1,2,\dots,n$. Equivalently, $\sigma_{k}[A]$ is the sum of all $k\times k$ principal minors of $A$.
Following the work of Caffarelli-Nirenberg-Spruck [2], a function $u\in C^{2}(\Omega)\cap C^{0}(\overline{\Omega})$ is called $k$-admissible if its Hessian eigenvalues satisfy
$\lambda\big(D^{2}u(x)\big)\in\Gamma_{k},\quad\forall x\in\Omega,$
where the Gårding cone is defined by
$\Gamma_{k}:=\left\{\lambda\in\mathbb{R}^{n}:\ \sigma_{j}(\lambda)>0\ \text{for all}\ 1\leq j\leq k\right\}.$
They proved that
$\frac{\partial\sigma_{k}}{\partial\lambda_{i}}>0\quad\text{in}\ \Gamma_{k},\quad\forall 1\leq i\leq n;$
that $\sigma_{k}^{1/k}(\lambda)$ is concave in $\Gamma_{k}$; and that the operator $\sigma_{k}^{1/k}[A]$ is elliptic and concave for any real symmetric matrix $A$ with $\lambda(A)\in\Gamma_{k}$. A hypersurface in $\mathbb{R}^{n}$ is called $k$-convex (for $k\in\{1,2,\dots,n-1\}$) if its principal curvatures $\kappa=(\kappa_{1},\kappa_{2},\dots,\kappa_{n-1})$ satisfy $\sigma_{j}(\kappa)\geq 0$ for all $j\leq k$ everywhere. It is strictly $k$-convex if there exists a constant $\delta>0$ such that $\sigma_{j}(\kappa)\geq\delta$ uniformly on the hypersurface for all $j\leq k$. As shown in [2], a necessary condition for the existence of a $C^{2}(\overline{\Omega})$ solution to (1.1) is that the boundary $\partial\Omega$ be strictly (k1)(k-1)-convex when regarded as a hypersurface in $\mathbb{R}^{n}$.
The $k$-Hessian equation is called degenerate if the nonnegative function $f$ is allowed to vanish at some points in $\overline{\Omega}$, and non-degenerate if $\inf_{\Omega}f>\delta$ for some positive constant $\delta$. This paper focuses on the degenerate case. The condition $f^{1/k}\in C^{1,1}(\overline{\Omega})$ is natural when exploiting the concavity of the operator. However, the condition $f^{1/k}\in C^{1,1}$ may not be optimal. Regarding the optimal condition, a counterexample for the Monge-Ampère equation was given by Wang [32]:
$\begin{cases}\det D^{2}u=\bar{f}(x):=\eta\left(\frac{x_{n}}{|x^{\prime}|^{\alpha}}\right)\left|x^{\prime}\right|^{\beta}\quad\text{in}\ B_{1},\\ u=0\quad\text{on}\ \partial B_{1},\end{cases}$ (1.2)
where $\alpha>1$, $\beta>0$, and
$\eta(t):=\begin{cases}e^{-1/(1-t^{2})},\quad|t|<1,\\ 0,\quad|t|\geq 1.\end{cases}$
Wang [32] proved that if $\beta<2(n-1)(\alpha-1)$, then (1.2) admits no $C^{1,1}$ solution. Taking $\beta=2(n-1)(\alpha-1)-1$, one has $\bar{f}\geq 0$ in $B_{2}$, and
$\bar{f}^{\frac{1}{n-1-\frac{2n-1}{2\alpha}}}\in C^{1,1}(\overline{B_{2}}),\quad\bar{f}^{\frac{1}{n-1}}\in C^{1,1-\frac{2n-1}{\alpha(n-1)}}(\overline{B_{2}}),\quad\forall\alpha\geq 4,$
and
$\bar{f}^{\frac{1}{\frac{2(n-1)}{3}-\frac{2n-1}{3\alpha}}}\in C^{2,1}(\overline{B_{2}}),\quad\bar{f}^{\frac{3}{2(n-1)}}\in C^{2,1-\frac{6n-3}{2\alpha(n-1)}}(\overline{B_{2}}),\quad\forall\alpha\geq 6.$
Ivochkina-Trudinger-Wang [16] pointed out that a modification of the counterexample (1.2) also applies to the $k$-Hessian equation for $2\leq k\leq n-1$, although no proof was provided. A detailed proof of this fact was given by Dinew-Pliś-Zhang [5]; namely, the problem
$\begin{cases}\sigma_{k}\big[D^{2}u\big]=\bar{f}\quad\text{in}\ B_{1},\\ u=0\quad\text{on}\ \partial B_{1},\end{cases}$
has no $C^{1,1}$ solution whenever $\beta<2(k-1)(\alpha-1)$, for all $2\leq k\leq n-1$. Consequently, for any $k\in\{2,3,\dots,n\}$,
$f^{\frac{1}{k-1}}\in C^{1,1}\quad\text{and}\quad f^{\frac{3}{2(k-1)}}\in C^{2,1}$
are optimal.
For the Monge-Ampère equation, Guan [9] reduced the estimate of second-order derivatives to the boundary under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, and obtained the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ under the homogeneous boundary condition ($\varphi=0$). Subsequently, by introducing a completely new method, Guan-Trudinger-Wang [12] established a global $C^{2}$ estimate for the Monge-Ampère equation under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, which in turn yielded the existence of a convex solution in $C^{1,1}(\overline{\Omega})$. For the case $2\leq k\leq n-1$ of equation (1.1), Ivochkina-Trudinger-Wang [16] proposed the following open problem:
$\text{existence of a global $C^{1,1}$ solution under $f^{1/(k-1)}\in C^{1,1}$}.$ (1.3)
Dong [6] solved problem (1.3) under the homogeneous boundary condition. For general boundary conditions, Jiao-Wang [17] derived a global $C^{2}$ estimate for convex solutions to problem (1.3). However, since smooth approximations of convex viscosity solutions do not necessarily preserve convexity, they [17] were unable to establish the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for equation (1.1). There is an extensive literature on degenerate $k$-Hessian equations; see, for instance, [3, 15, 18, 19, 28, 29, 30, 11, 8, 13, 31, 25, 4], although this list is far from exhaustive.
A common strategy is to derive global $C^{2}$ a priori estimate for the non-degenerate case that are independent of the positive lower bound of $f$. Then the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for the degenerate case is established by an approximation argument using solutions to non-degenerate problems. To establish global $C^{2}$ a priori estimate independent of the positive lower bound of $f$, the primary challenge is to derive the second-order normal derivative estimate on the boundary.
For the Monge-Ampère equation, using the convexity of the admissible solutions and of the domain, as well as the affine invariance of the equation, Guan-Trudinger-Wang [12] introduced a method to estimate the second-order normal derivatives on the boundary. Recently, Jiao-Wang [17] extended this method to convex solutions of the $k$-Hessian equation in uniformly convex domains. Applying the method of Guan-Trudinger-Wang [12], we establish the existence of a global $C^{1,1}$ solution for the Monge-Ampère equation under the optimal condition.
Theorem 1.1.
Let $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique convex solution in $C^{1,1}(\overline{\Omega})$ for (1.1) with $k=n$.
For general $k$-Hessian equations, we employ the method of Ivochkina-Trudinger-Wang [16], which is completely different from that of [12], to establish the boundary second-order normal derivative estimate. Ivochkina-Trudinger-Wang [16] provided a PDE-based proof of Krylov’s results [21, 22, 23, 24], and, roughly speaking, their ideas are similar. These results and ideas were clarified very well in [16], and we quote them below.
A global upper bound, independent of the positive lower bound of $f$, for the second-order derivatives of admissible solutions to the Dirichlet problem of Hessian equations, and more general Bellman equations, was obtained by Krylov [20], by a probabilistic argument, and by an analytic proof in a series of papers Krylov [21, 22, 23, 24]. Roughly speaking, Krylov’s proof [21, 22, 23, 24] consists of two steps. One is the weakly interior estimate, that is for any positive constants $\varepsilon,\ \delta>0$, there exists a constant $C_{\varepsilon,\delta}$, depending on $n$, $\varphi,\ \partial\Omega,\ \|u\|_{C^{1}(\overline{\Omega})},\ \|f\|_{C^{1,1}(\overline{\Omega})}$, and in particular the upper bound of the boundary second-order tangential and tangential-normal derivatives of $u$, such that
$\sup_{\Omega_{\delta}}\left|D^{2}u\right|\leq\varepsilon\sup_{\partial\Omega}\left|D^{2}u\right|+C_{\varepsilon,\delta},$ (1.4)
where $\Omega_{\delta}:=\{x\in\Omega:\ \operatorname{dist}(x,\partial\Omega)>\delta\}$. The other one is the boundary estimate in terms of the interior one, that is for any $\delta>0$, there exists a constant $C_{\delta}$, depending in addition on $C_{\varepsilon_{0},\delta}$ in (1.4), with arbitrarily given $\varepsilon_{0}>0$, such that
$\sup_{\partial\Omega}\left|D^{2}u\right|\leq C_{\delta}\left(1+\sup_{\partial\Omega_{\delta}}\left|D^{2}u\right|\right).$ (1.5)
Firstly, we establish the weakly interior estimate for degenerate $k$-Hessian equations.
Theorem 1.2.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Then, for any constant $\varepsilon>0$, there exists a constant $C_{\varepsilon}>0$ such that
$\left|D^{2}u(x)\right|\leq\frac{1}{d(x)}\left(\varepsilon\sup_{\partial\Omega}\left|D^{2}u\right|+C_{\varepsilon}\right),\quad\forall x\in\Omega,$ (1.6)
where $d(x):=\operatorname{dist}(x,\partial\Omega)$, and the constant $C_{\varepsilon}$ depends on $\varepsilon$, $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Secondly, the boundary estimate in terms of the interior one for degenerate $k$-Hessian equations is also obtained, but under an additional a priori condition: there exists a small positive constants $r_{0}$ such that
$\Delta u\geq 1\quad\text{in}\ \Omega\cap B_{r_{0}}(y),$ (1.7)
for all $y\in\partial\Omega$ at which $\Delta u(y)\gg 1$. Then the boundary second-order normal derivative estimate is established.
Theorem 1.3.
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, and $f\geq 0$ in $\Omega_{0}$ where $\Omega\Subset\Omega_{0}$. Let one of the following three conditions hold:
[…]
$(iii)\quad f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ 2\leq k\leq n-1.$
Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Remark 1.4 (On condition $(i)$).
If condition (1.7) in $(i)$ could be removed, then the longstanding open problem (1.3) would be resolved. When $\sup_{\partial\Omega}\Delta u$ is bounded from above, the $C^{2}$ estimate follows readily. We hope that Theorem 1.3 may provide some insight into the resolution of the open problem (1.3). Moreover, condition (1.7) is automatically satisfied if $f$ admits a positive lower bound near $\partial\Omega$.
Remark 1.5 (On condition $(iii)$).
As pointed out by Guan-Trudinger-Wang [12], for applications to problems in differential geometry, it is desirable to impose no restriction on the nonnegative function $f$ apart from smoothness. Guan [9] established the existence of a global $C^{1,1}$ solution for the Monge-Ampère equation with homogeneous boundary condition in dimension $n=3$, assuming $0\leq f\in C^{3,1}(\mathbb{R}^{3})$. In comparison with the condition $f^{1/(k-1)}\in C^{1,1}$, it is interesting that the exponent of $f$ can be improved when higher regularity of $f$ is available, for $k\geq 4$.
Remark 1.6.
There is no implication between the conditions $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{1/(k-q)}\in C^{2,1}(\overline{\Omega_{0}})$ for any $0<q<k$. This can be seen from the following two counterexamples. One counterexample is $f^{1/(k-q)}:=|x|^{2}\in C^{\infty}(\overline{\Omega_{0}})$ with $0\in\partial\Omega$. It is clear that $f^{1/k}=|x|^{2(k-q)/k}\notin C^{1,1}(\overline{\Omega_{0}})$. The other counterexample is
$f^{1/k}:=\left(1+g(x)\right)e^{-1/|x-x_{0}|}$
in $\overline{\Omega_{0}}$, where 0 and $x_{0}$ are two distinct points on $\partial\Omega$, and
$g(x):=\begin{cases}|x|^{4}\left(1+\sin(1/|x|)\right),\quad x\neq 0,\\ 0,\quad x=0.\end{cases}$
One verifies that $g\in W^{2,\infty}$, hence $g\in C^{1,1}(\overline{\Omega_{0}})$, while $D^{2}g$ is not continuous at 0. Thus $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$, but $D^{2}(f^{1/(k-q)})$ fails to be continuous at 0.
The rest of the paper is organized as follows. Section 2 presents preliminary results. Section 3 establishes the global $C^{1}$ estimate and the boundary second-order tangential and tangential-normal estimates, and reduces the estimate of second-order derivatives to the boundary. Section 4 deals with the Monge-Ampère equation under the optimal condition, proving Theorem 1.1. Section 5 is devoted to the weakly interior estimate. Section 6 develops the boundary estimate in terms of the interior one to finish Theorem 1.3.
Throughout the paper, the dependence of constants on the dimension $n$ and the number $k$ in (1.1) is omitted, and a constant depending only on $n$ and $k$ is called universal. This section ends with a list of notations.
Notation.
1. $e_{i}=(0,\dots,1,\dots,0)$ denotes the $i$-th standard coordinate vector.
2. $x^{\prime}=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$.
3. $B_{r}(x_{0})=\{x\in\mathbb{R}^{n}:\ |x-x_{0}|\leq r\}$ and $x_{0}$ is omitted when $x_{0}=0$.
4. $\operatorname{dist}(E,F)=\text{distance from $E$ to $F$}$, $\forall E,\ F\subset\mathbb{R}^{n}$.
5. $\mathbb{S}^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
6. $\mathbf{I}_{n}$ denotes the $n\times n$ identity matrix.
7. $A^{\prime}=A^{T}$ denotes the transpose of the matrix $A$.
8. $A^{1/2}$ denotes the square root of positive semi-definite matrix $A$.
9. $\nabla_{x^{\prime}}=(\partial_{1},\partial_{2},\dots,\partial_{n-1})$.
10. $D^{2}_{x^{\prime}x^{\prime}}$ denotes a matrix with elements $\partial_{ij}$ where $1\leq i,\ j\leq n-1$.
11. $\nu=\nu(x)$ denotes the unit interior normal vector of $\partial\Omega$ at $x\in\partial\Omega$.
2. Preliminaries
The following pointwise inequalities are fundamental in this paper.
Lemma 2.1.
Let $g$ be a function satisfying $g>0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. We have
$(i)$ If $g\in C^{1,1}(\overline{\Omega_{0}})$, then
$\frac{\left|\nabla g(x)\right|^{2}}{g(x)}\leq K,\quad\forall x\in\overline{\Omega},$ (2.1)
where $K$ depends on $\|g\|_{C^{1,1}(\overline{\Omega_{0}})}$ and $\operatorname{dist}(\Omega,\partial\Omega_{0})$, but is independent of $\inf_{\Omega_{0}}g$.
$(ii)$ If $g\in C^{2,1}(\overline{\Omega_{0}})$, then for any $\alpha<1/2$,
$\partial_{ee}g(x)-\alpha\frac{\left|\partial_{e}g(x)\right|^{2}}{g(x)}\geq-Kg^{\frac{1}{3}}(x),\quad\forall x\in\overline{\Omega},\ e\in\mathbb{S}^{n-1},$ (2.2)
where $K$ depends on $\alpha$, $\|g\|_{C^{2,1}(\overline{\Omega_{0}})}$, and $\operatorname{dist}(\Omega,\partial\Omega_{0})$, but is independent of $\inf_{\Omega_{0}}g$.
Proof.
In fact, we only need to consider the case of one dimension. Let $h(x)$ be a function in $C^{2,1}(\mathbb{R})$ and $h>0$ on $\mathbb{R}$. Denote $\beta:=h(0)>0$.
Step 1: we prove conclusion $(i)$.
It holds that
$\begin{cases}0\leq h\big(\sqrt{\beta}\big)=\beta+h^{\prime}(0)\sqrt{\beta}+O(\beta)\\ 0\leq h\big(-\sqrt{\beta}\big)=\beta-h^{\prime}(0)\sqrt{\beta}+O(\beta).\end{cases}$
Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq 1+K,$
where $K=\|h\|_{C^{1,1}(\mathbb{R})}$. By virtue of the above, (2.1) are valid at points where $(g(x))^{1/2}<\operatorname{dist}(\Omega,\partial\Omega_{0})$. And (2.1) hold clearly at points where $(g(x))^{1/2}\geq\operatorname{dist}(\Omega,\partial\Omega_{0})$.
Step 2: we prove conclusion $(ii)$.
It holds that
$0\leq h\big(\beta^{t_{1}}\big)+h\big(-\beta^{t_{1}}\big)=2\beta+h^{\prime\prime}(0)\beta^{2t_{1}}+O\big(\beta^{3t_{1}}\big),$
where $t_{1}>0$ is to be determined. Thus
$h^{\prime\prime}(0)\geq-2\beta^{1-2t_{1}}+O\big(\beta^{t_{1}}\big).$
Taking $t_{1}=1/3$ yields
$h^{\prime\prime}(0)\geq-C\beta^{\frac{1}{3}},$ (2.3)
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$.
It holds that
$\begin{cases}0\leq h\big(\beta^{t_{2}}\big)=\beta+h^{\prime}(0)\beta^{t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{2t_{2}}+O\big(\beta^{3t_{2}}\big)\\ 0\leq h\big(-\beta^{t_{2}}\big)=\beta-h^{\prime}(0)\beta^{t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{2t_{2}}+O\big(\beta^{3t_{2}}\big),\end{cases}$
where $t_{2}>0$ is to be determined. Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\beta^{\frac{1}{2}-t_{2}}+\frac{1}{2}h^{\prime\prime}(0)\beta^{t_{2}-\frac{1}{2}}+O\big(\beta^{2t_{2}-\frac{1}{2}}\big).$
Taking $t_{2}=1/3$ yields
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\frac{1}{2}h^{\prime\prime}(0)\beta^{-\frac{1}{6}}+C\beta^{\frac{1}{6}},$ (2.4)
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$.
Case 1: $h^{\prime\prime}(0)\leq K_{1}\beta^{1/3}$, where $K_{1}>0$ is to be determined. It follows from (2.4) that
$\frac{\left|h^{\prime}(0)\right|^{2}}{\beta}\leq\left(K_{1}/2+C\right)^{2}\beta^{\frac{1}{3}}.$
Combining this inequality and (2.3), we get
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq-\left(C+\alpha(K_{1}/2+C)^{2}\right)h^{\frac{1}{3}}(0)\quad\text{in Case 1}.$
Case 2: $h^{\prime\prime}(0)>K_{1}\beta^{1/3}$. It holds that
$\begin{cases}0\leq h\big(t_{3}\sqrt{\beta}\big)=\beta+h^{\prime}(0)t_{3}\sqrt{\beta}+\frac{1}{2}h^{\prime\prime}(0)\left(t_{3}\sqrt{\beta}\right)^{2}+O\big(\big(t_{3}\sqrt{\beta}\big)^{3}\big)\\ 0\leq h\big(-t_{3}\sqrt{\beta}\big)=\beta-h^{\prime}(0)t_{3}\sqrt{\beta}+\frac{1}{2}h^{\prime\prime}(0)\left(t_{3}\sqrt{\beta}\right)^{2}+O\big(\big(t_{3}\sqrt{\beta}\big)^{3}\big),\end{cases}$
where $t_{3}>0$ is to be determined. Thus
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}\leq\frac{1}{t_{3}}+\frac{1}{2}h^{\prime\prime}(0)t_{3}+Ct_{3}^{2}\beta^{\frac{1}{2}},$
where $C$ depends only on $\|h\|_{C^{2,1}(\mathbb{R})}$. Taking $t_{3}=\sqrt{2}\left(h^{\prime\prime}(0)\right)^{-1/2}$, which implies
$t_{3}\sqrt{\beta}=\sqrt{2}\left(h^{\prime\prime}(0)\right)^{-\frac{1}{2}}\sqrt{\beta}<\sqrt{2}K_{1}^{-\frac{1}{2}}\beta^{\frac{1}{3}},$
we have
$\frac{\left|h^{\prime}(0)\right|}{\sqrt{\beta}}$ $\leq\sqrt{2}\sqrt{h^{\prime\prime}(0)}+2C\left(h^{\prime\prime}(0)\right)^{-\frac{3}{2}}\beta^{\frac{1}{2}}\sqrt{h^{\prime\prime}(0)}$
$<\left(\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\right)\sqrt{h^{\prime\prime}(0)}.$
Given that $\alpha<1/2$, choose $K_{1}$ sufficiently large such that
$\sqrt{2}+2CK_{1}^{-\frac{3}{2}}\leq\frac{1}{\sqrt{\alpha}}.$
It follows that
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq 0\quad\text{in Case 2}.$
Combining Case 1 and Case 2 gives
$h^{\prime\prime}(0)-\alpha\frac{\left|h^{\prime}(0)\right|^{2}}{h(0)}\geq-Kh^{\frac{1}{3}}(0),$
where $K$ depends on $\|h\|_{C^{2,1}(\mathbb{R})}$ and $\alpha$. By virtue of the above, (2.2) are valid at points in $\overline{\Omega}$ satisfying
$\max\left\{1,\sqrt{2}K_{1}^{-1/2}\right\}g^{\frac{1}{3}}(x)<\operatorname{dist}(\Omega,\partial\Omega_{0}).$
And (2.2) hold clearly at other points in $\overline{\Omega}$. The proof of Lemma 2.1 is complete. ∎
Remark 2.2.
In general, it is impossible to weaken the condition $g>0$ in $\Omega_{0}$ to $g>0$ in $\Omega$. A counterexample was provided by Błocki [1]: the function $g(x):=x$ on $[0,1]$ is smooth but $\sqrt{g}\notin C^{0,1}([0,1])$. However, the condition $g>0$ in $\Omega$ suffices to estimate the derivatives of $g$ in directions tangential to $\partial\Omega$.
Remark 2.3.
In general, it is impossible to improve the exponent $\alpha$ to any $\alpha>1/2$. Consider the family of functions $g_{\beta}(x):=(x+\beta^{1/2})^{2}$, where $\beta>0$ is a parameter. It is clear that $g_{\beta}>0$ and $g_{\beta}\in C^{\infty}(\mathbb{R})$. Computing (2.2) at $x=0$ and letting $\beta\rightarrow 0$ yield that $\alpha\leq 1/2$.
By convention, we set $\sigma_{0}(\lambda)\equiv 1$. Some fundamental properties of $\sigma_{k}$ are listed below, and one can refer to [27, 14, 26, 33].
Proposition 2.4.
For any $k\in\{2,3,\dots,n\}$ and any symmetric matrix $A$ with $\lambda(A)\in\Gamma_{k}$, the following hold:
$\sigma_{k}^{\frac{1}{k}}[A]\leq C_{1}\sigma_{k-1}^{\frac{1}{k-1}}[A];$ (2.5)
$\sigma_{k-1}[A]\geq C_{2}\sigma_{1}^{\frac{1}{k-1}}[A]\sigma_{k}^{\frac{k-2}{k-1}}[A];$ (2.6)
$\sum_{i=1}^{n}\sigma_{k}^{ii}[A]=(n-k+1)\sigma_{k-1}[A],$ (2.7)
where $C_{1}>0$ and $C_{2}>0$ are universal.
The equation (1.1) is equivalent to
$F\big[D^{2}u\big]=g^{\frac{p}{k}}\quad\text{in}\ \Omega,$ (2.8)
where
$F[A]:=\sigma_{k}^{\frac{1}{k}}[A]\quad\text{and}\quad g:=f^{\frac{1}{p}},$
with
$p:=\begin{cases}k-1,\quad\text{if}\ f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}}),\\ (2k-2)/3,\quad\text{if}\ f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}}).\end{cases}$ (2.9)
Define $F^{ij}[A]:=\partial F/\partial A_{ij}$ and $F^{ij,st}[A]:=\partial^{2}F/\partial A_{ij}\partial A_{st}$. Differentiating equation (2.8) in the direction $e\in\mathbb{S}^{n-1}$ gives
$F^{ij}\big[D^{2}u\big]u_{eij}=\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{e}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}}$ (2.10)
for all $e\in\mathbb{S}^{n-1}$. A second differentiation combined with the concavity of $F$ yields
$\partial_{ee}\left(g^{\frac{p}{k}}\right)=F^{ij}\big[D^{2}u\big]u_{eeij}+F^{ij,st}\big[D^{2}u\big]u_{eij}u_{est}\leq F^{ij}\big[D^{2}u\big]u_{eeij},$
which implies
$F^{ij}\big[D^{2}u\big]u_{eeij}\geq\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{ee}g-\frac{k-p}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)$ (2.11)
for all $e\in\mathbb{S}^{n-1}$.
Define a matrix-valued function $G$ with entries
$G^{ij}(x):=\frac{F^{ij}\big[D^{2}u(x)\big]}{\operatorname{tr}F^{ij}\big[D^{2}u(x)\big]},\quad\forall x\in\Omega.$ (2.12)
Lemma 2.5.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), and $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$ hold. Then
$\operatorname{tr}F^{ij}\big[D^{2}u\big]\geq\frac{1}{C}\max\left\{1,\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}\right\};$ (2.13)
$0\leq\operatorname{tr}\left(GD^{2}u\right)\leq K;$ (2.14)
$\left|\operatorname{tr}\left(GD^{2}u_{\xi}\right)\right|\leq K\left|\xi\right|\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall\xi\in\mathbb{R}^{n},$ (2.15)
in $\Omega$, where $C>0$ is universal, and $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Step 1. It follows from (2.7) that
$\operatorname{tr}F^{ij}\big[D^{2}u\big]$ $=\frac{1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sum_{i=1}^{n}\sigma_{k}^{ii}\big[D^{2}u\big]$
$=\frac{n-k+1}{k}\sigma_{k}^{\frac{1-k}{k}}\big[D^{2}u\big]\sigma_{k-1}\big[D^{2}u\big].$
By (2.5) and (2.6), we get (2.13).
Step 2. It follows from (2.7) and the Euler theorem for homogeneous functions that
$\operatorname{tr}\left(GD^{2}u\right)=\frac{1}{n-k+1}\frac{\sigma_{k}^{ij}\big[D^{2}u\big]u_{ij}}{\sigma_{k-1}\big[D^{2}u\big]}=\frac{k}{n-k+1}\frac{\sigma_{k}\big[D^{2}u\big]}{\sigma_{k-1}\big[D^{2}u\big]}\geq 0.$
By (2.5) and (2.8), one derives
$\operatorname{tr}\left(GD^{2}u\right)\leq C\sigma_{k}^{\frac{1}{k}}\big[D^{2}u\big]=Cf^{\frac{1}{k}}\leq K,$
where $C$ is universal, and $K$ depends on $\|f\|_{C^{0}(\overline{\Omega})}$.
Step 3. The combination of (2.13) and (2.5) yields that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\left|F^{ij}\big[D^{2}u\big]u_{eij}\right|$
for all $e\in\mathbb{S}^{n-1}$. Then it follows from (2.10), (2.9), and (2.1) that
$\left|\operatorname{tr}\left(GD^{2}u_{e}\right)\right|\leq C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\left|\nabla g\right|}{\sqrt{g}}\leq K\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
for all $e\in\mathbb{S}^{n-1}$, where the constant $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Thus (2.15) is valid. ∎
Lemma 2.6.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), and $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Then
$\operatorname{tr}\left(GD^{2}u_{\xi\xi}\right)\geq-K\left|\xi\right|^{2}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],\quad\forall\xi\in\mathbb{R}^{n},$ (2.16)
in $\Omega$, where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
When $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it follows from (2.11), (2.13), and (2.1) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)$ $\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}\frac{k-1}{k}g^{-\frac{1}{k}}\left|\partial_{ee}g-\frac{1}{k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right|$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]$
for all $e\in\mathbb{S}^{n-1}$, where $C$ is universal, and $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. When $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, which implies $(k+2)/3k<1/2$, it follows from (2.11) and (2.2) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)$ $\geq\frac{1}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}\frac{p}{k}g^{-\frac{2}{3k}-\frac{1}{3}}\left(\partial_{ee}g-\frac{k+2}{3k}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)$
$\geq-\frac{1}{\operatorname{tr}F^{ij}\big[D^{2}u\big]}K_{2}g^{-\frac{2}{3k}}$
for all $e\in\mathbb{S}^{n-1}$, where $K_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Combining this and (2.13) yields
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq-C\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big]g^{\frac{p}{k(k-1)}}K_{2}g^{-\frac{2}{3k}}=-CK_{2}\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u\big],$
where $C$ is universal. Therefore (2.16) holds. ∎
Below, we construct an auxiliary function which depends only on $\Omega$.
Lemma 2.7.
Let $\Omega\subset\mathbb{R}^{n}$ be a domain of class $C^{3,1}$ with $\partial\Omega$ strictly (k1)(k-1)-convex. Then there exists a function $\psi\in C^{3,1}(\overline{\Omega})$ satisfying
$\begin{cases}\operatorname{tr}\left(GD^{2}\psi\right)\leq-1&\quad\text{in}\ \Omega,\\ \psi>0&\quad\text{in}\ \Omega,\\ \psi=0&\quad\text{on}\ \partial\Omega,\\ \frac{D\psi(x)}{|D\psi(x)|}=\nu(x)&\quad\text{on}\ \partial\Omega,\\ |D\psi|\geq 1&\quad\text{on}\ \partial\Omega.\end{cases}$ (2.17)
Proof.
Let $\psi^{*}$ be the function constructed in Section 2 of Caffarelli-Nirenberg-Spruck [2], which satisfies
$\begin{cases}\lambda\big(D^{2}\psi^{*}(x)\big)\in\Gamma_{k},&\quad\forall x\in\overline{\Omega},\\ \psi^{*}<0&\quad\text{in}\ \Omega,\\ \psi^{*}=0&\quad\text{on}\ \partial\Omega,\\ \frac{D\psi^{*}(x)}{|D\psi^{*}(x)|}=-\nu(x)&\quad\text{on}\ \partial\Omega,\\ \left|D\psi^{*}\right|\geq\varepsilon_{0}&\quad\text{on}\ \partial\Omega,\end{cases}$
for some constant $\varepsilon_{0}>0$ depending on $\Omega$. By the compactness of $\overline{\Omega}$, there exists a small constant $\varepsilon_{1}>0$ such that
$\left\{\lambda\big(D^{2}\psi^{*}(x)\big):\ x\in\overline{\Omega}\right\}-\varepsilon_{1}\lambda(\mathbf{I}_{n})\Subset\Gamma_{k}.$
Using the concavity of $F$, we derive
$F^{ij}\big[D^{2}u\big]\left(K\big(\psi^{*}_{ij}-\varepsilon_{1}\delta_{ij}\big)-u_{ij}\right)$ $\geq F\big[K\big(D^{2}\psi^{*}-\varepsilon_{1}\mathbf{I}_{n}\big)\big]-F\big[D^{2}u\big]$
$=K\sigma_{k}^{\frac{1}{k}}\big[D^{2}\psi^{*}-\varepsilon_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}$
$>0,$
where $K$ is sufficiently large. Consequently, by (2.14), we have
$F^{ij}\big[D^{2}u\big]\psi^{*}_{ij}\geq\varepsilon_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]+\frac{1}{K}F^{ij}\big[D^{2}u\big]u_{ij}\geq\varepsilon_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big].$
Thus
$\operatorname{tr}\left(GD^{2}\psi^{*}\right)\geq\varepsilon_{1}.$
Choosing $\psi=-t\psi^{*}$ with $t>0$ sufficiently large, we obtain (2.17). ∎
3. Reduction to the boundary second-order normal estimate
It follows from Caffarelli-Nirenberg-Spruck [2] that
$\sup_{\overline{\Omega}}|u|+\sup_{\partial\Omega}|Du|\leq C,$ (3.1)
where $C$ depends on $\Omega$, $\|f\|_{C^{0}(\overline{\Omega})}$, and $\|\varphi\|_{C^{3}(\partial\Omega)}$, but is independent of $\inf_{\Omega_{0}}f$.
Lemma 3.1.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\Omega}|Du|\leq\sup_{\partial\Omega}|Du|+C,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Without loss of generality, let $\Omega\subset\mathbb{R}^{n}\setminus B_{1}$. Suppose
$\sup_{x\in\overline{\Omega},e\in\mathbb{S}^{n-1}}\left(u_{e}+K|x|^{2}\right)$
is attained at an interior point $x_{0}\in\Omega$ in the direction $e=e_{1}$, where $K>0$ is a constant to be determined. Then at the point $x_{0}$,
$u_{1j}+2Kx_{j}=0,\quad\forall 1\leq j\leq n,$
and
$F^{ij}\big[D^{2}u\big]\left(u_{1ij}+2K\delta_{ij}\right)\leq 0.$ (3.2)
Since $\lambda(D^{2}u)\in\Gamma_{2}$, we derive at the point $x_{0}$,
$0\leq\sigma_{1}^{2}\big[D^{2}u\big]-\sum_{j=1}^{n}u_{1j}^{2}\leq\sigma_{1}^{2}\big[D^{2}u\big]-4K^{2}.$ (3.3)
By (2.10) and (2.1), we obtain
$F^{ij}\big[D^{2}u\big]u_{1ij}=\frac{p}{k}\frac{\partial_{1}g}{\sqrt{g}}f^{\frac{1}{4(k-1)}-\frac{1}{k(k-1)}}\geq-K_{1}f^{\frac{-1}{k(k-1)}}$ (3.4)
in $\Omega$, where $K_{1}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
It follows from (3.2), (3.4), (2.13), and (3.3) that at the point $x_{0}$,
$0$ $\geq\frac{2K}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}-K_{1}f^{\frac{-1}{k(k-1)}}$
$\geq\left(\frac{2K}{C}(2K)^{\frac{1}{k-1}}-K_{1}\right)f^{\frac{-1}{k(k-1)}},$
where $C>0$ is universal. Taking $K$ large enough, we get a contradiction. The proof of Lemma 3.1 is complete. ∎
Lemma 3.2.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{1,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\partial\Omega}\left(|u_{\nu\eta}|+|u_{\eta\eta}|\right)\leq C,$
where $\eta$ is any unit tangential vector; and $C$ depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Let the origin be a boundary point satisfying $e_{n}=\nu(0)$, and $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)$ (3.5)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. It is clear that $\nabla_{x^{\prime}}\rho(0)=0$. Let $\varphi(0)=0$ and $\nabla\varphi(0)=0$ by subtracting an affine function. Differentiating $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ on $\partial\Omega$ twice, one gets
$u_{ij}(0)=\varphi_{ij}(0)-u_{n}(0)\rho_{ij}(0),\quad\forall 1\leq i,\ j\leq n-1.$
Thus the second-order tangential estimate is obtained.
For any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, define an operator
$T_{\alpha}:=\partial_{\alpha}+\sum_{\beta=1}^{n-1}\rho_{\alpha\beta}(0)(x_{\beta}\partial_{n}-x_{n}\partial_{\beta}),\quad\forall 1\leq\alpha\leq n-1.$
It is standard that
$T_{\alpha}(u-\varphi)=O\Big(\big|x^{\prime}\big|^{2}\Big),\quad\forall 1\leq\alpha\leq n-1,$ (3.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. It follows from (2.10) and (2.1) that
$\left|F^{ij}[D^{2}u]\left(T_{\alpha}(u-\varphi)\right)_{ij}\right|$ $\leq\left|T_{\alpha}(g^{p})\right|+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$
$\leq C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]$ (3.7)
in $\Omega$, where $C_{1}$ depends on $\|\rho\|_{C^{3,1}}$, $\|\varphi\|_{C^{3}(\partial\Omega)}$; $C_{2}$ depends on $\|\rho\|_{C^{3,1}}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Define
$\omega_{r}:=\left\{x\in\Omega:\ \rho\big(x^{\prime}\big)<x_{n}<\rho\big(x^{\prime}\big)+r^{4},\ \big|x^{\prime}\big|<r\right\}$
for all $r\leq r_{0}$. Define a function
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{0}},$
where positive constants $\beta_{1}$ and $\beta_{2}$ are to be determined. It holds that
$D^{2}v=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]-\beta_{2}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big].$
Since $\partial\Omega\in C^{3,1}$ is strictly (k1)(k-1)-convex, one can fix $\beta_{1}$ sufficiently small such that $\lambda(D^{2}[x_{n}^{2}+\beta_{1}\rho](0))$ belongs to a compact subset of $\Gamma_{k}$ independent of $0\in\partial\Omega$. By
$|x|\rightarrow 0,\ \rho\big(x^{\prime}\big)\rightarrow 0,\ \nabla_{x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow 0,\ D^{2}_{x^{\prime}x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow D^{2}_{x^{\prime}x^{\prime}}\rho(0)\quad\text{as}\ r\rightarrow 0,$
there exist small positive constants $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\beta_{2}$, and $\delta_{1}$ such that
$\left\{\lambda\big(D^{2}v-\delta_{1}\mathbf{I}_{n}\big):\ x\in\omega_{r_{1}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big),$
and is thus contained in a compact subset of $\Gamma_{k}$. We emphasize that $\beta_{1}$, $r_{1}$, $\beta_{2}$, $\delta_{1}$ all depend only on $\Omega$. Using the concavity of $F$,
$F^{ij}\big[D^{2}u\big]\left(C_{3}(v_{ij}-\delta_{1}\delta_{ij})-u_{ij}\right)\geq C_{3}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{1}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}>0$
in $\omega_{r_{1}}$, where $C_{3}$ is sufficiently large. Then by virtue of (2.14) and (2.13), we derive
$F^{ij}\big[D^{2}u\big]v_{ij}\geq\delta_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\geq\frac{\delta_{1}}{C}\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \omega_{r_{1}},$ (3.8)
where $C>0$ is universal.
We claim that
$\pm T_{\alpha}(u-\varphi)\leq-Kv\quad\text{in}\ \overline{\omega_{r_{1}}},\quad\forall 1\leq\alpha\leq n-1,$ (3.9)
where the constant $K>0$ is to be determined.
Step 1: verification of (3.9) on $\partial\omega_{r_{1}}$. The boundary $\partial\omega_{r_{1}}$ consists of three parts: $\partial_{1}\omega_{r_{1}}\cup\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$, where $\partial_{1}\omega_{r_{1}}$ and $\partial_{2}\omega_{r_{1}}$ are respectively the graph parts of $\rho$ and $\rho+r_{1}^{4}$, and $\partial_{3}\omega_{r_{1}}$ is the boundary part on $\{|x^{\prime}|=r_{1}\}$. On $\partial_{1}\omega_{r_{1}}$, using (3.6) one gets
$Kv=K\beta_{2}\big|x^{\prime}\big|^{2}\geq\pm T_{\alpha}(u-\varphi)$
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$. It is clear that
$\pm T_{\alpha}(u-\varphi)\leq C_{4},$
where $C_{4}$ depends on $\|\rho\|_{C^{2}}$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$. It holds that
$Kv\geq K\left(\beta_{1}r_{1}^{4}-r_{1}^{8}\right)\geq\frac{\beta_{1}}{2}Kr_{1}^{4}\quad\text{on}\ \partial_{2}\omega_{r_{1}},$
and
$Kv\geq K\beta_{2}r_{1}^{4}\quad\text{on}\ \partial_{3}\omega_{r_{1}}.$
Then we obtain $Kv\geq\pm T_{\alpha}(u-\varphi)$ on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$ by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{1}(\partial\Omega)}$, and $\|u\|_{C^{1}(\overline{\Omega})}$.
Step 2. Suppose, for contradiction, that (3.9) does not hold. Then the maximum
$\max_{\overline{\omega_{r_{1}}}}\left(\pm T_{\alpha}(u-\varphi)+Kv\right)$
is attained at an interior point $x_{0}\in\omega_{r_{1}}$. At the point $x_{0}$, we have
$0=\left(\pm T_{\alpha}(u-\varphi)+Kv\right)_{n}=\left(\pm T_{\alpha}u\right)_{n}-\left[-Kv_{n}+\left(\pm T_{\alpha}\varphi\right)_{n}\right]$ (3.10)
and
$F^{ij}\big[D^{2}u\big]\left(\pm T_{\alpha}(u-\varphi)+Kv\right)_{ij}\leq 0.$ (3.11)
It follows from (3.10) that
$\left|\left(T_{\alpha}u\right)_{n}\right|\geq-Kv_{n}-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$ $=K\left(\beta_{1}-2x_{n}+2\rho\big(x^{\prime}\big)\right)-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
$\geq\frac{\beta_{1}}{2}K-\left|\left(T_{\alpha}\varphi\right)_{n}\right|$
at the point $x_{0}$. By taking $K$ large, which depends on $\Omega$ and $\|\varphi\|_{C^{2}(\partial\Omega)}$, we get
$\big|D^{2}u(x_{0})\big|\geq 1,$
which together with $\lambda(D^{2}u)\in\Gamma_{2}$ yields
$\sigma_{1}\big[D^{2}u(x_{0})\big]\geq\big|D^{2}u(x_{0})\big|\geq 1.$ (3.12)
Using (3.11), (3), (3.8), and (3.12), we obtain at the point $x_{0}$,
$\left(C_{2}f^{\frac{-1}{k(k-1)}}+C_{1}\operatorname{tr}F^{ij}\big[D^{2}u\big]\right)+K\left(\frac{\delta_{1}}{2}\operatorname{tr}F^{ij}\big[D^{2}u\big]+\frac{\delta_{1}}{2C}f^{\frac{-1}{k(k-1)}}\right)\leq 0,$
which leads to a contradiction by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Therefore, claim (3.9) holds with $K$ depending on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows immediately that
$\pm\left(T_{\alpha}(u-\varphi)\right)_{n}(0)\leq-Kv_{n}(0)=K\beta_{1}$
for all $1\leq\alpha\leq n-1$. The proof of Lemma 3.2 is complete. ∎
Inspired by Guan [9], Guan-Trudinger-Wang [12], and Jiao-Wang [17], we reduce the estimate of second-order derivatives to the boundary.
Lemma 3.3.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\sup_{\Omega}\left|D^{2}u\right|\leq\sup_{\partial\Omega}\left|D^{2}u\right|+C,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Define a function
$v:=\Delta u+|x|^{2}.$
Suppose $v$ attains its maximum at an interior point $x_{0}\in\Omega$, for otherwise we are done. Then at the point $x_{0}$, we have
$\partial_{x_{i}}(\Delta u)+2x_{i}=0,\quad\forall 1\leq i\leq n,$ (3.13)
and
$0\geq\sigma_{k}^{ij}\big[D^{2}u\big]v_{ij},$
which together with (2.7) and (2.6) implies
$0\geq\sigma_{k}^{ij}\big[D^{2}u\big](\Delta u)_{ij}+C\sigma_{1}^{\frac{1}{k-1}}\big[D^{2}u\big]\sigma_{k}^{\frac{k-2}{k-1}}\big[D^{2}u\big],$ (3.14)
where $C>0$ is universal. Differentiating equation (1.1), we derive
$\sigma_{k}^{ij}\big[D^{2}u\big](\Delta u)_{ij}+\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}=\Delta(g^{p}).$
Combining this equality with (3.14), we obtain at the point $x_{0}$,
$0$ $\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[(p-1)\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]$ (3.15)
$\hskip 12.80373pt-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}.$
It follows from Lemma 3.2 of Guan-Li-Li [10] that computing
$0\geq\left(\left(\frac{\sigma_{k}}{\sigma_{1}}\right)^{1/(k-1)}\right)^{ij,st}u_{ije}u_{ste}$
yields
$\frac{\sigma_{k}^{ij,st}u_{ije}u_{ste}}{\sigma_{k}}$ $\leq\frac{k-2}{k-1}\left(\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\right)^{2}$
$\hskip 12.80373pt+\frac{2}{k-1}\frac{\sigma_{k}^{ij}u_{ije}}{\sigma_{k}}\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}-\frac{k}{k-1}\left(\frac{\sigma_{1}^{ij}u_{ije}}{\sigma_{1}}\right)^{2}$
for all $e\in\mathbb{S}^{n-1}$, where operators take values at arbitrary matrix $A$ with $\lambda(A)\in\Gamma_{k}$. Thus by (3.13), we infer at the point $x_{0}$,
$-\sum_{l=1}^{n}\sigma_{k}^{ij,st}\big[D^{2}u\big]u_{ijl}u_{stl}$ $\geq-\sum_{l=1}^{n}\frac{k-2}{k-1}\frac{p^{2}g^{p-1}|\partial_{x_{l}}g|^{2}}{g}$ (3.16)
$\hskip 12.80373pt-\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{pg^{p-1}\partial_{x_{l}}g2x_{l}}{\Delta u}-\frac{kf}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right].$
Using (3.15), (3.16), (2.2), and (2.1), we get at the point $x_{0}$,
$0$ $\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}+pg^{p-1}\sum_{l=1}^{n}\left[-\frac{1}{3}\frac{|\partial_{x_{l}}g|^{2}}{g}+\partial_{x_{l}x_{l}}g\right]$
$\hskip 12.80373pt-\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{pg^{p-1}\partial_{x_{l}}g2x_{l}}{\Delta u}-\frac{kf}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right]$
$\geq C(\Delta u)^{\frac{1}{k-1}}f^{1-\frac{1}{k-1}}-K_{1}f^{1-\frac{1}{k-1}}$
$\hskip 12.80373pt-f^{1-\frac{1}{k-1}}\sum_{l=1}^{n}\left[\frac{2}{k-1}\frac{K_{2}f^{\frac{1}{4(k-1)}}2|x_{l}|}{\Delta u}-\frac{kf^{\frac{1}{k-1}}}{k-1}\left(\frac{2x_{l}}{\Delta u}\right)^{2}\right],$
where $K_{1}$ and $K_{2}$ depend on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Therefore,
$\Delta u(x_{0})\leq K,$
where $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. The proof of Lemma 3.3 is complete. ∎
Combining (3.1), Lemmas 3.1, 3.2 and 3.3, we reduce the $C^{2}$ estimate to the boundary second-order normal derivative estimate under the condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$. Under the homogeneous boundary condition, Dong [6] established the required boundary second-order normal derivative estimate. Thus the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$ is obtained. Then the existence is established by an approximation argument.
Theorem 3.4.
Let $\partial\Omega\in C^{3}$ be strictly (k1)(k-1)-convex, $\varphi=0$, $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
We emphasize here that results corresponding to Lemmas 3.1, 3.2, 3.3, and Theorem 3.4 were established by Jiao-Wang [17] under the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega})$ with $f\geq 0$ in $\Omega$.
4. Monge-Ampère equation
In this section, we assume
$f^{\frac{3}{2n-2}}\in C^{2,1}(\overline{\Omega}),\quad f\geq 0\ \text{in}\ \Omega,$
and apply the method of Guan-Trudinger-Wang [12] to establish the boundary second-order normal derivative estimate. We emphasize here that for the Monge-Ampère equation, Lemmas 3.1 and 3.2 hold under the weaker condition $\inf_{\Omega}f>0$ and $f^{3/(2n-2)}\in C^{1,1}(\overline{\Omega})$. Indeed, Lemma 3.1 follows directly from the convexity of the admissible solution, and Lemma 3.2 can be adapted by differentiating with respect to a tangential vector filed, with the same argument as in the proof of later Lemma 4.3.
Let the origin be a boundary point satisfying $\nu(0)=e_{n}$. By an affine transformation of $x^{\prime}$, let $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)=\frac{1}{2}\big|x^{\prime}\big|^{2}+\text{cubic of $x^{\prime}$}+O\Big(\big|x^{\prime}\big|^{4}\Big)$ (4.1)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. Then rotate $x^{\prime}$ such that $D^{2}_{x^{\prime}x^{\prime}}u(0)$ is a diagonal matrix with
$0<u_{11}(0)\leq u_{22}(0)\leq\dots\leq u_{n-1,n-1}(0).$
By subtracting an affine function, let
$u(0)=0\quad\text{and}\quad\nabla u(0)=0,$
which yields by the convexity of $u$,
$u(x)\geq 0,\quad\forall x\in\overline{\Omega}.$
Differentiating the boundary condition $u(x^{\prime},\rho(x^{\prime}))=\varphi(x^{\prime},\rho(x^{\prime}))$ yields
$D^{2}_{x^{\prime}x^{\prime}}u(0)=D^{2}_{x^{\prime}x^{\prime}}\varphi(0)+D^{2}_{x^{\prime}x^{\prime}}\rho(0)\partial_{n}\varphi(0).$
For convenience, denote
$b_{i}:=u_{ii}(0),\quad\forall 1\leq i\leq n-1.$
By Taylor expansion with respect to $x^{\prime}$ at the origin, we have
$\varphi\big(x^{\prime},\rho\big(x^{\prime}\big)\big)=\frac{1}{2}\sum_{i=1}^{n-1}b_{i}x_{i}^{2}+\text{cubic of $x^{\prime}$}+O\Big(\big|x^{\prime}\big|^{4}\Big)$ (4.2)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
The following lemma is the inequality (3.5) in Guan-Trudinger-Wang [12].
Lemma 4.1 ([12]).
Let $h\in C^{3,1}(\mathbb{R})$ be a nonnegative function on $\mathbb{R}$ satisfying
$h(t)=\alpha t^{2}+\beta t^{3}+R(t),\quad\forall t\in\mathbb{R},$
where $\alpha\geq 0$, $\beta\in\mathbb{R}$, and the remainder term satisfies $|R(t)|\leq\gamma|t|^{4}$ for some $\gamma>0$. Then
$|\beta|\leq(1+\gamma)\sqrt{\alpha}.$
Proof.
Lemma 4.1 is obtained by combining two facts $h(\sqrt{\alpha})\geq 0$ and $h(-\sqrt{\alpha})\geq 0$. ∎
4.1. Lower bounds for boundary second-order tangential derivatives
Lemma 4.2.
There exists a constant $\gamma_{0}>0$ such that
$\prod_{i=1}^{n-1}b_{i}\geq\gamma_{0}f(0),$
where $\gamma_{0}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
Step 1. Define
$M:=\frac{1}{b_{1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (4.3)
where
$M_{i}:=\sqrt{b_{i}}M\ \big(\geq\sqrt{M}\big)\ \ \text{for}\ i=1,2,\dots,n-1,\quad\text{and}\quad M_{n}:=M.$
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=\frac{f(T^{-1}y)}{\prod_{i=1}^{n-1}b_{i}}=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
It follows from (4.1) and (4.3) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}y_{1}^{2}+\frac{1}{2}\sum_{i=2}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (4.4)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$d_{i}\leq 1\ \ \text{for}\ i=2,\dots,n-1,\quad\big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\quad\big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (4.2) and (4.3) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}\big|y^{\prime}\big|^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (4.5)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 4.1, we conclude that the coefficients of the third-order terms in (4.5) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Step 2. Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1,\ |y_{i}|<1,\ \,\forall 2\leq i\leq n-1\right\}.$
We analyze the geometry of $\omega$. Clearly, the set $\omega$ is convex. Looking at $\partial\omega\cap\{y_{n}=1\}$, by (4.4) we have
$1\geq\frac{1}{2}y_{1}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|y_{1}|=2$ must be outside of $\omega$. The connectedness of $\omega$ yields $|y_{1}|\leq 2$, so the set $\omega$ is contained in the ball $B_{\sqrt{n+3}}$. It is trivial that
$y_{\star}:=\left(0,\frac{1}{4n},\dots,\frac{1}{4n},\frac{1}{2}\right)\in\omega,$ (4.6)
and
[…]
Then, since $\tilde{\rho}(y^{\prime})\leq|y^{\prime}|^{2}$, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where
$r_{\star}:=\min\left\{\frac{1}{2},1-\frac{1}{4n},\min\left\{\frac{1}{4n},\frac{1}{2}-\frac{1}{4n}\right\}\right\}.$ (4.7)
Using the convexity of $v$, we derive that the maximum of $v$ in $\overline{\omega}$ is attained on $\partial\omega\cap\partial\widetilde{\Omega}$. Thus, by (4.5) we obtain
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$ (4.8)
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Step 3. We claim that there exists a constant $K>0$, independent of $M$, such that for any ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$,
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\tilde{f}<K.$ (4.9)
Suppose, for contradiction, that (4.9) does not hold. Then
$\det D^{2}v\geq K\quad\text{in}\ B_{\frac{r_{\star}}{4n}}(y_{*}).$
Since (4.8) the function
$\frac{K^{1/n}}{2}\left((y-y_{*})^{2}-\frac{r_{\star}}{4n}\right)+C_{2}$
is an upper barrier of $v$ in $B_{\frac{r_{\star}}{4n}}(y_{*})$. Thus
$v(y_{*})\leq-K^{\frac{1}{n}}\frac{r_{\star}}{8n}+C_{2}<0$
by taking $K$ large, which depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$. But this contradicts with $v\geq 0$. Therefore, we obtain claim (4.9).
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
We have
$\left\|D_{y}^{3}\tilde{g}\right\|_{L^{\infty}(\widetilde{\Omega})}$ $\leq\left(\prod_{i=1}^{n-1}b_{i}\right)^{\frac{-3}{2n-2}}b_{1}^{\frac{3}{2}}\left\|D_{x}^{3}\left(f^{\frac{3}{2n-2}}\right)\right\|_{L^{\infty}(\Omega)}$
$\leq\left\|f^{\frac{3}{2n-2}}\right\|_{C^{2,1}(\overline{\Omega})}\leq C.$ (4.10)
From (4.9), there is a point $\tilde{y}\in B_{\frac{r_{\star}}{4n}}(y_{\star})$ satisfying $\tilde{f}(\tilde{y})\leq K$. The fact
$0\leq\tilde{g}\Big(\tilde{y}+\frac{r_{\star}}{2}e\Big)+\tilde{g}\Big(\tilde{y}-\frac{r_{\star}}{2}e\Big)\leq 2K^{\frac{3}{2(n-1)}}+\partial_{ee}\tilde{g}(\tilde{y})\left(\frac{r_{\star}}{2}\right)^{2}+O\big(r_{\star}^{3}\big)$
yields
$\partial_{ee}\tilde{g}(\tilde{y})\geq-\frac{C}{r_{\star}^{2}}\left(K^{\frac{3}{2(n-1)}}+1\right),\quad\forall e\in\mathbb{S}^{n-1}.$ (4.11)
Combining (4.1) and (4.11) yields that there exists a constant $K_{1}$ such that
$D^{2}\tilde{g}+K_{1}\mathbf{I}_{n}\quad\text{is positive definite in $\omega$},$
that is $\tilde{g}$ is semi-convex in $\omega$. Define
$\bar{g}:=\tilde{g}+\frac{K_{1}}{2}|y|^{2}\quad\text{in}\ \overline{\omega}.$
Then $\bar{g}$ is convex in $\omega$, and there exists a constant $\overline{K}>0$ such that
$\inf_{B_{\frac{r_{\star}}{4n}}(y_{*})}\bar{g}<\overline{K}$ (4.12)
for any ball $B_{\frac{r_{\star}}{4n}}(y_{*})$ contained in $B_{r_{\star}}(y_{\star})$. By (4.12) and the convexity of $\bar{g}$, it follows (see Figure 1) that
$|D\bar{g}|\leq\frac{2\overline{K}}{r_{\star}/4}\quad\text{in}\ B_{\frac{r_{\star}}{4}}(y_{\star}),$
and then by the mean value theorem,
$\partial_{ee}\bar{g}(y_{e})\leq\frac{4\overline{K}}{r_{\star}/4}\frac{1}{r_{\star}/4},\quad\forall e\in\mathbb{S}^{n-1},$
for some point $y_{e}\in\omega$ depending on $e$. Combining this and (4.1), we obtain
$\|\tilde{g}\|_{C^{2,1}(\overline{\omega})}\leq K_{2},$
where $K_{2}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
Figure 1. Use the convexity of $\bar{g}$ to get a uniform bound on its first-order derivatives in $B_{\frac{r_{\star}}{4}}(y_{\star})$, and further a uniform bound on its second-order derivatives at some points is obtained by the mean value theorem.
Therefore,
$K_{2}^{\frac{2(n-1)}{3}}\geq\tilde{f}(0)=\frac{f(0)}{\prod_{i=1}^{n-1}b_{i}}.$
This completes the proof of Lemma 4.2. ∎
4.2. Lower bounds for boundary tangential-normal derivatives
Lemma 4.3.
For any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\frac{1}{2}\sqrt{b_{n-1}}$, we have
$\left|u_{\nu\eta}(x)\right|\leq C\sqrt{b_{n-1}},$
where $\eta$ is any unit tangential vector to $\partial\Omega$ at $x$, and $C>0$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
Define
$M:=\frac{1}{b_{n-1}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y^{\prime}=\sqrt{M}x^{\prime}\quad\text{and}\quad y_{n}=Mx_{n}.$ (4.13)
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=M^{n-1}f\big(T^{-1}y\big)=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
It follows from (4.1) and (4.13) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}\big|y^{\prime}\big|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (4.14)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\quad\big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (4.2) and (4.13) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}\sum_{i=1}^{n-1}\tilde{b}_{i}y_{i}^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (4.15)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\tilde{b}_{i}\leq 1\ \ \text{for}\ i=1,2,\dots,n-1,\quad\text{and}\quad\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 4.1, we conclude that the coefficients of the third-order terms in (4.15) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<1\right\}.$
Looking at $\partial\omega\cap\{y_{n}=1\}$, by (4.14) we have
$1=\frac{1}{2}\big|y^{\prime}\big|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|y^{\prime}|=2$ must be outside of $\omega$. So the set $\omega$ is contained in the ball $B_{2}$ by the convexity of $\omega$. Similar to Step 2 in the proof of Lemma 4.2, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where $y_{\star}$ and $r_{\star}$ are given in (4.6) and (4.7), respectively. Furthermore, using the convexity of $v$ and (4.15), we derive
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$.
Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega}$.
By the same argument as in Step 3 of the proof of Lemma 4.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq K,$
where $K$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Note that $\partial\widetilde{\Omega}$ is uniformly convex on $\partial\widetilde{\Omega}\cap\partial\omega$ by (4.14). The first-order normal estimate on the boundary follows from Theorem 17.21 of Gilbarg-Trudinger [7] that
$\left|v_{\tilde{\nu}}(y)\right|\leq C_{3},\quad\forall y\in\partial\widetilde{\Omega}\cap\left\{y_{n}\leq 7/8\right\},$ (4.16)
where $\tilde{\nu}=\tilde{\nu}(y)$ denotes the unit interior normal vector to $\partial\widetilde{\Omega}$ at $y$, and $C_{3}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Using the convexity of $v$ and the uniform convexity of $\partial\widetilde{\Omega}$, it follows from (4.16) that
$|Dv|\leq C_{4}\quad\text{in}\ \widetilde{\Omega}\cap\left\{y_{n}\leq 3/4\right\},$
where $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. In view of Remark 2.2, to weaken the condition to $\inf_{\Omega}\tilde{g}>0$ when applying Lemma 2.1, we adapt the proof of Lemma 3.2 by differentiating with respect to a tangential vector filed to establish the boundary tangential-normal estimate. The convexity of $v$ and the uniform convexity of $\partial\widetilde{\Omega}$ play key roles in this argument. Define tangential vector fields
$\tilde{\eta}^{l}(y):=e_{l}+\tilde{\rho}_{l}\big(y^{\prime}\big)e_{n}\quad\text{for}\ l=1,2,\dots,n-1,$
in $\overline{\omega}$. Consider functions
$\Psi^{l}(y):=\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A|y|^{2}-B_{A}y_{n}\quad\text{in}\ \overline{\omega},$
for all $l=1,2,\dots,n-1$, where positive constants $A$ and $B_{A}$ are to be determined. We can obtain
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
and similarly
$\left|v_{\tilde{\nu}\tilde{\eta}^{l}}(y)\right|\leq C,\quad\forall y\in\partial\widetilde{\Omega}\cap\left\{y_{n}\leq 1/2\right\},\quad\forall 1\leq l\leq n-1.$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. Pulling back to the $x$-coordinates, we finish the proof of Lemma 4.3. Since a more complicated version of these arguments will be needed later, in the absence of uniform convexity of $\partial\widetilde{\Omega}$, we omit the details here. ∎
We proceed from Lemma 4.3 by induction to obtain a refinement. Define
$\eta^{l}=\eta^{l}(x):=e_{l}+\rho_{l}\big(x^{\prime}\big)e_{n},\quad\forall x\in\partial\Omega\cap\left\{\big|x^{\prime}\big|\leq r_{0}\right\},$ (4.17)
for all $l=1,2,\dots,n-1$. In fact, $\eta^{l}(x)$ is the tangential direction of $\partial\Omega$, which lies in the two-dimensional plane parallel to the $x_{l}$- and $x_{n}$-axes and passes through the point $x$. Our induction hypothesis is the following.
Assumption 4.4.
For some fixed $k\in\{1,2,\dots,n-2\}$ and each $i=k+1,\dots,n-1$, there exists a constant $\theta_{i}>0$ such that for any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\theta_{i}\sqrt{b_{i}}$, we have
$\left|u_{\nu\eta^{j}}(x)\right|\leq C\sqrt{b_{i}},\quad\forall 1\leq j\leq i,$
where constants $\theta_{i}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Lemma 4.5.
Let Assumption 4.4 hold for some fixed $k\in\{1,2,\dots,n-2\}$. Then there exists a constant $\theta_{k}>0$ such that for any $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ satisfying $|x|\leq\theta_{k}\sqrt{b_{k}}$, we have
$\left|u_{\nu\eta^{j}}(x)\right|\leq C\sqrt{b_{k}},\quad\forall 1\leq j\leq k,$ (4.18)
where constants $\theta_{k}$ and $C$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$, but are independent of $\inf_{\Omega}f$.
Proof.
Define
$M:=\frac{1}{b_{k}}.$
Make a dilation $x\rightarrow y=T(x)$ defined by
$y_{i}:=M_{i}x_{i}\quad\text{for}\ i=1,2,\dots,n,$ (4.19)
where
$M_{i}:=\begin{cases}\sqrt{M},\quad i=1,\dots,k,\\ \sqrt{b_{i}}M\ (\geq\sqrt{M}),\quad i=k+1,\dots,n-1,\\ M,\quad i=n.\end{cases}$
Without loss of generality, we assume $M\gg 1$, which depends on the $C^{3,1}$-norm of $\partial\Omega$. Thus, the transformation $T$ can be regarded as a blow-up transformation near the origin. Define
$v(y):=M^{2}u(x).$
Then
$\det D^{2}v=\frac{M^{k}f(T^{-1}y)}{\prod_{i=k+1}^{n-1}b_{i}}=:\tilde{f}(y)\quad\text{in}\ \widetilde{\Omega}:=T(\Omega).$
Denote
$\hat{y}:=(y_{1},\dots,y_{k})\quad\text{and}\quad\tilde{y}:=(y_{k+1},\dots,y_{n-1}).$
It follows from (4.1) and (4.19) that $\partial\widetilde{\Omega}$ is represented by
$y_{n}=\tilde{\rho}\big(y^{\prime}\big)=\frac{1}{2}|\hat{y}|^{2}+\frac{1}{2}\sum_{i=k+1}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)$ (4.20)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$d_{i}=\frac{b_{k}}{b_{i}}\ (\leq 1)\ \ \forall k+1\leq i\leq n-1,\ \ \big|D^{3}\tilde{\rho}\big|\leq\frac{C_{1}}{\sqrt{M}},\ \ \big|D^{4}\tilde{\rho}\big|\leq\frac{C_{1}}{M},$
and $C_{1}$ depends only on $\partial\Omega$. Combining (4.2) and (4.19) yields
$v=\widetilde{\varphi}\big(y^{\prime},\tilde{\rho}\big(y^{\prime}\big)\big)=\frac{1}{2}|\tilde{y}|^{2}+\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}y_{i}^{2}+\text{cubic of $y^{\prime}$}+O\Big(\big|y^{\prime}\big|^{4}\Big)$ (4.21)
on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, where
$\tilde{b}_{i}=\frac{b_{i}}{b_{k}}\ (\leq 1)\ \ \text{for}\ i=1,2,\dots,k,\quad\text{and}\quad\big|D_{y}^{4}\widetilde{\varphi}\big|\leq\big|D_{x}^{4}\varphi\big|\leq C.$
By applying Lemma 4.1, we conclude that the coefficients of the third-order terms in (4.21) are also bounded. Hence both $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ are $C^{3,1}$-smooth on $\partial\widetilde{\Omega}\cap\{|y^{\prime}|\leq\sqrt{M}r_{0}\}$, and their $C^{3,1}$-norms are independent of $M$.
Define a set
$\omega:=\left\{y\in\widetilde{\Omega}:\ y_{n}<\beta^{2},\ |y_{i}|<\beta,\ \,\forall k+1\leq i\leq n-1\right\},$
where $\beta$ is a small positive constant to be determined. We analyze the geometry of $\omega$. Clearly, the set $\omega$ is convex. Looking at $\partial\omega\cap\{y_{n}=\beta^{2}\}$, by (4.20) we have
$\beta^{2}\geq\frac{1}{2}|\hat{y}|^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big).$
Since $M\gg 1$, any point with $|\hat{y}|=2\beta$ must be outside of $\omega$. So the set $\omega$ is contained in the ball $B_{\sqrt{n+3}\beta}$ by the convexity of $\omega$. It is trivial that
$y_{\star}:=\left(0,\dots,0,\frac{\beta}{4n},\dots,\frac{\beta}{4n},\frac{\beta^{2}}{2}\right)\in\omega,$
and
[…]
Then, since $\tilde{\rho}(y^{\prime})\leq|y^{\prime}|^{2}$, the ball $B_{r_{\star}}(y_{\star})$ is contained in $\omega$, where
$r_{\star}:=\min\left\{\frac{\beta^{2}}{2},\beta-\frac{\beta}{4n},\min\left\{\frac{\beta}{4n},\frac{\beta^{2}}{2}-\frac{\beta^{2}}{4n}\right\}\right\}.$
Using the convexity of $v$ and (4.21), we derive
$0\leq v\leq\sup_{\partial\omega\cap\partial\widetilde{\Omega}}\widetilde{\varphi}\leq C_{2}\quad\text{in}\ \overline{\omega},$ (4.22)
where $C_{2}$ depends only on $\|\varphi\|_{C^{3,1}(\partial\Omega)}$. Define
$\tilde{g}:=\tilde{f}^{\frac{3}{2n-2}}\quad\text{in}\ \widetilde{\Omega},\quad\text{and}\quad p=\frac{2n-2}{3}.$
By the same argument as in Step 3 of the proof of Lemma 4.2, we get
$\left\|\tilde{g}\right\|_{C^{2,1}(\overline{\omega})}\leq C_{3},$
where $C_{3}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$.
The main difficulty is that we cannot control the convexity of $\partial\widetilde{\Omega}$ in (4.20). An observation is that the directions, in which the convexity is under control, with respect to $\partial\widetilde{\Omega}$ and $\widetilde{\varphi}$ complement each other.
Step 1: we prove $|Dv|\leq C$ in $\omega$. Define
$\underline{v}(y):=\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}+\frac{1}{2}Ky_{n}^{2}-Ky_{n},$
where constants $\sigma>0$ and $K>0$ are to be determined. Let $K$ large depending on $\sigma$ such that
$\det D^{2}\underline{v}=\sigma^{n-1}K\geq 2^{n}\sup_{\omega}\tilde{f}$. (4.23)
We now verify that
$\underline{v}\leq v\quad\text{on}\ \partial\omega.$ (4.24)
The boundary $\partial\omega$ consists of three parts: $\partial_{1}\omega\cup\partial_{2}\omega\cup\partial_{3}\omega$, where $\partial_{1}\omega:=\partial\omega\cap\partial\widetilde{\Omega}$, $\partial_{2}\omega:=\partial\omega\cap\{y_{n}=\beta^{2}\}$, and $\partial_{3}\omega:=\partial\omega\cap\{|y_{i}|=\beta\ \text{for some}\ i=k+1,\dots,n-1\}$. It follows from (4.21) that on $\partial_{1}\omega$,
$v=\frac{1}{2}|\tilde{y}|^{2}+\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}y_{i}^{2}+O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}|\tilde{y}|^{2}-|\hat{y}|^{2}$
by taking $\beta$ small. By (4.20), we have
$y_{n}=\frac{1}{2}|\hat{y}|^{2}+\frac{1}{2}\sum_{i=k+1}^{n-1}d_{i}y_{i}^{2}+\frac{1}{\sqrt{M}}O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}|\hat{y}|^{2}-\beta|\tilde{y}|^{2}$
on $\partial_{1}\omega$, since $M\gg 1$. Hence
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}-\frac{1}{2}Ky_{n}\leq\left(\frac{1}{2}\sigma+\frac{1}{2}K\beta\right)\left|\tilde{y}\right|^{2}+\left(\frac{1}{2}\sigma-\frac{1}{8}K\right)\left|\hat{y}\right|^{2}\leq v$
on $\partial_{1}\omega$, by taking $\sigma<1/4$ and $\beta$ small depending on $K$. It holds that
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}-\frac{1}{2}K\beta^{2}\leq\frac{1}{2}\beta^{2}((n+3)\sigma-K)\leq 0\leq v\quad\text{on}\ \partial_{2}\omega.$
We only consider the piece $\partial^{\prime}_{3}\omega:=\partial\omega\cap\{y_{n-1}=\beta\}$, since other pieces of $\partial_{3}\omega$ can be handled in the same way. Let $\varepsilon_{0}>0$ be a small parameter to be determined later. It holds that
$\underline{v}\leq\frac{1}{2}\beta^{2}((n+3)\sigma-K\varepsilon_{0})\leq 0\leq v\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}\geq\varepsilon_{0}\beta^{2}\right\}$
by taking $\sigma$ small depending on $\varepsilon_{0}$.
We claim that
$v\geq\frac{1}{4}\beta^{2}\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\}.$ (4.25)
It follows from (4.21) that
$v\geq\frac{1}{2}\beta^{2}+O\big(\beta^{3}\big)\geq\frac{1}{4}\beta^{2}\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\}\cap\partial\widetilde{\Omega}$,
by taking $\beta$ small. It remains to consider points $p=(\hat{p},\tilde{p},p_{n})$ on $\partial^{\prime}_{3}\omega\cap\{y_{n}<\varepsilon_{0}\beta^{2}\}\cap\widetilde{\Omega}$, where $\hat{p}:=(p_{1},\dots,p_{k})$ and $\tilde{p}:=(p_{k+1},\dots,p_{n-1})$. Define
[…]
and
$p^{*}:=(0,\tilde{p},p_{n}+\delta),\quad p^{0}:=(0,\tilde{p},\tilde{\rho}(0,\tilde{p}))\in\partial\widetilde{\Omega}$.
From (4.20), we deduce
$\left|\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})\right|=O\big(\beta^{2}\big).$
Then $\delta=O(\beta^{3})$. Since $\widetilde{\Omega}$ is convex,
$p^{*}_{n}=p_{n}+\delta>p^{0}_{n}+\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})\cdot\hat{p}+\delta\geq p^{0}_{n},$
which implies $p^{*}\in\widetilde{\Omega}$. Extend the segment $\overline{pp^{*}}$ until it intersects $\partial\widetilde{\Omega}$ at a point denoted by $\bar{p}$ (see Figure 2).
Figure 2. The convexity of $\partial\widetilde{\Omega}$ in the direction $\tilde{y}$ is not under control. Cut $\widetilde{\Omega}$ by the hyperplane $\tilde{y}=\tilde{p}$ through the point $p$, and relate it to points on the boundary $\partial\widetilde{\Omega}$.
From (4.20), the graph of $\partial\widetilde{\Omega}$ near $p^{0}$ in the direction $\hat{y}$ is a perturbation of the paraboloid $|\hat{y}|^{2}/2$, by taking $\beta$ small. Hence we have
$\frac{1}{2}|p^{*}-p|\leq|\bar{p}-p^{*}|.$ (4.26)
Since the slope of $\overline{pp^{*}}$ is $|\nabla_{\hat{y}}\tilde{\rho}(0,\tilde{p})|=O(\beta^{2})$, we solve
$\frac{1}{4}t^{2}\leq\left(\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)\right)+O\big(\beta^{2}\big)t$
to deduce $t=O(\beta)$. Thus
$\bar{p}_{n}\leq\left(\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)\right)+O\big(\beta^{3}\big)\leq 2\varepsilon_{0}\beta^{2}$ (4.27)
by taking $\beta$ small depending on $\varepsilon_{0}$. Using (4.20) and (4.27) we derive
$\frac{1}{2}\sum_{i=1}^{k}\left|\bar{p}_{i}\right|^{2}\leq\bar{p}_{n}+O\big(\beta^{3}\big)\leq 3\varepsilon_{0}\beta^{2}.$ (4.28)
For any $T^{-1}x=y=(0,\tilde{y},\tilde{\rho}(0,\tilde{y}))\in\partial\widetilde{\Omega}\cap B_{\theta_{k+1}}(0)$ with $\theta_{k+1}$ given in Assumption 4.4, we have
$\hat{x}=0\quad\text{and}\quad|x_{i}|=\left|\frac{y_{i}}{M_{i}}\right|\leq\theta_{k+1}\frac{b_{k}}{\sqrt{b_{i}}},\quad\forall k+1\leq i\leq n-1.$
Since $|\nabla u(x)-\nabla u(0)|=O(|x^{\prime}|)$ and $|u_{\nu}(x)-u_{n}(x)|=O(|x^{\prime}|^{2})$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ (see later (6.7), (6.8)), using Assumption 4.4, we infer
$\left|v_{y_{n}}(y)\right|$ $=\frac{1}{b_{k}}\left(\left|u_{\nu}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)\right|+O\Big(\big|x^{\prime}\big|^{2}\Big)\right)$
$=\frac{1}{b_{k}}\left|u_{n}(0)+\nabla_{x^{\prime}}(u_{\nu})\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\cdot x^{\prime}\right|+C\theta_{k+1}^{2}$
$=\frac{1}{b_{k}}\sum_{i=k+1}^{n-1}(u_{\nu\eta_{i}}+\partial_{x_{i}}\nu\cdot\nabla u)\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)x_{i}+C\theta_{k+1}^{2}$
$\leq\frac{1}{b_{k}}\sum_{i=k+1}^{n-1}\left(C\sqrt{b_{i}}x_{i}+C\big|x^{\prime}\big|^{2}\right)+C\theta_{k+1}^{2}$
$\leq C_{4}$
for some $t\in(0,1)$, where $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. By the convexity of $v$, (4.21), and (4.27), we get
$v(p^{*})$ $\geq v\big(p^{0}\big)+\nabla v\big(p^{0}\big)\cdot\big(p^{*}-p^{0}\big)$
$\geq\widetilde{\varphi}\big(p^{0}\big)-C_{4}\bar{p}_{n}$
$\geq\frac{1}{2}\beta^{2}-2C_{4}\varepsilon_{0}\beta^{2}+O\big(\beta^{3}\big)$
$\geq\frac{1}{3}\beta^{2}$ (4.29)
by taking $\varepsilon_{0}$ small. Furthermore, we obtain from (4.21) and (4.27)–(4.2) that
$v(\bar{p})=\widetilde{\varphi}(\bar{p})$ $=\frac{1}{2}\sum_{i=1}^{k}\tilde{b}_{i}\left|\bar{p}_{i}\right|^{2}+\widetilde{\varphi}\big(p^{0}\big)+O\big(\beta^{3}\big)$
$\leq v(p^{*})+C_{4}\bar{p}_{n}+4\varepsilon_{0}\beta^{2}$
$\leq v(p^{*})+(2C_{4}+4)\varepsilon_{0}\beta^{2}$
$\leq\frac{9}{8}v(p^{*})$ (4.30)
by taking $\varepsilon_{0}$ small. Using the convexity of $v$, (4.30), (4.26), and (4.2), we get
$v(p)\geq-\frac{|p^{*}-p|}{|\bar{p}-p^{*}|}v(\bar{p})+\frac{|\bar{p}-p|}{|\bar{p}-p^{*}|}v(p^{*})\geq\left(1-\frac{1}{8}\frac{|p^{*}-p|}{|\bar{p}-p^{*}|}\right)v(p^{*})\geq\frac{1}{4}\beta^{2}.$
In conclusion, claim (4.25) is valid.
We deduce from (4.25) that
$\underline{v}\leq\frac{1}{2}\sigma\big|y^{\prime}\big|^{2}\leq\frac{n+3}{2}\sigma\beta^{2}\leq v\quad\text{on}\ \partial^{\prime}_{3}\omega\cap\left\{y_{n}<\varepsilon_{0}\beta^{2}\right\},$
by taking $\sigma$ small. Thus (4.24) holds. It follows from the comparison principle that
$\underline{v}\leq v\quad\text{in}\ \overline{\omega}.$
Since $\underline{v}(0)=v(0)=0$,
$v_{n}(0)\geq\underline{v}_{n}(0).$
Using the convexity of $v$ and (4.22), we get
$v_{n}(0)\leq\frac{v((0,\dots,0,\beta^{2}))-v(0)}{\beta^{2}}\leq\frac{C_{2}}{\beta^{2}}.$
So $|v_{n}(0)|\leq C$. For any point $y_{0}\in\partial\omega\cap\partial\widetilde{\Omega}$, letting $y_{0}$ be the origin and $\tilde{\nu}(y_{0})=e_{n}$ by translation and rotation, letting $v(y_{0})=0$ and $\nabla v(y_{0})=0$ by subtracting an affine function, with similar arguments as above, we get
$|v_{\tilde{\nu}}(y_{0})|\leq C,\quad\forall y_{0}\in\partial\omega\cap\partial\widetilde{\Omega}$, (4.31)
where $\tilde{\nu}=\tilde{\nu}(y)$ denotes the unit interior normal vector to $\partial\widetilde{\Omega}$ at $y$. From (4.31) and the convexity of $v$, it follows that
$|Dv|\leq C\quad\text{in}\ \overline{\omega}$
for a smaller $\beta$, where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$.
Step 2. Define tangential vector fields
$\tilde{\eta}^{l}(y):=e_{l}+\tilde{\rho}_{l}\big(y^{\prime}\big)e_{n}\quad\text{for}\ l=1,2,\dots,n-1,$
in $\overline{\omega}$. Indeed, for any $l\in\{1,2,\dots,n-1\}$,
$T\eta^{l}=T(e_{l}+\partial_{x_{l}}\rho e_{n})=M_{l}e_{l}+\partial_{x_{l}}\rho M_{n}e_{n}=M_{l}\left(e_{l}+\partial_{y_{l}}\tilde{\rho}e_{n}\right)=M_{l}\tilde{\eta}^{l}.$ (4.32)
Construct functions
$\Psi^{l}(y):=\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A|y|^{2}-B_{A}\bigg(v-\frac{1}{2}\underline{v}+(C_{5}+1)y_{n}\bigg)\quad\text{in}\ \overline{\omega},$
for $l=1,2,\dots,n-1$, where positive constants $A$ and $B_{A}$ are to be determined, and $C_{5}:=\sup_{\omega}|D(v-\underline{v}/2)|$.
First, we choose $A$ large so that the functions $\Psi^{l}$ cannot attain its maximum in $\omega$ for $l=1,2,\dots,n-1$. Fix $l\in\{1,2,\dots,n-1\}$, and suppose, for contradiction, that $\Psi^{l}$ attains its maximum at an interior point $y_{*}\in\omega$. Then at the point $y_{*}$, we have
$\partial_{n}\Psi^{l}=0\quad\text{and}\quad F^{ij}\big[D^{2}v\big]\Psi^{l}_{ij}\leq 0.$ (4.33)
From (4.33), at the point $y_{*}$,
$\pm\partial_{n}\Big(\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})\Big)=B_{A}\bigg(\bigg(v-\frac{1}{2}\underline{v}\bigg)_{n}+C_{5}+1\bigg)-2Ay_{n}\geq B_{A}-2A,$
which yields $|D^{2}v(y_{*})|>1$ by taking $B_{A}$ large depending on $A$. Since $\lambda(D^{2}v(y_{*}))\in\Gamma_{2}$,
$\sigma_{1}\big[D^{2}v(y_{*})\big]\geq\left|D^{2}v(y_{*})\right|>1.$ (4.34)
For convenience, introduce a coordinate system $z$, obtained by an orthogonal transformation of the $y$-coordinates, such that at the point $y_{*}$ in the $z$-coordinates,
$D_{z}^{2}v\ \ \text{is diagonal, which implies}\ \ F^{ij}_{z}\big[D^{2}_{z}v\big]\ \ \text{is also diagonal}.$
Computing (4.33) at the point $y_{*}$ in the $z$-coordinates, by the concavity and homogeneity of $F$ and (4.23), we obtain
$0$ $\geq F^{ij}_{z}\big[D^{2}_{z}v\big]\Psi^{l}_{z_{i}z_{j}}$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}(v-\widetilde{\varphi})\right)_{z_{i}z_{j}}+2A\operatorname{tr}F^{ij}\big[D^{2}v\big]$
$\hskip 12.80373pt+B_{A}\bigg(\frac{1}{2}F\big[D^{2}\underline{v}\big]-F\big[D^{2}v\big]\bigg)$
$\geq\pm F^{ij}_{z}\big[D^{2}_{z}v\big]\left(\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}z_{j}}+2\partial_{z_{j}}\tilde{\eta}^{l}_{z}\cdot\nabla_{z}v_{z_{i}}\right)+(2A-C)\operatorname{tr}F^{ij}$
$=\pm\tilde{\eta}^{l}\cdot\nabla\left(\tilde{g}^{\frac{p}{n}}\right)\pm 2\sum_{i=1}^{n}F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\times\partial_{z_{i}}\tilde{\eta}^{l,i}_{z}+(2A-C)\operatorname{tr}F^{ij}$
$\geq-\left|\frac{p}{n}\tilde{g}^{\frac{p}{n}-\frac{1}{2}}\frac{\tilde{\eta}^{l}\cdot\nabla\tilde{g}}{\sqrt{\tilde{g}}}\right|-C\sum_{i=1}^{n}\left|F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\right|+(2A-C)\operatorname{tr}F^{ij},$ (4.35)
[…]
$\tilde{f}^{\frac{1}{n}}=\sum_{j=1}^{n}F^{jj}_{z}\big[D^{2}_{z}v\big]v_{z_{j}z_{j}}\geq\left|F^{ii}_{z}\big[D^{2}_{z}v\big]v_{z_{i}z_{i}}\right|\quad\text{for}\ i=1,2,\dots,n,$ (4.36)
at the point $y_{*}$ in the $z$-coordinates. Combining (4.35), (4.36), Lemma 2.1, (2.13), and (4.34), we derive at the point $y_{*}$,
$0$ $\geq-C_{6}\tilde{f}^{\frac{1}{4(n-1)}-\frac{1}{n(n-1)}}-C\tilde{f}^{\frac{1}{n}}+\frac{2A-C}{C_{7}}\sigma_{1}^{\frac{1}{n-1}}\big[D^{2}v\big]\tilde{f}^{\frac{-1}{n(n-1)}}$
$\geq\frac{2A-C}{C_{7}}\tilde{f}^{\frac{-1}{n(n-1)}}-C_{6}\tilde{f}^{\frac{1}{4(n-1)}-\frac{1}{n(n-1)}}-C\tilde{f}^{\frac{1}{n}}$
$>0$
by taking $A$ large, which is a contradiction, where $C_{7}>0$ is universal, and $C_{6}$ depends on $\|\tilde{g}\|_{C^{1,1}(\overline{\omega})}$ . Thus $\Psi^{l}$ cannot attain its maximum in $\omega$ for $l=1,2,\dots,n-1$.
Second, we choose $B_{A}$ large, depending on $A$, so that
$\Psi^{l}\leq 0\quad\text{on}\ \partial\omega,\quad\forall 1\leq l\leq n-1.$
Fix $l\in\{1,2,\dots,n-1\}$, it follows from (4.24) that
$\Psi^{l}\leq\pm\tilde{\eta}^{l}\cdot\nabla(v-\widetilde{\varphi})+A\big|y^{\prime}\big|^{2}-\frac{1}{2}B_{A}(v+y_{n}).$ (4.37)
Combining (4.20) and (4.21) yields
$v+y_{n}\geq\frac{1}{2}\big|y^{\prime}\big|^{2}+O\Big(\big|y^{\prime}\big|^{3}\Big)\geq\frac{1}{4}\big|y^{\prime}\big|^{2}\quad\text{on}\ \partial_{1}\omega,$
which implies
$\Psi^{l}\leq A\big|y^{\prime}\big|^{2}-\frac{1}{8}B_{A}\big|y^{\prime}\big|^{2}\leq 0\quad\text{on}\ \partial_{1}\omega,$
by taking $B_{A}$ large. It follows from (4.25) that
$v+y_{n}\geq\min\left\{\frac{1}{4}\beta^{2},\varepsilon_{0}\beta^{2}\right\}=\varepsilon_{0}\beta^{2}\quad\text{on}\ \partial_{2}\omega\cup\partial_{3}\omega.$ (4.38)
Then using (4.37) and (4.38), we have
$\Psi^{l}\leq C+A-\frac{1}{2}B_{A}\varepsilon_{0}\beta^{2}\leq 0\quad\text{on}\ \partial_{2}\omega\cup\partial_{3}\omega,$
[…]
By the comparison principle,
$\Psi^{l}\leq 0\quad\text{in}\ \overline{\omega},\quad\forall 1\leq l\leq n-1.$
Since $\Psi^{l}(0)=0$, we get $\Psi^{l}_{n}(0)\leq 0$, which yields
$|v_{nl}(0)|\leq C,\quad\forall 1\leq l\leq n-1,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. For any point
$T^{-1}y_{0}=x_{0}\in\partial\Omega\cap\left\{\big|x^{\prime}\big|\leq r_{0}\right\}\cap\left\{|x|\leq\theta_{k}\sqrt{b_{k}}\right\},$
where $\theta_{k}>0$ is chosen small so that $|x_{0}|\leq\theta_{i}\sqrt{b_{i}}/2$ for all $k+1\leq i\leq n-1$, let $v(y_{0})=0$, $\nabla v(y_{0})=0$ by subtracting an affine function, and let $y_{0}$ be the origin, $\tilde{\nu}(y_{0})=e_{n}$, $\tilde{\eta}^{l}(y_{0})=e_{l}$ for all $1\leq l\leq n-1$ by translation and rotation. With similar arguments as above, we derive
$\left|v_{\tilde{\nu}\tilde{\eta}^{l}}(y_{0})\right|\leq C,\quad\forall 1\leq l\leq n-1,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. Pulling back to the $x$-coordinates, we obtain from (4.32) that
$C\geq M^{2}\left|\tilde{\nu}^{\prime}D^{2}_{y}u(y_{0})\tilde{\eta}^{l}\right|$ $=M^{2}\left|\big(T^{-1}\tilde{\nu}\big)^{\prime}D^{2}_{x}u(x_{0})T^{-1}\tilde{\eta}^{l}\right|$
$=M^{2}\left|u_{\nu\eta^{l}}(x_{0})\right|\frac{1}{|T\nu|}\frac{1}{M_{l}}$
for $l=1,2,\dots,n-1$, where $\tilde{\nu}=T\nu/|T\nu|$. Thus by (4.19), we have
$\left|u_{\nu\eta^{l}}(x_{0})\right|\leq C\frac{|T\nu|M_{l}}{M^{2}}\leq C\frac{1}{\sqrt{M}}=C\sqrt{b_{k}},\quad\forall 1\leq l\leq k,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $|f^{3/(2n-2)}|_{C^{2,1}(\overline{\Omega})}$. This completes the proof of Lemma 4.5. ∎
Theorem 4.6.
Let $u\in C^{3,1}(\overline{\Omega})$ be a convex solution of equation (1.1) with $k=n$, $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega}f>0$, and $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega})$. Then
$\sup_{\partial\Omega}|u_{\nu\nu}|\leq C,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$, but is independent of $\inf_{\Omega}f$.
Proof.
At the origin, we have
$u_{nn}(0)\prod_{i=1}^{n-1}b_{i}-\sum_{j=1}^{n-1}\bigg(\prod_{i\neq j}b_{i}\bigg)u_{nj}^{2}(0)=f(0).$
Lemma 4.3 ensures that Assumption 4.4 holds for k=n2k=n-2. Thus using Lemma 4.5, by induction we obtain
$u_{ni}(0)\leq C\sqrt{b_{i}}\quad\text{for}\ i=1,2,\dots,n-2.$ (4.39)
Then it follows from Lemma 4.2, Lemma 4.3, and (4.39) that
$0\leq u_{nn}(0)\leq\frac{1}{\prod_{i=1}^{n-1}b_{i}}f(0)+\sum_{j=1}^{n-1}\frac{1}{b_{j}}u_{nj}^{2}(0)\leq C,$
where $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega})}$. This completes the proof of Theorem 4.6. ∎
Combining Lemmas 3.1, 3.2, 3.3, and Theorem 4.6, we obtain the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$, and then Theorem 1.1 follows from an approximation argument.
5. The weakly interior estimate
We recall the following comparison principle established by Ivochkina-Trudinger-Wang [16], and provide a proof for completeness.
Lemma 5.1 (Lemma 3.1 of [16]).
Let $w(x,y)$, $v(x,y)$ be functions in $\overline{\Omega}\times\mathbb{R}^{N}$ that are both $p$-homogeneous in $y\in\mathbb{R}^{N}$ for some $p\in\mathbb{R}$. Let $w$, $v\in C^{2}(\Omega\times\mathbb{S}^{N-1})\cap C^{0}(\overline{\Omega}\times\mathbb{S}^{N-1})$ with $v>0$ in $\overline{\Omega}\times\mathbb{S}^{N-1}$. Define the linear operator $\mathscr{L}$ by
$\mathscr{L}(w):=\operatorname{tr}\big(\widetilde{G}D^{2}w\big)+\langle b,Dw\rangle,$
where $\widetilde{G}=\widetilde{G}(x,y)$ is an $(n+N)\times(n+N)$ positive semi-definite matrix and $b=b(x,y)$ is an $(n+N)$-dimensional vector. If there exist two positive constants $\mu_{1}$, $\mu_{2}$ such that
$\begin{cases}\mathscr{L}(w)\geq-\mu_{1}\quad\text{in}\ \Omega\times\mathbb{S}^{N-1},\\ \mathscr{L}(v)\leq-\mu_{2}\quad\text{in}\ \Omega\times\mathbb{S}^{N-1},\end{cases}$ (5.1)
then
$\sup_{\overline{\Omega}\times\mathbb{S}^{N-1}}\frac{w}{v}\leq\frac{\mu_{1}}{\mu_{2}}+\sup_{\partial\Omega\times\mathbb{S}^{N-1}}\frac{w}{v}.$
Proof.
Since $w$, $v$ are $p$-homogeneous in $y\in\mathbb{R}^{N}$, we have
$\frac{w}{v}(x,y)\in C^{2}(Q)\cap C^{0}(\overline{Q})\quad\text{and}\quad\frac{w}{v}(x_{0},y_{0})=\sup_{\overline{Q}}\frac{w}{v},$
where $Q:=\Omega\times(B_{2}\setminus B_{1/2})$ and $(x_{0},y_{0})\in\overline{\Omega}\times\mathbb{S}^{N-1}$. We only need to consider the case of $(x_{0},y_{0})\in\Omega\times\mathbb{S}^{N-1}$ with $w(x_{0},y_{0})>0$. At the maximum point $(x_{0},y_{0})$,
$\partial_{i}\Big(\frac{w}{v}\Big)=\frac{\partial_{i}w}{v}-\frac{w\partial_{i}v}{v^{2}}=0,\quad\forall 1\leq i\leq n+N,$ (5.2)
and the Hessian
$\partial_{ij}\Big(\frac{w}{v}\Big)=\frac{\partial_{ij}w}{v}-\frac{\partial_{i}w\partial_{j}v+\partial_{j}w\partial_{i}v}{v^{2}}+\frac{2w\partial_{i}v\partial_{j}v}{v^{3}}-\frac{w\partial_{ij}v}{v^{2}}$
is an $(n+N)\times(n+N)$ negative semi-definite matrix. Then at the point $(x_{0},y_{0})$,
$0\geq\widetilde{G}^{ij}\partial_{ij}\Big(\frac{w}{v}\Big)$ $=\widetilde{G}^{ij}\left(\frac{\partial_{ij}w}{v}-\frac{w\partial_{ij}v}{v^{2}}+\left(\frac{2w\partial_{i}v}{v^{3}}-\frac{2\partial_{i}w}{v^{2}}\right)\partial_{j}v\right)$
[…]
$=\frac{\mathscr{L}(w)-b^{i}\partial_{i}w}{v}-\frac{w}{v^{2}}\left(\mathscr{L}(v)-b^{i}\partial_{i}v\right)$
[…]
$\overset{w,v>0}{\geq}\frac{1}{v}\left(-\mu_{1}+\frac{w}{v}\mu_{2}\right).$
This yields
$\frac{w}{v}(x_{0},y_{0})\leq\frac{\mu_{1}}{\mu_{2}}.$
The proof of Lemma 5.1 is complete. ∎
We now proceed to prove the weakly interior estimate.
Proof of Theorem 1.2.
We will apply Lemma 5.1 with $N=n+1$ and coordinates $(x,\xi,\eta)\in\overline{\Omega}\times\mathbb{R}^{n}\times\mathbb{R}$. Define functions
$w:=u_{\xi\xi}+2\eta u_{\xi}+\eta^{2}u$
and
$v:=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\varrho(\psi+\beta)\eta^{2},$
where $0<\alpha<1$, $0<\varrho<1$, and $0<\beta\ll 1$ are parameters to be determined; and the function $\psi$ is given in Lemma 2.7. Define a linear operator
$\mathscr{L}(w):=\operatorname{tr}\big(\widetilde{G}D^{2}w\big)+\langle b,Dw\rangle,$
where
$\widetilde{G}:=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix},\quad b:=\begin{pmatrix}0_{n\times 1}\\ -2Gq\\ 0_{1\times 1}\end{pmatrix}$,
with $G$ being the $n\times n$ matrix-valued function defined in (2.12). The scalar function $r$ and the $n$-dimensional vector-valued function $q$ are defined by
$r:=\frac{\alpha\psi_{\xi}}{\psi+\beta},\quad q:=\frac{1}{\psi+\beta}\left(\frac{\xi}{4}+\frac{\alpha\psi_{\xi}}{\psi+\beta}\nabla\psi\right).$ (5.3)
It holds that the quadratic form
$\left(\xi_{1},\xi_{2},\eta\right)^{T}\widetilde{G}\left(\xi_{1},\xi_{2},\eta\right)=\left(\xi_{1}+r\xi_{2}+\eta q\right)^{T}G\left(\xi_{1}+r\xi_{2}+\eta q\right)\geq 0$
is valid for all $(\xi_{1},\xi_{2},\eta)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$, which implies that $\widetilde{G}$ is positive semi-definite. Due to the substantial computations involved, we divide the proof into five steps for clarity.
Step 1. For convenience, denote
$\hat{v}:=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}.$
We claim that
$\mathscr{L}(\hat{v})\leq-\frac{1}{C_{\psi}}\frac{1}{8\alpha}|\xi|^{2}$ (5.4)
in $\Omega\times\mathbb{R}^{n}\times\mathbb{R}$ for all $\alpha\in(0,\alpha_{0})$, where constants $C_{\psi}$ and $\alpha_{0}$ depend on $\|\psi\|_{C^{3}(\overline{\Omega})}$ to be determined. It holds that
$\widetilde{G}D^{2}\hat{v}=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix}\begin{pmatrix}\partial_{x_{i}x_{j}}&\partial_{x_{i}\xi_{j}}&0\\ \partial_{\xi_{i}x_{j}}&\partial_{\xi_{i}\xi_{j}}&0\\ 0&0&0\end{pmatrix}[\hat{v}$],
where $1\leq i,\ j\leq n$. Since $G$ is symmetric, we derive
$\mathscr{L}(\hat{v})=G^{ij}\partial_{x_{i}x_{j}}\hat{v}+2rG^{ij}\partial_{x_{i}\xi_{j}}\hat{v}+r^{2}G^{ij}\partial_{\xi_{i}\xi_{j}}\hat{v}-2G^{ij}q_{j}\partial_{\xi_{i}}\hat{v}$. (5.5)
Compute derivatives
$\partial_{x_{i}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\left(\frac{1-\alpha}{4\alpha}|\xi|^{2}\psi_{x_{i}}+2\psi_{\xi}\psi_{\xi x_{i}}-\alpha\frac{\psi_{\xi}^{2}\psi_{x_{i}}}{\psi+\beta}\right);$
$\partial_{x_{i}x_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\left(\frac{1-\alpha}{4\alpha}|\xi|^{2}\psi_{x_{i}x_{j}}+2\psi_{\xi x_{i}}\psi_{\xi x_{j}}+2\psi_{\xi}\psi_{\xi x_{i}x_{j}}\right)$ (5.6)
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}\right)$
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(-2\alpha\psi_{\xi}\psi_{x_{i}}\psi_{\xi x_{j}}-\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$r\partial_{x_{i}\xi_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{1-\alpha}{2}\psi_{\xi}\psi_{x_{i}}\xi_{j}+2\alpha\psi_{\xi}\psi_{\xi x_{i}}\psi_{x_{j}}+2\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}\right)$ (5.7)
$\hskip 12.80373pt-\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$\partial_{\xi_{i}}\hat{v}$ $=\frac{(\psi+\beta)^{1-\alpha}}{2\alpha}\xi_{i}+(\psi+\beta)^{-\alpha}2\psi_{\xi}\psi_{x_{i}};$
$q_{j}\partial_{\xi_{i}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha}}\frac{\xi_{i}\xi_{j}}{8\alpha}+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{1}{2}\psi_{\xi}\xi_{i}\psi_{x_{j}}+\frac{1}{2}\psi_{\xi}\psi_{x_{i}}\xi_{j}\right)$ (5.8)
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}};$
$r^{2}\partial_{\xi_{i}\xi_{j}}\hat{v}$ $=\frac{1}{(\psi+\beta)^{\alpha+1}}\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}+\frac{1}{(\psi+\beta)^{\alpha+2}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}.$ (5.9)
Combining (5.5)–(5.9) with the symmetry of $G$ yields that
[…]
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}\psi_{x_{i}}\xi_{j}\right)$
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha+1}}\left(3\alpha\psi_{\xi}^{2}\psi_{x_{i}x_{j}}+\frac{1}{2}\alpha\psi_{\xi}^{2}\delta_{ij}\right)$
$\hskip 12.80373pt-\frac{G^{ij}}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+3\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}.$
Since $\operatorname{tr}\left(GD^{2}\psi\right)\leq-1$ from (2.17), $\operatorname{tr}G=1$, and $G$ is positive definite, we get that
[…]
$\hskip 12.80373pt+\frac{1}{(\psi+\beta)^{\alpha+1}}\left(\frac{\alpha-1}{4}|\xi|^{2}G^{ij}\psi_{x_{i}}\psi_{x_{j}}-(1+\alpha)\psi_{\xi}G^{ij}\psi_{x_{i}}\xi_{j}-\frac{5}{2}\alpha\psi_{\xi}^{2}\right)$
$\hskip 12.80373pt-\frac{G^{ij}}{(\psi+\beta)^{\alpha+2}}\big(\alpha^{2}+3\alpha\big)\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}$
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{1}\|\psi\|_{C^{3}(\overline{\Omega})}^{2}\right)|\xi|^{2}$ (5.10)
$\hskip 12.80373pt+\frac{G^{ij}}{(\psi+\beta)^{\alpha}}\left[-\frac{1}{4\alpha}\xi_{i}\xi_{j}-\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}-\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right],$
where the constant $C_{1}>0$ is universal. It holds that
$0$ $\leq\sum_{s=1}^{n}\left[\sum_{i=1}^{n}\Big(G^{1/2}\Big)^{si}\left(\sqrt{\alpha^{2}+3\alpha}\frac{\psi_{\xi}}{\psi+\beta}\psi_{x_{i}}+\frac{1+\alpha}{2\sqrt{\alpha^{2}+3\alpha}}\xi_{i}\right)\right]^{2}$
$=\sum_{i,j=1}^{n}G^{ij}\left[\frac{\alpha^{2}+3\alpha}{(\psi+\beta)^{2}}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}+\frac{1+\alpha}{\psi+\beta}\psi_{\xi}\psi_{x_{i}}\xi_{j}+\frac{(1+\alpha)^{2}}{4\alpha(3+\alpha)}\xi_{i}\xi_{j}\right].$ (5.11)
By (5) and (5.11), we obtain that
[…]
$\leq\frac{1}{(\psi+\beta)^{\alpha}}\left(-\frac{1-\alpha}{4\alpha}+C_{\psi}\right)|\xi|^{2}$
$\leq-\frac{1}{\psi+1}\frac{1}{8\alpha}|\xi|^{2}$
for all $\alpha\in(0,\alpha_{0})$ with $\alpha_{0}>0$ chosen sufficiently small, where constants $C_{\psi}$ and $\alpha_{0}$ depend on $\|\psi\|_{C^{3}(\overline{\Omega})}$. This establishes claim (5.4).
Step 2. It holds that
$\widetilde{G}D^{2}\left[(\psi+\beta)\eta^{2}\right]=\begin{pmatrix}G&rG&Gq\\ rG&r^{2}G&rGq\\ q^{\prime}G&rq^{\prime}G&q^{\prime}Gq\end{pmatrix}\begin{pmatrix}\partial_{x_{i}x_{j}}&0&\partial_{x_{i}\eta}\\ 0&0&0\\ \partial_{\eta x_{j}}&0&\partial_{\eta\eta}\end{pmatrix}\left[(\psi+\beta)\eta^{2}\right],$
where $1\leq i,\ j\leq n$. By the symmetry of $G$, we compute
$\hskip 12.80373pt\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$
$=G^{ij}\left(\partial_{x_{i}x_{j}}+2q_{j}\partial_{\eta x_{i}}+q_{i}q_{j}\partial_{\eta\eta}\right)\left[(\psi+\beta)\eta^{2}\right]$
$=\eta^{2}G^{ij}\psi_{x_{i}x_{j}}+G^{ij}\left(\frac{1}{\psi+\beta}\eta\psi_{x_{i}}\xi_{j}+\frac{1}{(\psi+\beta)^{2}}4\alpha\eta\psi_{\xi}\psi_{x_{i}}\psi_{x_{j}}\right)$
$\hskip 12.80373pt+G^{ij}\left(\frac{1}{\psi+\beta}\frac{1}{8}\xi_{i}\xi_{j}+\frac{1}{(\psi+\beta)^{2}}\alpha\psi_{\xi}\psi_{x_{i}}\xi_{j}+\frac{1}{(\psi+\beta)^{3}}2\alpha^{2}\psi_{\xi}^{2}\psi_{x_{i}}\psi_{x_{j}}\right).$
Since $\operatorname{tr}\left(GD^{2}\psi\right)\leq-1$ from (2.17), $\operatorname{tr}G=1$, and $G$ is positive definite, we get that
$\mathscr{L}\big((\psi+\beta)\eta^{2}\big)$ $\leq-\eta^{2}+\frac{1}{\beta^{2}}C_{\psi}|\eta||\xi|+\frac{1}{\beta^{3}}C_{\psi}|\xi|^{2}$
$\leq-\frac{1}{2}\eta^{2}+\frac{1}{2\beta^{4}}C_{\psi}^{2}|\xi|^{2}+\frac{1}{\beta^{3}}C_{\psi}|\xi|^{2}$
$\leq-\frac{1}{2}\eta^{2}+\frac{1}{\beta^{4}}C_{\psi}^{2}|\xi|^{2},$ (5.12)
where the constant $C_{\psi}>0$ depends on $\|\psi\|_{C^{3}(\overline{\Omega})}$. Then it follows from (5.4) and (5.12) that
$\mathscr{L}(v)$ $\leq-\frac{1}{C_{\psi}}\frac{1}{8\alpha}|\xi|^{2}+\varrho\left(-\frac{1}{2}\eta^{2}+\frac{1}{\beta^{4}}C_{\psi}^{2}|\xi|^{2}\right)$
$=\left(-\frac{1}{8C_{\psi}}\frac{1}{\alpha}+C_{\psi}^{2}\right)|\xi|^{2}-\frac{1}{2}\beta^{4}\eta^{2}$
$\leq-\frac{1}{2}\beta^{4}\left(|\xi|^{2}+\eta^{2}\right)$ (5.13)
in $\Omega\times\mathbb{R}^{n}\times\mathbb{R}$ for all $\alpha\in(0,\alpha_{0})$, where we set
$\varrho:=\beta^{4}$
and $\alpha_{0}>0$ chosen sufficiently small depending on $\|\psi\|_{C^{3}(\overline{\Omega})}$. Fix $\alpha\in(0,\alpha_{0})$ depending on $\|\psi\|_{C^{3}(\overline{\Omega})}$ in subsequent arguments.
Step 3. It holds that
[…]
where $1\leq i,\ j\leq n$. By the symmetry of $G$, we compute
$\hskip 12.80373pt\mathscr{L}(w)$
$=G^{ij}\left(\partial_{x_{i}x_{j}}+2r\partial_{x_{i}\xi_{j}}+2q_{j}\partial_{x_{i}\eta}+r^{2}\partial_{\xi_{i}\xi_{j}}+2rq_{j}\partial_{\xi_{i}\eta}+q_{i}q_{j}\partial_{\eta\eta}-2q_{j}\partial_{\xi_{i}}\right)[w]$
$=G^{ij}2q_{j}(\partial_{x_{i}\eta}w-\partial_{\xi_{i}}w)+G^{ij}\big(u_{\xi\xi x_{i}x_{j}}+2\eta u_{\xi x_{i}x_{j}}+\eta^{2}u_{x_{i}x_{j}}\big)$
$\hskip 12.80373pt+G^{ij}2r\big(2u_{\xi x_{i}x_{j}}+2\eta u_{x_{i}x_{j}}\big)+G^{ij}r^{2}2u_{x_{i}x_{j}}+G^{ij}2rq_{j}2u_{x_{i}}+G^{ij}q_{i}q_{j}2u$
$=G^{ij}\big(u_{\xi\xi x_{i}x_{j}}+2\eta u_{\xi x_{i}x_{j}}+\eta^{2}u_{x_{i}x_{j}}\big)+4rG^{ij}\big(u_{\xi x_{i}x_{j}}+\eta u_{x_{i}x_{j}}\big)$
$\hskip 12.80373pt+2r^{2}G^{ij}u_{x_{i}x_{j}}+4rG^{ij}u_{x_{i}}q_{j}+2uG^{ij}q_{i}q_{j}.$
It follows from (2.13)–(2.16), (5.3), and $0\leq G\leq\mathbf{I}_{n}$ that
$\mathscr{L}(w)$ $\geq-K_{1}|\xi|^{2}\sigma_{1}^{\frac{-1}{k-1}}-2|\eta|K_{1}|\xi|\sigma_{1}^{\frac{-1}{k-1}}-4|r|K_{1}|\xi|\sigma_{1}^{\frac{-1}{k-1}}-4|r||\eta|K_{1}$
$\hskip 12.80373pt-4|r|\|u\|_{C^{1}(\overline{\Omega})}|q|-2\|u\|_{C^{0}(\overline{\Omega})}|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+r^{2}\right)-K_{1}\left(4\eta^{2}+r^{2}\right)-4r^{2}$
$\hskip 12.80373pt-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)|q|^{2}$
$\geq-K_{1}\sigma_{1}^{\frac{-1}{k-1}}\left(6|\xi|^{2}+\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)-K_{1}\left(4\eta^{2}+\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}\right)$
$\hskip 12.80373pt-4\frac{C_{\psi}}{\beta^{2}}|\xi|^{2}-\left(\|u\|_{C^{1}(\overline{\Omega})}^{2}+2\|u\|_{C^{0}(\overline{\Omega})}\right)\frac{C_{\psi}}{\beta^{4}}|\xi|^{2}$
$\geq-K\left(1+\sigma_{1}^{\frac{-1}{k-1}}\right)\frac{1}{\beta^{4}}\left(|\xi|^{2}+\eta^{2}\right),$ (5.14)
where $K>0$ depends on $\|\psi\|_{C^{1}(\overline{\Omega})}$, $\|u\|_{C^{1}(\overline{\Omega})}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Step 4. It holds that
$w$ $=\left(\theta\nu+\tau\sqrt{|\xi|^{2}-\theta^{2}}\right)^{T}D^{2}u\left(\theta\nu+\tau\sqrt{|\xi|^{2}-\theta^{2}}\right)+2\eta u_{\xi}+\eta^{2}u$
$\leq\theta^{2}\sup_{\partial\Omega}u_{\nu\nu}+C$ (5.15)
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$, where $\tau$ is a unit tangential direction, $\theta:=\langle\xi,\nu\rangle$, and the constant $C>0$ depends on $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Moreover, by (2.17) we get that
$v$ $=\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\beta^{4}(\psi+\beta)\eta^{2}$
$=\frac{\beta^{1-\alpha}}{4\alpha}|\xi|^{2}+\frac{1}{\beta^{\alpha}}(\left|\nabla\psi\right|\langle\nu,\xi\rangle)^{2}+\beta^{5}\eta^{2}$
$\geq\frac{1}{\beta^{\alpha}}\theta^{2}+\beta^{5}\big(|\xi|^{2}+\eta^{2}\big)$
$=\frac{1}{\beta^{\alpha}}\theta^{2}+\beta^{5}$ (5.16)
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$. Therefore, it follows from (5.15) and (5.16) that
$\frac{w}{v}\leq\frac{\theta^{2}\sup_{\partial\Omega}u_{\nu\nu}+C}{\theta^{2}\beta^{-\alpha}+\beta^{5}}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{C}{\beta^{5}}$
for all $(x,\xi,\eta)\in\partial\Omega\times\mathbb{S}^{n}$. Thus
$\sup_{\partial\Omega\times\mathbb{S}^{n}}\frac{w}{v}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{C}{\beta^{5}}.$ (5.17)
Step 5. Let
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})=\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v},$
where $(x_{*},\xi_{*},\eta_{*})\in\overline{\Omega}\times\mathbb{S}^{n}$. Applying Lemma 5.1 with (5.13), (5.14), and (5.17), we derive
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{2K+C}{\beta^{8}}\left(1+\sigma_{1}^{\frac{-1}{k-1}}\big[D^{2}u(x_{*})\big]\right).$ (5.18)
Suppose
$\frac{w}{v}(x_{*},\xi_{*},\eta_{*})>\frac{1}{\beta^{8}},$
then we have at the point $(x_{*},\xi_{*},\eta_{*})$,
$u_{\xi\xi}+2\eta u_{\xi}+\eta^{2}u$ $\geq\frac{1}{\beta^{8}}\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}|\xi|^{2}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}+\beta^{4}(\psi+\beta)\eta^{2}\right)$
$\geq\frac{1}{\beta^{8}}\left(\frac{1}{4}\beta|\xi|^{2}+\beta^{5}\eta^{2}\right)$
$\geq\frac{1}{\beta^{3}}\left(|\xi|^{2}+\eta^{2}\right)$
$=\frac{1}{\beta^{3}}.$
Since $\lambda(D^{2}u(x_{*}))\in\Gamma_{2}$, one gets at the point $x_{*}$,
$\sigma_{1}\big[\lambda\big(D^{2}u\big)\big]\geq\lambda_{\operatorname{max}}\big(D^{2}u\big)\geq u_{\xi_{*}\xi_{*}}\geq\frac{1}{\beta^{3}}-3\|u\|_{C^{1}(\overline{\Omega})}>1,$ (5.19)
where $0<\beta\ll 1$. Combining (5.18) and (5.19) yields
$\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v}\leq\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}.$
Therefore, for any $x\in\Omega$ and $\xi\in\mathbb{S}^{n-1}$,
$w(x,\xi,0)$ $\leq v(x,\xi,0)\sup_{\overline{\Omega}\times\mathbb{S}^{n}}\frac{w}{v}$
$=\left(\frac{1}{4\alpha}(\psi+\beta)^{1-\alpha}+\frac{\psi_{\xi}^{2}}{(\psi+\beta)^{\alpha}}\right)\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right)$
$\leq\frac{C_{\psi}}{\psi^{\alpha}}\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{4K+2C}{\beta^{8}}\right),$ (5.20)
where the constant $C_{\psi}>0$ depends on $\|\psi\|_{C^{3}(\overline{\Omega})}$. From the construction of $\psi^{*}=-t^{-1}\psi$ in Lemma 2.7, there exist small positive constants $\delta$ and $\delta^{*}$ such that
$\psi^{*}\leq-\delta^{*}d(x)\quad\text{in}\ \Omega_{\delta},$ (5.21)
and since $\psi^{*}$ is subharmonic,
$\psi^{*}\leq-\delta^{*}\delta\quad\text{in}\ \Omega\setminus\Omega_{\delta},$ (5.22)
where $t$, $\delta$, $\delta^{*}$ depend on $\Omega$. It follows from (5.20)–(5.22) that
$u_{\xi\xi}(x)\leq C_{*}\frac{1}{d(x)}\left(\beta^{\alpha}\sup_{\partial\Omega}u_{\nu\nu}+\frac{1}{\beta^{8}}\right),\quad\forall 0<\beta\ll 1,$
for all $x\in\Omega$ and $\xi\in\mathbb{S}^{n-1}$, where the constant $C_{*}>0$ depends on $\Omega$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Finally, taking
$\varepsilon:=C_{*}\beta^{\alpha}\quad\text{and}\quad C_{\varepsilon}:=C_{*}\beta^{-8},$
using the results in Section 3, we complete the proof of Theorem 1.2. ∎
6. The boundary estimate in terms of the interior one
Let $A=(A_{ij})$ be a skew-symmetric matrix, which implies that the matrix $\exp{A}$ is orthogonal. Define a vector field $\bar{\tau}:\ \overline{\Omega}\rightarrow\mathbb{R}^{n}$ by
$\bar{\tau}(x):=Ax.$
We adopt notations
$u_{\bar{\tau}\bar{\tau}}:=u_{ij}\bar{\tau}_{i}\bar{\tau}_{j}\quad\text{and}\quad u_{(\bar{\tau})(\bar{\tau})}:=\big(u_{\bar{\tau}}\big)_{\bar{\tau}}=u_{ij}\bar{\tau}_{i}\bar{\tau}_{j}+u_{i}\bar{\tau}_{j}\partial_{x_{j}}(\bar{\tau}_{i}).$
The following derivative exchange formula is due to Ivochkina-Trudinger-Wang [16].
Lemma 6.1 (Lemma 2.1 of [16]).
For any function $u\in C^{4}(\overline{\Omega})$, it holds that
$F^{ij}\big[D^{2}u\big]\left(u_{(\tau)(\tau)}\right)_{ij}=\left(F\big[D^{2}u\big]\right)_{(\tau)(\tau)}-F^{ij,st}\big[D^{2}u\big](u_{\tau})_{ij}(u_{\tau})_{st},$
where $\tau(x):=\vec{a}+\bar{\tau}(x)$ for any fixed vector $\vec{a}\in\mathbb{R}^{n}$.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution to the Dirichlet problem (1.1). Let the origin be a boundary point satisfying $\nu(0)=e_{n}$, and $\partial\Omega$ be represented by
$x_{n}=\rho\big(x^{\prime}\big)$ (6.1)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ for some $C^{3,1}$ function $\rho$ and constant $r_{0}>0$. It is clear that $\nabla_{x^{\prime}}\rho(0)=0$ and $x_{n}=O(|x^{\prime}|^{2})$. Let $\varphi(0)=0$ and $\nabla\varphi(0)=0$ by subtracting an affine function. It follows from (6.1) that
$\nu(x)=\frac{\left(-\rho_{1}(x^{\prime}),\dots,-\rho_{n-1}(x^{\prime}),1\right)}{\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}}$ (6.2)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. We introduce a tangential vector field $\eta$ defined by
$\eta(x):=\left(\frac{x^{\prime}}{|x^{\prime}|},\sum_{i=1}^{n-1}\frac{x_{i}}{|x^{\prime}|}\rho_{i}\big(x^{\prime}\big)\right)\quad\text{in}\ B_{r_{0}}\setminus\{0\}.$ (6.3)
We construct a vector field
$\tau=\tau(x):=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}+\begin{pmatrix}~&~&~&-\rho_{11}(0)\\ ~&0_{(n-1)\times(n-1)}\hskip-28.45274pt&~&\vdots\\ ~&~&~&-\rho_{1,n-1}(0)\\ \rho_{11}(0)\hskip-19.91692pt&\cdots\hskip-17.07164pt&\rho_{1,n-1}(0)\hskip-8.53581pt&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{n}\end{pmatrix}$ (6.4)
in $\overline{\Omega}$. Define
$w(x):=u_{(\tau)(\tau)}(x)-u_{(\tau)(\tau)}(0)\quad\text{in}\ \overline{\Omega},$
and
$M:=\sup_{x\in\partial\Omega}u_{\nu\nu}(x).$
Without loss of generality, we assume $M\gg 1$, and $C_{\varepsilon}>1$ for all $\varepsilon$ in (1.6).
Lemma 6.2.
There exist a positive constant $K_{0}$ and a linear function $h$ of $x^{\prime}$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)$ (6.5)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but are independent of $\inf_{\Omega_{0}}f$.
Proof.
For any point $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, let $\xi(x)$ be the projection of $\tau(x)$ onto the tangent plane of $\partial\Omega$ at $x$. By (6.2) and (6.4), one has
$\langle\tau,\nu\rangle=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.6)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. It is clear that for any $i\in\{1,\dots,n\}$,
$u_{i}(x)-u_{i}(0)$ $=u_{i}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)-u_{i}(0,\rho(0))$
$=e_{i}^{T}D^{2}u\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\eta\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\times\left|x^{\prime}\right|$
$=O\big(\big|x^{\prime}\big|\big),$ (6.7)
for all $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $t\in(0,1)$ depends on $x$, and the tangential vector field $\eta$ is defined in (6.3). Since $\nabla_{x^{\prime}}u(0)=\nabla_{x^{\prime}}\varphi(0)=0$, using (6.7) one gets
$|u_{\nu}(x)-u_{n}(x)|$ $\leq\sum_{i=1}^{n-1}\left|\left(u_{i}(x)-u_{i}(0)\right)\rho_{i}\big(x^{\prime}\big)\right|+\frac{|u_{n}(x)|\sum_{i=1}^{n-1}\rho_{i}^{2}(x^{\prime})}{1+\sqrt{1+\sum_{i=1}^{n-1}\rho_{i}^{2}}}$
$=O\Big(\big|x^{\prime}\big|^{2}\Big)$ (6.8)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$.
Combining (6.6), (6.4), and (6.7), one derives
$u_{(\tau)(\tau)}(x)$ $=(\xi+\langle\tau,\nu\rangle\nu)^{T}D^{2}u\left(\xi+\langle\tau,\nu\rangle\nu\right)+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=u_{\xi\xi}+u_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=u_{\xi\xi}+\rho_{11}(0)u_{n}(x)+h_{1}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$ (6.9)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where
$h_{1}\big(x^{\prime}\big):=-\sum_{i=1}^{n-1}\bigg(\rho_{1i}(0)\sum_{k=1}^{n-1}\rho_{1k}(0)u_{k}(0)\bigg)x_{i}.$
Using $\partial\Omega\in C^{3,1}$ and $u=\varphi$ on $\partial\Omega$, one gets
$u_{\xi\xi}(x)$ $=\varphi_{\xi\xi}(x)+\kappa(x,\xi(x))\left(\varphi_{\nu}(x)-u_{\nu}(x)\right)$
$=[\varphi_{11}+\rho_{11}\varphi_{n}](0)+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$
$=[u_{11}+\rho_{11}u_{n}](0)+h_{2}\big(x^{\prime}\big)+O\Big(\big|x^{\prime}\big|^{2}\Big)-\kappa(x,\xi(x))u_{\nu}(x)$ (6.10)
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $\kappa(x,\xi)$ is the curvature of $\partial\Omega$ along the direction $\xi$ at $x\in\partial\Omega$, and $h_{2}(x^{\prime})$ is the linear part in the Taylor expansion of $\varphi_{\xi\xi}(x)+\kappa(x,\xi(x))\varphi_{\nu}(x)$ with respect to $x^{\prime}$. By (6.9), (6), (6.8), and (6.7), we have
$w(x)$ $=u_{(\tau)(\tau)}(x)-[u_{11}+\rho_{11}u_{n}](0)$
$=\rho_{11}(0)u_{n}(x)-\kappa(x,\xi(x))u_{\nu}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(x)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=\left[\rho_{11}(0)-\kappa(x,\xi(x))\right]u_{n}(0)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big)$
$=h_{3}\big(x^{\prime}\big)+h_{1}+h_{2}+O\Big(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\Big),$
on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h_{3}(x^{\prime})$ is the linear part in the Taylor expansion of $\kappa(x,\xi(x))u_{n}(0)$ with respect to $x^{\prime}$.
Consequently, there exists a constant $K_{0}>0$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)$
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $h(x^{\prime}):=h_{1}(x^{\prime})+h_{2}(x^{\prime})+h_{3}(x^{\prime})$; $K_{0}$ and the coefficients of $h$ depend on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. Using the results in Section 3, we complete the proof of Lemma 6.2. ∎
Define the linear elliptic operator $\mathcal{L}$ by
$\mathcal{L}(w):=F^{ij}\big[D^{2}u\big]w_{ij}.$
Define
$\omega_{r}:=\left\{x\in\Omega:\ \rho\big(x^{\prime}\big)<x_{n}<\rho\big(x^{\prime}\big)+r^{4},\ \big|x^{\prime}\big|<r\right\},\quad\forall r\leq r_{0}.$
6.1. The case under condition $(i)$ or $(ii)$
In this subsection, we assume condition (1.7) is satisfied, and either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ holds or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold.
Lemma 6.3.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. We have two conclusions:
$(i)$ If $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, then
$\mathcal{L}(w)\geq-C_{1}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$
where $C_{1}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
$(ii)$ If $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, then
$\mathcal{L}(w)\geq-C_{2}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$
where $C_{2}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
From Lemma 6.1, we have
$\mathcal{L}(w)$ $=F^{ij}\big[D^{2}u\big]\left(u_{(\tau)(\tau)}\right)_{ij}$
$\geq\left(g^{\frac{p}{k}}\right)_{\tau\tau}+\left(g^{\frac{p}{k}}\right)_{i}\tau_{j}\partial_{x_{j}}(\tau_{i})$
$=\frac{p}{k}g^{\frac{p}{k}-1}\left(\partial_{\tau\tau}g-\frac{k-p}{k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+\left(\frac{p}{k}g^{\frac{p}{k}-\frac{1}{2}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$ (6.11)
in $\Omega$. When $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it follows from (6.11) and (2.1) that
$\mathcal{L}(w)$ $\geq-\frac{p}{k}g^{-\frac{1}{k}}\left|\partial_{\tau\tau}g-\frac{k-p}{k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right|-g^{-\frac{1}{k}}\left|\frac{p}{k}\partial_{x_{i}}g\tau_{j}\partial_{x_{j}}(\tau_{i})\right|$
$\geq-C_{1}f^{\frac{-1}{k(k-1)}}$
in $\Omega$, where $C_{1}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
When $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$, which implies $(k+2)/3k<1/2$, it follows from (6.11), (2.2), and (2.1) that
$\mathcal{L}(w)$ $\geq\frac{p}{k}g^{\frac{-p}{k(k-1)}-\frac{1}{3}}\left(\partial_{\tau\tau}g-\frac{k+2}{3k}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+g^{\frac{-p}{k(k-1)}}\left(\frac{p}{k}g^{\frac{1}{6}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$
$\geq-C_{2}f^{\frac{-1}{k(k-1)}}$
in $\Omega$, where $C_{2}>0$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. This completes the proof of Lemma 6.3. ∎
Lemma 6.4.
There exist positive constants $r_{2}$ and $K_{2}$ such that
$w(x)-h\big(x^{\prime}\big)\leq K_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)+K_{2}M\left(x_{n}-\rho\big(x^{\prime}\big)\right)\quad\text{in}\ \omega_{r_{2}},$
where $r_{2}$ depends on $\Omega$; and the constant $K_{2}$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Define a function
$v:=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}-K_{0}\big|x^{\prime}\big|^{4}$
in $\omega_{r_{0}}$, where the constant $\beta_{1}>0$ is to be determined. It holds that
$D^{2}v$ $=D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big(x^{\prime}\big)\big]+D^{2}\big[\rho^{2}\big(x^{\prime}\big)-2x_{n}\rho\big(x^{\prime}\big)\big]$
$\hskip 12.80373pt-\frac{K_{0}}{M}D^{2}\Big[\big|x^{\prime}\big|^{2}\Big]-D^{2}\Big[K_{0}\big|x^{\prime}\big|^{4}\Big].$
Since $\partial\Omega\in C^{3,1}$ is strictly (k1)(k-1)-convex, one can fix $\beta_{1}$ sufficiently small such that $\lambda(D^{2}[x_{n}^{2}+\beta_{1}\rho](0))$ belongs to a compact subset of $\Gamma_{k}$ independent of $0\in\partial\Omega$. By
$|x|\rightarrow 0,\ \rho\big(x^{\prime}\big)\rightarrow 0,\ \nabla_{x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow 0,\ D^{2}_{x^{\prime}x^{\prime}}\rho\big(x^{\prime}\big)\rightarrow D^{2}_{x^{\prime}x^{\prime}}\rho(0)\quad\text{as}\ r\rightarrow 0,$
there exist two small positive constants $r_{2}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{2}$, and a large constant $C_{1}>0$ such that
$\left\{\lambda\big(D^{2}v-\delta_{2}\mathbf{I}_{n}\big):\ x\in\omega_{r_{2}}\right\}\ \ \text{is a perturbation of}\ \ \lambda\big(D^{2}\big[x_{n}^{2}+\beta_{1}\rho\big](0)\big)$
for all $M\geq C_{1}$, and is thus contained in a compact subset of $\Gamma_{k}$ for all $M\geq C_{1}$. We emphasize that $\beta_{1}$, $r_{2}$, $\delta_{2}$, $C_{1}$ all depend only on $\Omega$. Using the concavity of $F$, we have
$F^{ij}\big[D^{2}u\big]\left(C_{2}(v_{ij}-\delta_{2}\delta_{ij})-u_{ij}\right)\geq C_{2}\sigma_{k}^{\frac{1}{k}}\big[D^{2}v-\delta_{2}\mathbf{I}_{n}\big]-f^{\frac{1}{k}}>0$
in $\omega_{r_{2}}$, where $C_{2}$ is sufficiently large. Then by virtue of (2.14) and (2.13), we derive
[…]
in $\omega_{r_{2}}$, where $C>0$ is universal. Thus using (6.12) and Lemma 6.3, there exists a large constant $K>0$ such that
$\mathcal{L}(-KMv)<\mathcal{L}(-Kv)<\mathcal{L}(w-h)\quad\text{in}\ \omega_{r_{1}},$ (6.13)
where $K$ depends on $\delta_{0}$, $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
The boundary $\partial\omega_{r_{2}}$ consists of three parts: $\partial_{1}\omega_{r_{2}}\cup\partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}}$, where $\partial_{1}\omega_{r_{2}}$ and $\partial_{2}\omega_{r_{2}}$ are respectively the graph parts of $\rho$ and $\rho+r_{2}^{4}$, and $\partial_{3}\omega_{r_{2}}$ is the boundary part on $\{|x^{\prime}|=r_{2}\}$. On $\partial_{1}\omega_{r_{2}}$, using (6.5) one gets
$KMv=KK_{0}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h.$
It follows from the results in Section 3 that
$w-h\leq C_{3}M+C_{4}\quad\text{on}\ \partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}},$
where $C_{3}$ depends on $\Omega$; $C_{4}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. It holds that
$KMv\geq KM\left(\beta_{1}r_{2}^{4}-r_{2}^{8}\right)\geq\frac{\beta_{1}}{2}KMr_{2}^{4}\quad\text{on}\ \partial_{2}\omega_{r_{2}},$
and
$KMv\geq KK_{0}\left(r_{2}^{2}+Mr_{2}^{4}\right)\quad\text{on}\ \partial_{3}\omega_{r_{2}}.$
Then we obtain $KMv\geq w-h$ on $\partial_{2}\omega_{r_{2}}\cup\partial_{3}\omega_{r_{2}}$ by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
Therefore, it follows from the maximum principle that
$KMv\geq w-h\quad\text{in}\ \omega_{r_{2}},$
which implies
$w-h\leq KM\left(\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)+\frac{K_{0}}{M}\big|x^{\prime}\big|^{2}+K_{0}\big|x^{\prime}\big|^{4}\right)$
in $\omega_{r_{2}}$. This completes the proof of Lemma 6.4. ∎
Lemma 6.5.
For any $\sigma>0$, there exists a positive constant $K_{\sigma}$ such that
$\left(u_{(\tau)(\tau)}\right)_{n}(0)\leq\sigma M+K_{\sigma},$
where constant $K_{\sigma}$ depends on $\sigma$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
Let $v$ be the function defined in the proof of Lemma 3.2, which is
$v=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{1}}.$
Recalling (3.8), we derive
[…]
where $C>0$ is universal. We emphasize that $\beta_{1}$, $\beta_{2}$, $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{1}$ all depend only on $\Omega$. It follows from Lemma 6.3 and (6.14) that there exists a large constant $K>0$ such that
$\mathcal{L}(-Kv)<\mathcal{L}(w-h)\quad\text{in}\ \omega_{r_{1}},$ (6.15)
where constant $K$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
For any $r\leq\min\{r_{1},r_{2}\}$, the boundary of $\omega_{r}$ consists of three parts: $\partial_{1}\omega_{r}\cup\partial_{2}\omega_{r}\cup\partial_{3}\omega_{r}$, where $\partial_{1}\omega_{r}$ and $\partial_{2}\omega_{r}$ are respectively the graph parts of $\rho$ and $\rho+r^{4}$, and $\partial_{3}\omega_{r}$ is the boundary part on $\{|x^{\prime}|=r\}$. Using (6.5), one gets
$K(rM+1)v=K\beta_{2}(rM+1)\big|x^{\prime}\big|^{2}\geq K\beta_{2}\left(\big|x^{\prime}\big|^{2}+M\big|x^{\prime}\big|^{4}\right)\geq w-h$ (6.16)
on $\partial_{1}\omega_{r}$, by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from Lemma 6.4 that
$w-h\leq 2K_{2}Mr^{4}+K_{2}r^{2}\quad\text{on}\ \partial_{3}\omega_{r}.$
We have
$K(rM+1)v\geq K(rM+1)\beta_{2}r^{2}\geq K\beta_{2}Mr^{4}+K\beta_{2}r^{2}\geq w-h$ (6.17)
on $\partial_{3}\omega_{r}$, by taking $K$ large, which depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$.
Without loss of generality, we assume $|\nabla_{x^{\prime}}\rho|\leq 1$ on $\{|x^{\prime}|\leq r_{0}\}$. The distance function satisfies
$\operatorname{dist}(x,\partial\Omega)\geq\inf_{|y^{\prime}|\leq r_{0}}\left(\big|x^{\prime}-y^{\prime}\big|+r^{4}-\big|\nabla_{x^{\prime}}\rho\cdot\big(x^{\prime}-y^{\prime}\big)\big|\right)\geq r^{4},\quad\forall x\in\partial_{2}\omega_{r}.$
By virtue of the weakly interior estimate, namely Theorem 1.2, we infer
$w-h\leq\left|u_{ij}\tau_{i}\tau_{j}\right|+C_{1}$ $\leq 2\sup_{\Omega_{r^{4}}}\left|D^{2}u\right|+C_{1}$
$\leq\frac{2}{r^{4}}\left(r^{9}\sup_{\partial\Omega}\left|D^{2}u\right|+C_{r^{9}}\right)+C_{1}$
$\leq 2Mr^{5}+\frac{2C_{r^{9}}}{r^{4}}+C_{1}$
on $\partial_{2}\omega_{r}$, where $C_{r^{9}}$ is given in (1.6); and $C_{1}>0$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Then we obtain
$-K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)\frac{\beta_{1}}{2}r^{4}$ $=\frac{K\beta_{1}}{2}Mr^{5}+\frac{K\beta_{1}}{2}\frac{C_{r^{9}}}{r^{4}}$
$\geq w-h$ (6.18)
on $\partial_{2}\omega_{r}$, by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Note that $\mathcal{L}(v)>0$ and $v<0$ in $\omega_{r}$. Applying the maximum principle with (6.15)–(6.1) yields
$K\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)v\geq w-h\quad\text{in}\ \omega_{r},\quad\forall r\leq\min\{r_{1},r_{2}\},$
where the constant $K$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. Therefore,
$\left(u_{(\tau)(\tau)}\right)_{n}(0)=\lim_{x_{n}\rightarrow 0^{+}}\frac{w(0,x_{n})-h(0)}{x_{n}}\leq K\beta_{1}\left(rM+\frac{C_{r^{9}}}{r^{8}}\right)$
for all $r\leq\min\{r_{1},r_{2}\}$. Let
$\sigma:=K\beta_{1}r\quad\text{and}\quad K_{\sigma}:=K\beta_{1}\frac{C_{r^{9}}}{r^{8}}.$
This completes the proof of Lemma 6.5. ∎
Lemma 6.6.
It holds that
$\left(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right)(x)\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$
for all points $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, where $K_{\sigma}$ is given in Lemma 6.5.
Proof.
It is clear that the expression $(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i}))(x)$ is invariant under orthogonal transformations.
Step 1. Let $\tilde{\eta}(x)$ be an arbitrary tangential vector field defined in $B_{r}(0)$, where the radius $r>0$ can be arbitrarily small depending on the origin $0\in\partial\Omega$. Furthermore, let $\tilde{\eta}$ satisfy
$\tilde{\eta}(0)=e_{1}\quad\text{and}\quad\left|\partial_{x_{1}}(\tilde{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$
for some constant $C$.
Differentiating $\nu\cdot\tilde{\eta}(x^{\prime},\rho(x^{\prime}))=0$ with respect to $x_{1}$ at the origin yields
$\partial_{x_{1}}(\tilde{\eta}_{n})(0)-\rho_{11}(0)=0.$
Then we have
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)=u_{n11}(0)+u_{ni}(0)\partial_{x_{1}}(\tilde{\eta}_{i})(0)\leq u_{n11}(0)+u_{nn}(0)\rho_{11}(0)+K_{1},$
where $K_{1}$ depends on $C$ and tangential-normal derivatives of $u$ on $\partial\Omega$. It holds that
$\left(u_{(\tau)(\tau)}\right)_{n}(0)$ $=u_{n11}(0)+u_{in}(0)\tau_{j}\partial_{x_{j}}(\tau_{i})(0)$
$\hskip 12.80373pt+u_{ij}\partial_{x_{n}}[\tau_{i}\tau_{j}]+u_{i}\partial_{x_{n}}\big[\tau_{j}\partial_{x_{j}}(\tau_{i})\big]$
$\geq u_{n11}(0)+u_{nn}(0)\rho_{11}(0)-K_{2},$
where $K_{2}$ depends on $\Omega$, $\|u\|_{C^{1}(\overline{\Omega})}$, second-order tangential and tangential-normal derivatives of $u$ on $\partial\Omega$. As a conclusion,
$u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\left(u_{(\tau)(\tau)}\right)_{n}(0)+K_{1}+K_{2}\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$ (6.19)
by Lemma 6.5 and increasing $K_{\sigma}$.
Step 2. For any $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, by rotation and translation, we introduce a new coordinate system $y$ such that $x_{0}$ is the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, and $\eta(x_{0})$ is in the direction $e_{1}$. Denote the expression of vector field $\eta(x)$ in the new coordinate system by $\bar{\eta}(y)$. It is clear that $\bar{\eta}$ is well-defined in $B_{|x_{0}|/2}(0)$. Recalling (6.3), by calculation, one can verify
$\left|\partial_{y_{1}}(\bar{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1,$
where $C$ depends only on the $C^{2}$-norm of $\partial\Omega$.
Define $\tilde{\eta}(y):=\bar{\eta}(y)/|\eta(x_{0})|$. Then
$\tilde{\eta}(0)=e_{1}\quad\text{and}\quad\left|\partial_{y_{1}}(\tilde{\eta}_{i})(0)\right|\leq C,\quad\forall 1\leq i\leq n-1.$
By (6.19) and $|\eta(x_{0})|\leq 2$, we get
$\left(u_{nij}\bar{\eta}_{i}\bar{\eta}_{j}+u_{ni}\bar{\eta}_{j}\partial_{y_{j}}(\bar{\eta}_{i})\right)(0)\leq 4u_{n(\tilde{\eta})(\tilde{\eta})}(0)\leq\sigma M+4K_{\sigma/4}$
for all $\sigma>0$. Coming back to the original coordinate system (6.1), by increasing $K_{\sigma}$, we derive
$\left(u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right)(x_{0})\leq\sigma M+K_{\sigma},\quad\forall\sigma>0,$
for all $x_{0}\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$. The proof of Lemma 6.6 is complete. ∎
We now proceed to establish the boundary second-order normal derivative estimate.
Proof of Theorem 1.3.
Without loss of generality, let $M$ be attained at the origin $0\in\partial\Omega$, $e_{n}=\nu(0)$, (6.1) hold, $\varphi(0)=0$, and $\nabla\varphi(0)=0$.
Applying Taylor expansion to $h(t):=u_{n}(tx^{\prime},\rho(tx^{\prime}))$, we obtain that for any $x\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$,
$u_{n}\big(x^{\prime},\rho\big(x^{\prime}\big)\big)-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}=\frac{1}{2}u_{n(\eta)(\eta)}\big(tx^{\prime},\rho\big(tx^{\prime}\big)\big)\left|x^{\prime}\right|^{2},$ (6.20)
where $t\in(0,1)$ depends on $x$, and $\eta$ is defined in (6.3). Differentiating $u=\varphi$ on $\partial\Omega$ three times, yields that the third-order tangential derivatives of $u$ are bounded. Using Lemma 6.6, we derive
$u_{n(\eta)(\eta)}(y)$ $=\left[u_{n\eta\eta}+u_{ni}\eta_{j}\partial_{x_{j}}(\eta_{i})\right](y)$
$=\langle e_{n},\nu\rangle\left[u_{\nu ij}\eta_{i}\eta_{j}+u_{\nu i}\eta_{j}\partial_{x_{j}}(\eta_{i})\right](y)+C_{1}$
$\leq\sigma M+K_{\sigma}+C_{1}$ (6.21)
for all $y\neq 0$ on $\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$ and for all $\sigma>0$, where $C_{1}$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. It follows from (6.20) and (6.1) that
$u_{n}(x)-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\leq(\sigma M+K_{\sigma})\big|x^{\prime}\big|^{2}$ (6.22)
for all $x\in\partial\Omega\cap\{|x^{\prime}|\leq r_{0}\}$, by increasing $K_{\sigma}$.
Let $v$ be the function defined in the proof of Lemma 3.2, which is
$v=\big(x_{n}-\rho\big(x^{\prime}\big)\big)^{2}-\beta_{1}\big(x_{n}-\rho\big(x^{\prime}\big)\big)-\beta_{2}\big|x^{\prime}\big|^{2}\quad\text{in}\ \omega_{r_{1}}.$
Recalling (3.8), we derive
[…]
where $C>0$ is universal. We emphasize that $\beta_{1}$, $\beta_{2}$, $r_{1}\leq\min\{r_{0},\beta_{1}\}$, $\delta_{1}$ all depend only on $\Omega$. By (2.10) and (2.1), we infer
$\mathcal{L}(u_{n})=\left(\frac{p}{k}g^{\frac{p}{k-1}-\frac{1}{2}}\frac{\partial_{n}g}{\sqrt{g}}\right)f^{\frac{-1}{k(k-1)}}\geq-C_{2}f^{\frac{-1}{k(k-1)}}\quad\text{in}\ \Omega,$ (6.24)
where the constant $C_{2}$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Define
$\tilde{v}:=-K(\sigma M+K_{\sigma})v\quad\text{in}\ \overline{\omega_{r_{1}}},$
where $K>0$ is a constant to be determined. By (6.23) and (6.24), we obtain
$\mathcal{L}(\tilde{v})\leq\mathcal{L}\left(u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\right)\quad\text{in}\ \omega_{r_{1}},$
by taking the constant $K$ large, which depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
The boundary $\partial\omega_{r_{1}}$ consists of three parts: $\partial_{1}\omega_{r_{1}}\cup\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$, defined in the proof of Lemma 3.2. Using (6.22), one has
$\tilde{v}=K\beta_{2}(\sigma M+K_{\sigma})\big|x^{\prime}\big|^{2}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{on}\ \partial_{1}\omega_{r_{1}},$
by taking $K$ large, which depends on $\Omega$. It is clear that
$\tilde{v}\geq K(\sigma M+K_{\sigma})\min\left\{\frac{\beta_{1}}{2}r_{1}^{4},\beta_{2}r_{1}^{2}\right\}$
on $\partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}}$. According to the results in Section 3, we get
$\tilde{v}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{on}\ \partial_{2}\omega_{r_{1}}\cup\partial_{3}\omega_{r_{1}},$
by taking $K$ large, which depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$. Hence applying the comparison principle yields
$\tilde{v}\geq u_{n}-u_{n}(0)-\sum_{i=1}^{n-1}u_{ni}(0)x_{i}\quad\text{in}\ \omega_{r_{1}},$
where the constant $K$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$.
Therefore,
$M=u_{nn}(0)=\lim_{x_{n}\rightarrow 0^{+}}\frac{u_{n}(0,x_{n})-u_{n}(0)}{x_{n}}\leq\partial_{x_{n}}\tilde{v}(0)=K\beta_{1}(\sigma M+K_{\sigma}),$
where $K_{\sigma}$ is given in Lemma 6.5. Taking $\sigma=1/(2K\beta_{1})$, we obtain the upper bound of $M$. Then the $C^{2}$ estimate independent of $\inf_{\Omega_{0}}f$ is established, and Theorem 1.3 follows from an approximation argument. ∎
6.2. The case under condition $(iii)$
In this subsection, we assume $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. In fact, Lemmas 2.6 and 6.3 can be improved.
Lemma 6.7.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$, and $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. Then
$\operatorname{tr}\left(GD^{2}u_{\xi\xi}\right)\geq-K|\xi|^{2},\quad\forall\xi\in\mathbb{R}^{n},$
in $\Omega$, where $K$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
It follows from (2.11), (2.2), and (2.13) that
$\operatorname{tr}\left(GD^{2}u_{ee}\right)\geq\frac{1}{\operatorname{tr}F^{ij}}\frac{2}{3}g^{-\frac{1}{3}}\left(\partial_{ee}g-\frac{1}{3}\frac{\left|\partial_{e}g\right|^{2}}{g}\right)\geq-\frac{K}{\operatorname{tr}F^{ij}}\geq-CK$
for all $e\in\mathbb{S}^{n-1}$, where $g=f^{3/(2k)}$, $C>0$ is universal, and $K>0$ depends on $\operatorname{dist}(\Omega,\partial\Omega_{0})$ and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. ∎
Lemma 6.8.
Let $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$, and $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$. Then
$\mathcal{L}(w)\geq-C\quad\text{in}\ \Omega,$
where $C$ depends on $\Omega$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Proof.
It follows from (6.11) that
$\mathcal{L}(w)\geq\frac{2}{3}g^{-\frac{1}{3}}\left(\partial_{\tau\tau}g-\frac{1}{3}\frac{\left|\partial_{\tau}g\right|^{2}}{g}\right)+\left(\frac{2}{3}g^{\frac{1}{6}}\frac{\partial_{x_{i}}g}{\sqrt{g}}\right)\tau_{j}\partial_{x_{j}}(\tau_{i})$
in $\Omega$, where $g=f^{3/(2k)}$. Using Lemma 2.1, we obtain
$\mathcal{L}(w)\geq-C,$
where $C$ depends on $\|\rho\|_{C^{3}}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$. ∎
Using Lemmas 6.7 and 6.8, with the same proof as in subsection 6.1, we prove Theorem 1.3.